GNU Findutils

Finding files

version 4.10.0, 27 May 2024

by David MacKenzie and James Youngman

This manual documents version 4.10.0 of the GNU utilities for finding files that match certain criteria and performing various operations on them. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994–2024 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Scope . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 Finding Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2.1 find Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2 Starting points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

2.3.1 Base Name Patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3.2 Full Name Patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3.3 Fast Full Name Search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.3.4 Shell Pattern Matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2.4 Links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.4.1 Symbolic Links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.4.2 Hard Links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

2.5 Time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.5.1 Age Ranges . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.2 Comparing Timestamps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

2.6 Size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.7 Type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.8 Owner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.9 File Mode Bits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.10 Contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.11 Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.12 Filesystems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.13 Combining Primaries With Operators . . . . . . . . . . . . . . . . . . . . . . . . . 22

3 Actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

3.1 Print File Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.2 Print File Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

3.2.1 Escapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.2.2 Format Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

3.2.2.1 Name Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.2.2.2 Ownership Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.2.2.3 Size Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.2.2.4 Location Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.2.2.5 Time Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.2.2.6 Other Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.2.2.7 Reserved and Unknown Directives . . . . . . . . . . . . . . . . . . . . 28

3.2.3 Time Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.2.3.1 Time Components . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.2.3.2 Date Components . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

3.2.3.3 Combined Time Formats. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.2.4 Formatting Flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

3.3 Run Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.3.1 Single File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 3.3.2 Multiple Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

3.3.2.1 Unsafe File Name Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 3.3.2.2 Safe File Name Handling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 3.3.2.3 Unusual Characters in File Names . . . . . . . . . . . . . . . . . . . . 33 3.3.2.4 Limiting Command Size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 3.3.2.5 Controlling Parallelism. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 3.3.2.6 Interspersing File Names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

3.3.3 Querying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.4 Delete Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 3.5 Adding Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

4 File Name Databases . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

4.1 Database Locations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.2 Database Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

4.2.1 LOCATE02 Database Format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.2.2 Sample LOCATE02 Database. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.2.3 slocate Database Format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.2.4 Old Database Format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

4.3 Newline Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

5 File Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

5.1 Structure of File Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.2 Symbolic Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

5.2.1 Setting Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 5.2.2 Copying Existing Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.2.3 Changing Special Permissions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.2.4 Conditional Executability. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 5.2.5 Making Multiple Changes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 5.2.6 The Umask and Protection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

5.3 Numeric Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

6 Date input formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

6.1 General date syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 6.2 Calendar date items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 6.3 Time of day items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 6.4 Time zone items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 6.5 Combined date and time of day items . . . . . . . . . . . . . . . . . . . . . . . . . 54 6.6 Day of week items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 6.7 Relative items in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 6.8 Pure numbers in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 6.9 Seconds since the Epoch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 6.10 Specifying time zone rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 6.11 Authors of parse_datetime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

7 Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

7.1 Leaf Optimisation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 7.2 d type Optimisation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

8 Reference. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

8.1 Invoking find . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 8.1.1 Filesystem Traversal Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 8.1.2 Warning Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 8.1.3 Optimisation Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 8.1.4 Debug Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 8.1.5 Find Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

8.2 Invoking locate. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 8.3 Invoking updatedb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 8.4 Invoking xargs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

8.4.1 xargs options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 8.4.2 Conflicting options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 8.4.3 Invoking the shell from xargs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

8.5 Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 8.5.1 ‘findutils-default’ regular expression syntax . . . . . . . . . . . 71 8.5.2 ‘emacs’ regular expression syntax . . . . . . . . . . . . . . . . . . . . . . . . . 72 8.5.3 ‘gnu-awk’ regular expression syntax. . . . . . . . . . . . . . . . . . . . . . . 73 8.5.4 ‘grep’ regular expression syntax . . . . . . . . . . . . . . . . . . . . . . . . . . 74 8.5.5 ‘posix-awk’ regular expression syntax . . . . . . . . . . . . . . . . . . . . 75 8.5.6 ‘awk’ regular expression syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 8.5.7 ‘posix-basic’ regular expression syntax . . . . . . . . . . . . . . . . . . 77 8.5.8 ‘posix-egrep’ regular expression syntax . . . . . . . . . . . . . . . . . . 78 8.5.9 ‘egrep’ regular expression syntax . . . . . . . . . . . . . . . . . . . . . . . . . 79 8.5.10 ‘posix-extended’ regular expression syntax . . . . . . . . . . . . . . 79

8.6 Environment Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

9 Common Tasks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

9.1 Viewing And Editing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 9.2 Archiving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 9.3 Cleaning Up. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 9.4 Strange File Names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 9.5 Fixing Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 9.6 Classifying Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

10 Worked Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

10.1 Deleting Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 10.1.1 The Traditional Way . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 10.1.2 Making Use of xargs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 10.1.3 Unusual characters in filenames . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 10.1.4 Going back to -exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 10.1.5 A more secure version of -exec . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 10.1.6 Using the -delete action. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 10.1.7 Improving things still further . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 10.1.8 Conclusion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

10.2 Copying A Subset of Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 10.3 Updating A Timestamp File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

10.3.1 Updating the Timestamp The Wrong Way . . . . . . . . . . . . . . . . 90 10.3.2 Using the test utility to compare timestamps. . . . . . . . . . . . . . 90 10.3.3 A combined approach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 10.3.4 Using -printf and sort to compare timestamps. . . . . . . . . . 90 10.3.5 Solving the problem with make . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 10.3.6 Coping with odd filenames too. . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

10.4 Finding the Shallowest Instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

11 Security Considerations. . . . . . . . . . . . . . . . . . . . . . . 93

11.1 Levels of Risk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 11.2 Security Considerations for find. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

11.2.1 Problems with -exec and filenames . . . . . . . . . . . . . . . . . . . . . . . 94 11.2.2 Changing the Current Working Directory . . . . . . . . . . . . . . . . . 95

11.2.2.1 O_NOFOLLOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 11.2.2.2 Systems without O_NOFOLLOW . . . . . . . . . . . . . . . . . . . . . . . . 96

11.2.3 Race Conditions with -exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 11.2.4 Race Conditions with -print and -print0 . . . . . . . . . . . . . . . 97

11.3 Security Considerations for xargs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 11.4 Security Considerations for locate . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

11.4.1 Race Conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 11.5 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 11.6 Further Reading on Security. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

12 Error Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

12.1 Error Messages From find . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 12.2 Error Messages From xargs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 12.3 Error Messages From locate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 12.4 Error Messages From updatedb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

13 History. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

13.1 History of find . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 13.2 History of xargs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 13.3 History of locate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

Appendix A GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 find Primary Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

1 Introduction

This manual shows how to find files that meet criteria you specify, and how to perform various actions on the files that you find. The principal programs that you use to perform these tasks are find, locate, and xargs. Some of the examples in this manual use capabilities specific to the GNU versions of those programs.

See Chapter 13 [History], page 104, for a history of find, locate and xargs. The current maintainers of GNU findutils (and this manual) are Bernhard Voelker and James Youngman. Many other people have contributed bug fixes, small improvements, and helpful suggestions. Thanks!

To report a bug in GNU findutils, please use the form on the Savannah web site at

https://savannah.gnu.org/bugs/?group=findutils. Reporting bugs this way means

that you will then be able to track progress in fixing the problem.

If you don’t have web access, you can also just send mail to the mailing list. The mailing list bug-findutils@gnu.org carries discussion of bugs in findutils, questions and answers about the software and discussion of the development of the programs. To join the list,

send email to bug-findutils-request@gnu.org.

Please read any relevant sections of this manual before asking for help on the mailing list. You may also find it helpful to read the NON-BUGS section of the find manual page.

If you ask for help on the mailing list, people will be able to help you much more effectively if you include the following things:

• The version of the software you are running. You can find this out by running ‘locate

    --version’.

• What you were trying to do

• The exact command line you used

• The exact output you got (if this is very long, try to find a smaller example which exhibits the same problem)

• The output you expected to get

It may also be the case that the bug you are describing has already been fixed, if it is a bug. Please check the most recent findutils releases at ftp://ftp.gnu.org/gnu/findutils and, if possible, the development branch at ftp://alpha.gnu.org/gnu/findutils. If you take the time to check that your bug still exists in current releases, this will greatly help people who want to help you solve your problem. Please also be aware that if you obtained findutils as part of the GNU/Linux ’distribution’, the distributions often lag seriously behind findutils releases, even the stable release. Please check the GNU FTP site.

1.1 Scope

For brevity, the word file in this manual means a regular file, a directory, a symbolic link, or any other kind of node that has a directory entry. A directory entry is also called a file name. A file name may contain some, all, or none of the directories in a path that leads to the file. These are all examples of what this manual calls “file names”:

     parser.c
     README

   ./budget/may-94.sc
   fred/.cshrc
   /usr/local/include/termcap.h

A directory tree is a directory and the files it contains, all of its subdirectories and the files they contain, etc. It can also be a single non-directory file.

These programs enable you to find the files in one or more directory trees that:

• have names that contain certain text or match a certain pattern;

• are links to certain files;

• were last used during a certain period of time;

• are within a certain size range;

• are of a certain type (regular file, directory, symbolic link, etc.);

• are owned by a certain user or group;

• have certain access permissions or special mode bits;

• contain text that matches a certain pattern;

• are within a certain depth in the directory tree;

• or some combination of the above. Once you have found the files you’re looking for (or files that are potentially the ones

you’re looking for), you can do more to them than simply list their names. You can get any combination of the files’ attributes, or process the files in many ways, either individually or in groups of various sizes. Actions that you might want to perform on the files you have found include, but are not limited to:

• view or edit

• store in an archive

• remove or rename

• change access permissions

• classify into groups This manual describes how to perform each of those tasks, and more.

1.2 Overview

The principal programs used for making lists of files that match given criteria and running commands on them are find, locate, and xargs. An additional command, updatedb, is used by system administrators to create databases for locate to use.

find searches for files in a directory hierarchy and prints information about the files it found. It is run like this:

   find [file...] [expression]

Here is a typical use of find. This example prints the names of all files in the directory tree rooted in /usr/src whose name ends with ‘.c’ and that are larger than 100 KiB.

   find /usr/src -name '*.c' -size +100k -print

Notice that the wildcard must be enclosed in quotes in order to protect it from expansion by the shell.

locate searches special file name databases for file names that match patterns. The system administrator runs the updatedb program to create the databases. locate is run like this:

     locate [option...] pattern...

This example prints the names of all files in the default file name database whose name ends with ‘Makefile’ or ‘makefile’. Which file names are stored in the database depends on how the system administrator ran updatedb.

     locate '*[Mm]akefile'

The name xargs, pronounced EX-args, means “combine arguments.” xargs builds and executes command lines by gathering together arguments it reads on the standard input. Most often, these arguments are lists of file names generated by find. xargs is run like this:

     xargs [option...] [command [initial-arguments]]

The following command searches the files listed in the file file-list and prints all of the lines in them that contain the word ‘typedef’.

     xargs grep typedef < file-list

2 Finding Files

By default, find prints to the standard output the names of the files that match the given criteria. See Chapter 3 [Actions], page 23, for how to get more information about the matching files.

2.1 find Expressions

The expression that find uses to select files consists of one or more primaries, each of which is a separate command line argument to find. find evaluates the expression each time it processes a file. An expression can contain any of the following types of primaries:

options affect overall operation rather than the processing of a specific file;

tests return a true or false value, depending on the file’s attributes;

actions have side effects and return a true or false value; and

operators connect the other arguments and affect when and whether they are evaluated.

You can omit the operator between two primaries; it defaults to ‘-and’. See Section 2.13 [Combining Primaries With Operators], page 22, for ways to connect primaries into more complex expressions.

The ‘-print’ action is performed on all files for which the entire expression is true (see Section 3.1 [Print File Name], page 23), unless the expression contains an action other than ‘-prune’ or ‘-quit’. Actions which inhibit the default ‘-print’ are ‘-delete’, ‘-exec’,

‘-execdir’, ‘-ok’, ‘-okdir’, ‘-fls’, ‘-fprint’, ‘-fprintf’, ‘-ls’, ‘-print’ and ‘-printf’.

Options take effect immediately, rather than being evaluated for each file when their place in the expression is reached. Therefore, for clarity, it is best to place them at the beginning of the expression. There are exceptions to this; ‘-regextype’, ‘-daystart’ and ‘-follow’ have different effects depending on where in the command line they appear. This can be confusing, so it’s best to keep them at the beginning, too.

Many of the primaries take arguments, which immediately follow them in the next command line argument to find. Some arguments are file names, patterns, or other strings; others are numbers. Numeric arguments can be specified as

+n for greater than n,

-n for less than n,

n for exactly n.

2.2 Starting points

GNU find searches the directory tree rooted at each given starting-point by evaluating the given expression from left to right, according to the rules of operator precedence, until the outcome is known (the left hand side is false for ‘and’ operations, true for ‘or’), at which point find moves on to the next file name.

If no starting-point is specified, the current directory ‘.’ is assumed.

A double dash ‘--’ could theoretically be used to signal that any remaining arguments are not options, but this does not really work due to the way find determines the end of the list of starting point arguments: it does that by reading until an expression argument comes (which also starts with a ‘-’). Now, if a starting point argument would begin with a ‘-’, then find would treat it as expression argument instead. Thus, to ensure that all start points are taken as such, and especially to prevent that wildcard patterns expanded by the calling shell are not mistakenly treated as expression arguments, it is generally safer to prefix wildcards or dubious path names with either ‘./’, or to use absolute path names starting with ‘/’.

Alternatively, it is generally safe though non-portable to use the GNU option ‘-files0-from’ to pass arbitrary starting points to find.

[Option] -files0-from file

Read the starting points from file instead of getting them on the command line. In contrast to the known limitations of passing starting points via arguments on the command line, namely the limitation of the amount of file names, and the inherent ambiguity of file names clashing with option names, using this option allows to safely pass an arbitrary number of starting points to find. Using this option and passing starting points on the command line is mutually exclusive, and is therefore not allowed at the same time. The file argument is mandatory. One can use ‘-files0-from -’ to read the list of starting points from the standard input stream, and e.g. from a pipe. In this case, the actions ‘-ok’ and ‘-okdir’ are not allowed, because they would obviously interfere with reading from standard input in order to get a user confirmation. The starting points in file have to be separated by ASCII NUL characters. Two consecutive NUL characters, i.e., a starting point with a Zero-length file name is not allowed and will lead to an error diagnostic followed by a non-Zero exit code later. In the case the given file is empty, find does not process any starting point and therefore will exit immediately after parsing the program arguments. This is unlike the standard invocation where find assumes the current directory as starting point if no path argument is passed. The processing of the starting points is otherwise as usual, e.g. find will recurse into subdirectories unless otherwise prevented. To process only the starting points, one can additionally pass ‘-maxdepth 0’. Further notes: if a file is listed more than once in the input file, it is unspecified whether it is visited more than once. If the file is mutated during the operation of find, the result is unspecified as well. Finally, the seek position within the named ‘file’ at the time find exits, be it with ‘-quit’ or in any other way, is also unspecified. By "unspecified" here is meant that it may or may not work or do any specific thing, and that the behavior may change from platform to platform, or from findutils release to release. Example: Given that another program proggy pre-filters and creates a huge NULseparated list of files, process those as starting points, and find all regular, empty files among them:

     $ proggy | find -files0-from - -maxdepth 0 -type f -empty

The use of ‘-files0-from -’ means to read the names of the starting points from standard input, i.e., from the pipe; and ‘-maxdepth 0’ ensures that only explicitly those entries are examined without recursing into directories (in the case one of the starting points is one).

2.3 Name

Here are ways to search for files whose name matches a certain pattern. See Section 2.3.4 [Shell Pattern Matching], page 9, for a description of the pattern arguments to these tests.

Each of these tests has a case-sensitive version and a case-insensitive version, whose name begins with ‘i’. In a case-insensitive comparison, the patterns ‘fo*’ and ‘F??’ match the file names Foo, ‘FOO’, ‘foo’, ‘fOo’, etc.

2.3.1 Base Name Patterns

[Test] -name pattern

[Test] -iname pattern

True if the base of the file name (the path with the leading directories removed) matches shell pattern pattern. For ‘-iname’, the match is case-insensitive.^{1} To ignore a whole directory tree, use ‘-prune’ (see Section 2.11 [Directories], page 19). As an example, to find Texinfo source files in /usr/local/doc:

     find /usr/local/doc -name '*.texi'

Notice that the wildcard must be enclosed in quotes in order to protect it from expansion by the shell.

As of findutils version 4.2.2, patterns for ‘-name’ and ‘-iname’ match a file name with a leading ‘.’. For example the command ‘find /tmp -name \*bar’ match the file /tmp/.foobar. Braces within the pattern (‘{}’) are not considered to be special (that is, find . -name 'foo{1,2}' matches a file named foo{1,2}, not the files foo1

and foo2.

Because the leading directories are removed, the file names considered for a match with ‘-name’ will never include a slash, so ‘-name a/b’ will never match anything (you probably need to use ‘-path’ instead).

2.3.2 Full Name Patterns

[Test] -path pattern

[Test] -wholename pattern

True if the entire file name, starting with the command line argument under which the file was found, matches shell pattern pattern. To ignore a whole directory tree, use ‘-prune’ rather than checking every file in the tree (see Section 2.11 [Directories], page 19). The “entire file name” as used by find starts with the starting-point specified on the command line, and is not converted to an absolute pathname, so for example cd /; find tmp -wholename /tmp will never match anything.

Find compares the ‘-path’ argument with the concatenation of a directory name and the base name of the file it’s considering. Since the concatenation will never end with a slash, ‘-path’ arguments ending in ‘/’ will match nothing (except perhaps a start point specified on the command line). The name ‘-wholename’ is GNU-specific, but ‘-path’ is more portable; it is supported by HP-UX find and is part of the POSIX 2008 standard.

1 Because we need to perform case-insensitive matching, the GNU fnmatch implementation is always used;

if the C library includes the GNU implementation, we use that and otherwise we use the one from gnulib

[Test] -ipath pattern

[Test] -iwholename pattern

These tests are like ‘-wholename’ and ‘-path’, but the match is case-insensitive.

In the context of the tests ‘-path’, ‘-wholename’, ‘-ipath’ and ‘-iwholename’, a “full path” is the name of all the directories traversed from find’s start point to the file being tested, followed by the base name of the file itself. These paths are often not absolute paths; for example

$ cd /tmp
$ mkdir -p foo/bar/baz
$ find foo -path foo/bar -print
foo/bar
$ find foo -path /tmp/foo/bar -print
$ find /tmp/foo -path /tmp/foo/bar -print
/tmp/foo/bar

Notice that the second find command prints nothing, even though /tmp/foo/bar exists and was examined by find.

Unlike file name expansion on the command line, a ‘*’ in the pattern will match both ‘/’ and leading dots in file names:

$ find .  -path '*f'
./quux/bar/baz/f
$ find .  -path '*/*config'
./quux/bar/baz/.config

[Test] -regex expr

[Test] -iregex expr

True if the entire file name matches regular expression expr. This is a match on the whole path, not a search. For example, to match a file named ./fubar3, you can use the regular expression ‘.*bar.’ or ‘.*b.*3’, but not ‘f.*r3’. For ‘-iregex’, the match is case-insensitive. As for ‘-path’, the candidate file name never ends with a slash, so regular expressions which only match something that ends in slash will always fail. There are several varieties of regular expressions; by default this test uses GNU Emacs regular expressions, but this can be changed with the option ‘-regextype’.

[Option] -regextype name

This option controls the variety of regular expression syntax understood by the ‘-regex’ and ‘-iregex’ tests. This option is positional; that is, it only affects regular expressions which occur later in the command line. If this option is not given, GNU Emacs regular expressions are assumed. Currently-implemented types are

‘emacs’ Regular expressions compatible with GNU Emacs; this is also the default behaviour if this option is not used.

     ‘posix-awk’

Regular expressions compatible with the POSIX awk command (not GNU awk)

     ‘posix-basic’

POSIX Basic Regular Expressions.

     ‘posix-egrep’

Regular expressions compatible with the POSIX egrep command

     ‘posix-extended’

POSIX Extended Regular Expressions

Section 8.5 [Regular Expressions], page 70, for more information on the regular expression dialects understood by GNU findutils.

2.3.3 Fast Full Name Search

To search for files by name without having to actually scan the directories on the disk (which can be slow), you can use the locate program. For each shell pattern you give it, locate searches one or more databases of file names and displays the file names that contain the pattern. See Section 2.3.4 [Shell Pattern Matching], page 9, for details about shell patterns.

If a pattern is a plain string – it contains no metacharacters – locate displays all file names in the database that contain that string. If a pattern contains metacharacters, locate only displays file names that match the pattern exactly. As a result, patterns that contain metacharacters should usually begin with a ‘*’, and will most often end with one as well. The exceptions are patterns that are intended to explicitly match the beginning or end of a file name.

If you only want locate to match against the last component of the file names (the “base name” of the files) you can use the ‘--basename’ option. The opposite behaviour is the default, but can be selected explicitly by using the option ‘--wholename’.

The command

     locate pattern

is almost equivalent to

     find directories -name pattern

where directories are the directories for which the file name databases contain information. The differences are that the locate information might be out of date, and that by default locate matches wildcards against the whole file name (not just its base name) (see Section 2.3.4 [Shell Pattern Matching], page 9).

The file name databases contain lists of files that were on the system when the databases were last updated. The system administrator can choose the file name of the default database, the frequency with which the databases are updated, and the directories for which they contain entries.

Here is how to select which file name databases locate searches. The default is system-dependent. At the time this document was generated, the default was

/usr/local/var/locatedb.
--database=path

-d path Instead of searching the default file name database, search the file name databases in path, which is a colon-separated list of database file names. You can also use the environment variable LOCATE_PATH to set the list of database files to search. The option overrides the environment variable if both are used.

GNU locate can read file name databases generated by the slocate package. However, these generally contain a list of all the files on the system, and so when using this database, locate will produce output only for files which are accessible to you. See Section 8.2 [Invoking locate], page 63, for a description of the ‘--existing’ option which is used to do this.

The updatedb program can also generate database in a format compatible with slocate. See Section 8.3 [Invoking updatedb], page 65, for a description of its ‘--dbformat’ and ‘--output’ options.

2.3.4 Shell Pattern Matching

find and locate can compare file names, or parts of file names, to shell patterns. A shell pattern is a string that may contain the following special characters, which are known as wildcards or metacharacters.

You must quote patterns that contain metacharacters to prevent the shell from expanding them itself. Double and single quotes both work; so does escaping with a backslash.

* Matches any zero or more characters.

? Matches any one character.

[string] Matches exactly one character that is a member of the string string. This is called a character class. As a shorthand, string may contain ranges, which consist of two characters with a dash between them. For example, the class ‘[a-z0-9_]’ matches a lowercase letter, a number, or an underscore. You can negate a class by placing a ‘!’ or ‘^’ immediately after the opening bracket. Thus, ‘[^A-Z@]’ matches any character except an uppercase letter or an at sign.

\ Removes the special meaning of the character that follows it. This works even in character classes.

In the find tests that do shell pattern matching (‘-name’, ‘-wholename’, etc.), wildcards in the pattern will match a ‘.’ at the beginning of a file name. This is also the case for locate. Thus, ‘find -name '*macs'’ will match a file named .emacs, as will ‘locate

'*macs'’.

Slash characters have no special significance in the shell pattern matching that find and locate do, unlike in the shell, in which wildcards do not match them. Therefore, a pattern ‘foo*bar’ can match a file name ‘foo3/bar’, and a pattern ‘./sr*sc’ can match a file name

‘./src/misc’.

If you want to locate some files with the ‘locate’ command but don’t need to see the full list you can use the ‘--limit’ option to see just a small number of results, or the ‘--count’ option to display only the total number of matches.

2.4 Links

There are two ways that files can be linked together. Symbolic links are a special type of file whose contents are a portion of the name of another file. Hard links are multiple directory entries for one file; the file names all have the same index node (inode) number on the disk.

2.4.1 Symbolic Links

Symbolic links are names that reference other files. GNU find will handle symbolic links in one of two ways; firstly, it can dereference the links for you - this means that if it comes across a symbolic link, it examines the file that the link points to, in order to see if it matches the criteria you have specified. Secondly, it can check the link itself in case you might be looking for the actual link. If the file that the symbolic link points to is also within the directory hierarchy you are searching with the find command, you may not see a great deal of difference between these two alternatives.

By default, find examines symbolic links themselves when it finds them (and, if it later comes across the linked-to file, it will examine that, too). If you would prefer find to dereference the links and examine the file that each link points to, specify the ‘-L’ option to find. You can explicitly specify the default behaviour by using the ‘-P’ option. The ‘-H’ option is a half-way-between option which ensures that any symbolic links listed on the command line are dereferenced, but other symbolic links are not.

Symbolic links are different from “hard links” in the sense that you need permission to search the directories in the linked-to file name to dereference the link. This can mean that even if you specify the ‘-L’ option, find may not be able to determine the properties of the file that the link points to (because you don’t have sufficient permission). In this situation, find uses the properties of the link itself. This also occurs if a symbolic link exists but points to a file that is missing, or where the symbolic link points to itself (directly or indirectly).

The options controlling the behaviour of find with respect to links are as follows:

‘-P’ find does not dereference symbolic links at all. This is the default behaviour. This option must be specified before any of the file names on the command line.

‘-H’ find does not dereference symbolic links (except in the case of file names on the command line, which are dereferenced). If a symbolic link cannot be dereferenced, the information for the symbolic link itself is used. This option must be specified before any of the file names on the command line.

‘-L’ find dereferences symbolic links where possible, and where this is not possible it uses the properties of the symbolic link itself. This option must be specified before any of the file names on the command line. Use of this option also implies the same behaviour as the ‘-noleaf’ option. If you later use the ‘-H’ or ‘-P’ options, this does not turn off ‘-noleaf’. Actions that can cause symbolic links to become broken while ‘find’ is executing (for example ‘-delete’) can give rise to confusing behaviour. Take for example the command line ‘find -L . -type d -delete’. This will delete empty directories. If a subtree includes only directories and symbolic links to directoires, this command may still not successfully delete it, since deletion of the target of the symbolic link will cause the symbolic link to become broken and ‘-type d’ is false for broken symbolic links.

‘-follow’ This option forms part of the “expression” and must be specified after the file names, but it is otherwise equivalent to ‘-L’. The ‘-follow’ option affects only those tests which appear after it on the command line. This option is deprecated. Where possible, you should use ‘-L’ instead.

The following differences in behaviour occur when the ‘-L’ option is used:

• find follows symbolic links to directories when searching directory trees.

• ‘-lname’ and ‘-ilname’ always return false (unless they happen to match broken symbolic links).

• ‘-type’ reports the types of the files that symbolic links point to. This means that in combination with ‘-L’, ‘-type l’ will be true only for broken symbolic links. To check for symbolic links when ‘-L’ has been specified, use ‘-xtype l’.

• Implies ‘-noleaf’ (see Section 2.11 [Directories], page 19).

If the ‘-L’ option or the ‘-H’ option is used, the file names used as arguments to ‘-newer’, ‘-anewer’, and ‘-cnewer’ are dereferenced and the timestamp from the pointed-to file is used instead (if possible – otherwise the timestamp from the symbolic link is used).

[Test] -lname pattern

[Test] -ilname pattern

True if the file is a symbolic link whose contents match shell pattern pattern. For ‘-ilname’, the match is case-insensitive. See Section 2.3.4 [Shell Pattern Matching], page 9, for details about the pattern argument. If the ‘-L’ option is in effect, this test will always return false for symbolic links unless they are broken. So, to list any symbolic links to sysdep.c in the current directory and its subdirectories, you can do:

     find . -lname '*sysdep.c'

2.4.2 Hard Links

Hard links allow more than one name to refer to the same file on a file system, i.e., to the same inode. To find all the names which refer to the same file as name, use ‘-samefile

name’.
[Test] -samefile name

True if the file is a hard link to the same inode as name. This implies that name and the file reside on the same file system, i.e., they have the same device number. Unless the ‘-L’ option is also given to follow symbolic links, one may confine the search to one file system by using the ‘-xdev’ option. This is useful because hard links cannot point outside a single file system, so this can cut down on needless searching. If the ‘-L’ option is in effect, then dereferencing of symbolic links applies both to the name argument of the ‘-samefile’ primary and to each file examined during the traversal of the directory hierarchy. Therefore, ‘find -L -samefile name’ will find both hard links and symbolic links pointing to the file referenced by name.

find also allows searching for files by inode number. This can occasionally be useful in diagnosing problems with file systems; for example,

fsck and lsof tend to print inode numbers. Inode numbers also occasionally turn up in log messages for some types of software.

You can learn a file’s inode number and the number of links to it by running ‘ls -li’,

‘stat’ or ‘find -ls’.

You can search for hard links to inode number NUM by using ‘-inum NUM’. If there are any file system mount points below the directory where you are starting the search, use the ‘-xdev’ option unless you are also using the ‘-L’ option. Using ‘-xdev’ saves needless searching, since hard links to a file must be on the same file system. See Section 2.12 [Filesystems], page 21.

[Test] -inum n

True if the file has inode number n. The ‘+’ and ‘-’ qualifiers also work, though these are rarely useful.

Please note that the ‘-inum’ primary simply compares the inode number against the given n. This means that a search for a certain inode number in several file systems may return several files with that inode number, but as each file system has its own device number, those files are not necessarily hard links to the same file.

Therefore, it is much of the time easier to use ‘-samefile’ rather than this option.

find also allows searching for files that have a certain number of links, with ‘-links’.

A directory normally has at least two hard links: the entry named in its parent directory, and the . entry inside of the directory. If a directory has subdirectories, each of those also has a hard link called .. to its parent directory.

The . and .. directory entries are not normally searched unless they are mentioned on the find command line.

[Test] -links n

File has n hard links.

[Test] -links +n

File has more than n hard links.

[Test] -links -n

File has fewer than n hard links.

2.5 Time

Each file has three timestamps, which record the last time that certain operations were performed on the file:

1. access (read the file’s contents)

2. change the status (modify the file or its attributes)

3. modify (change the file’s contents)

Some systems also provide a timestamp that indicates when a file was created. For example, the UFS2 filesystem under NetBSD-3.1 records the birth time of each file. This information is also available under other versions of BSD and some versions of Cygwin. However, even on systems which support file birth time, files may exist for which this information was not recorded (for example, UFS1 file systems simply do not contain this information).

You can search for files whose timestamps are within a certain age range, or compare them to other timestamps.

2.5.1 Age Ranges

These tests are mainly useful with ranges (‘+n’ and ‘-n’).

[Test] -atime n

[Test] -ctime n

[Test] -mtime n

True if the file was last accessed (or its status changed, or it was modified) n*24 hours ago. The number of 24-hour periods since the file’s timestamp is always rounded down; therefore 0 means “less than 24 hours ago”, 1 means “between 24 and 48 hours ago”, and so forth. Fractional values are supported but this only really makes sense for the case where ranges (‘+n’ and ‘-n’) are used.

[Test] -amin n

[Test] -cmin n

[Test] -mmin n

True if the file was last accessed (or its status changed, or it was modified) n minutes ago. These tests provide finer granularity of measurement than ‘-atime’ et al., but rounding is done in a similar way (again, fractions are supported). For example, to list files in /u/bill that were last read from 2 to 6 minutes ago:

     find /u/bill -amin +2 -amin -6
[Option] -daystart

Measure times from the beginning of today rather than from 24 hours ago. So, to list the regular files in your home directory that were modified yesterday, do

     find ~/ -daystart -type f -mtime 1

The ‘-daystart’ option is unlike most other options in that it has an effect on the way that other tests are performed. The affected tests are ‘-amin’, ‘-cmin’, ‘-mmin’, ‘-atime’, ‘-ctime’ and ‘-mtime’. The ‘-daystart’ option only affects the behaviour of any tests which appear after it on the command line.

2.5.2 Comparing Timestamps

[Test] -newerXY reference

Succeeds if timestamp ‘X’ of the file being considered is newer than timestamp ‘Y’ of the file reference. The letters ‘X’ and ‘Y’ can be any of the following letters:

‘a’ Last-access time of reference

‘B’ Birth time of reference (when this is not known, the test cannot succeed)

‘c’ Last-change time of reference

‘m’ Last-modification time of reference

‘t’ The reference argument is interpreted as a literal time, rather than the name of a file. See Chapter 6 [Date input formats], page 51, for a description of how the timestamp is understood. Tests of the form ‘-newerXt’ are valid but tests of the form ‘-newertY’ are not.

For example the test -newerac /tmp/foo succeeds for all files which have been accessed more recently than /tmp/foo was changed. Here ‘X’ is ‘a’ and ‘Y’ is ‘c’.

Not all files have a known birth time. If ‘Y’ is ‘B’ and the birth time of reference is not available, find exits with an explanatory error message. If ‘X’ is ‘B’ and we do not know the birth time the file currently being considered, the test simply fails (that is, it behaves like -false does). Some operating systems (for example, most implementations of Unix) do not support file birth times. Some others, for example NetBSD-3.1, do. Even on operating systems which support file birth times, the information may not be available for specific files. For example, under NetBSD, file birth times are supported on UFS2 file systems, but not UFS1 file systems.

There are two ways to list files in /usr modified after February 1 of the current year. One uses ‘-newermt’:

find /usr -newermt "Feb 1"

The other way of doing this works on the versions of find before 4.3.3:

touch -t 02010000 /tmp/stamp$$
find /usr -newer /tmp/stamp$$
rm -f /tmp/stamp$$

[Test] -anewer reference

[Test] -cnewer reference

[Test] -newer reference

True if the time of the last access (or status change or data modification) of the current file is more recent than that of the last data modification of the reference file. As such, ‘-anewer’ is equivalent to ‘-neweram’, ‘-cnewer’ to ‘-newercm’, and

‘-newer’ to ‘-newermm’.

If reference is a symbolic link and the ‘-H’ option or the ‘-L’ option is in effect, then the time of the last data modification of the file it points to is always used. These tests are affected by ‘-follow’ only if ‘-follow’ comes before them on the command line. See Section 2.4.1 [Symbolic Links], page 10, for more information on

‘-follow’.

As an example, to list any files modified since /bin/sh was last modified:

     find . -newer /bin/sh

[Test] -used n

True if the file was last accessed n days after its status was last changed. Useful for finding files that are not being used, and could perhaps be archived or removed to save disk space.

2.6 Size

[Test] -size n[bckwMG]

True if the file uses n units of space, rounding up. The units are 512-byte blocks by default, but they can be changed by adding a one-character suffix to n:

b 512-byte blocks (never 1024)

c bytes w 2-byte words

k Kibibytes (KiB, units of 1024 bytes)

M Mebibytes (MiB, units of 1024 * 1024 = 1048576 bytes)

G Gibibytes (GiB, units of 1024 * 1024 * 1024 = 1073741824 bytes)

The ‘b’ suffix always considers blocks to be 512 bytes. This is not affected by the setting (or non-setting) of the POSIXLY_CORRECT environment variable. This behaviour is different from the behaviour of the ‘-ls’ action). If you want to use 1024-byte units, use the ‘k’ suffix instead. The number can be prefixed with a ‘+’ or a ‘-’. A plus sign indicates that the test should succeed if the file uses at least n units of storage (a common use of this test) and a minus sign indicates that the test should succeed if the file uses less than n units of storage; i.e., an exact size of n units does not match. Bear in mind that the size is rounded up to the next unit. Therefore ‘-size -1M’ is not equivalent to ‘-size -1048576c’. The former only matches empty files, the latter matches files from 0 to 1,048,575 bytes. There is no ‘=’ prefix, because that’s the default anyway. The size is simply the st size member of the struct stat populated by the lstat (or stat) system call, rounded up as shown above. In other words, it’s consistent with the result you get for ‘ls -l’. This handling of sparse files differs from the output of the ‘%k’ and ‘%b’ format specifiers for the ‘-printf’ predicate.

[Test] -empty

True if the file is empty and is either a regular file or a directory. This might help determine good candidates for deletion. This test is useful with ‘-depth’ (see Section 2.11 [Directories], page 19) and ‘-delete’ (see Section 3.3.1 [Single File], page 30).

2.7 Type

[Test] -type c

True if the file is of type c:

b block (buffered) special

c character (unbuffered) special

d directory

p named pipe (FIFO)

f regular file

l symbolic link; if ‘-L’ is in effect, this is true only for broken symbolic links. If you want to search for symbolic links when ‘-L’ is in effect, use ‘-xtype’ instead of ‘-type’.

s socket

D door (Solaris)

As a GNU extension, multiple file types can be provided as a combined list separated by comma ‘,’. For example, ‘-type f,d,l’ is logically interpreted as ‘( -type f -o

-type d -o -type l )’.

[Test] -xtype c

This test behaves the same as ‘-type’ unless the file is a symbolic link. If the file is a symbolic link, the result is as follows (in the table below, ‘X’ should be understood to represent any letter except ‘l’):

‘`-P -xtype l'’

True if the symbolic link is broken or has an infinite loop

‘`-P -xtype X'’

True if the (ultimate) target file is of type ‘X’.

‘`-L -xtype l'’

Always true

‘`-L -xtype X'’

False unless the symbolic link is broken or has an infinite loop

In other words, for non-broken symbolic links, ‘-xtype’ checks the type of the file that ‘-type’ does not check. For broken symbolic links (or loops), ‘-xtype’ behaves like ‘-type’ does. Symbolic links pointing to things the user has no access to are not considered to be broken.

The ‘-H’ option also affects the behaviour of ‘-xtype’. When ‘-H’ is in effect, ‘-xtype’ behaves as if ‘-L’ had been specified when examining files listed on the command line, and as if ‘-P’ had been specified otherwise. If neither ‘-H’ nor ‘-L’ was specified, ‘-xtype’ behaves as if ‘-P’ had been specified.

See Section 2.4.1 [Symbolic Links], page 10, for more information on ‘-follow’ and ‘-L’.

2.8 Owner

[Test] -user uname

[Test] -group gname

True if the file is owned by user uname (belongs to group gname). A numeric ID is allowed.

[Test] -uid n

[Test] -gid n

True if the file’s numeric user ID (group ID) is n. These tests support ranges (‘+n’ and ‘-n’), unlike ‘-user’ and ‘-group’.

[Test] -nouser
[Test] -nogroup

True if no user corresponds to the file’s numeric user ID (no group corresponds to the numeric group ID). These cases usually mean that the files belonged to users who have since been removed from the system. You probably should change the ownership of such files to an existing user or group, using the chown or chgrp program.

2.9 File Mode Bits

See Chapter 5 [File Permissions], page 45, for information on how file mode bits are structured and how to specify them.

Four tests determine what users can do with files. These are ‘-readable’, ‘-writable’, ‘-executable’ and ‘-perm’. The first three tests ask the operating system if the current user can perform the relevant operation on a file, while ‘-perm’ just examines the file’s mode. The file mode may give a misleading impression of what the user can actually do, because the file may have an access control list, or exist on a read-only filesystem, for example. Of these four tests though, only ‘-perm’ is specified by the POSIX standard.

The ‘-readable’, ‘-writable’ and ‘-executable’ tests are implemented via the access system call. This is implemented within the operating system itself. If the file being considered is on an NFS filesystem, the remote system may allow or forbid read or write operations for reasons of which the NFS client cannot take account. This includes user-ID mapping, either in the general sense or the more restricted sense in which remote superusers are treated by the NFS server as if they are the local user ‘nobody’ on the NFS server.

None of the tests in this section should be used to verify that a user is authorised to perform any operation (on the file being tested or any other file) because of the possibility of a race condition. That is, the situation may change between the test and an action being taken on the basis of the result of that test.

[Test] -readable

True if the file can be read by the invoking user.

[Test] -writable

True if the file can be written by the invoking user. This is an in-principle check, and other things may prevent a successful write operation; for example, the filesystem might be full.

[Test] -executable

True if the file can be executed/searched by the invoking user.

[Test] -perm pmode

True if the file’s mode bits match pmode, which can be either a symbolic or numeric mode (see Chapter 5 [File Permissions], page 45) optionally prefixed by ‘-’ or ‘/’. Note that pmode starts with all file mode bits cleared, i.e., does not relate to the process’s file creation bit mask (also known as umask). A pmode that starts with neither ‘-’ nor ‘/’ matches if mode exactly matches the file mode bits. (To avoid confusion with an obsolete GNU extension, mode must not start with a ‘+’ immediately followed by an octal digit.) A pmode that starts with ‘-’ matches if all the file mode bits set in mode are set for the file; bits not set in mode are ignored. A pmode that starts with ‘/’ matches if any of the file mode bits set in mode are set for the file; bits not set in mode are ignored. This is a GNU extension. If you don’t use the ‘/’ or ‘-’ form with a symbolic mode string, you may have to specify a rather complex mode string. For example ‘-perm g=w’ will only match files that have mode 0020 (that is, ones for which group write permission is the only file mode bit set). It is more likely that you will want to use the ‘/’ or ‘-’ forms, for example ‘-perm -g=w’, which matches any file with group write permission.

‘-perm 664’

Match files that have read and write permission for their owner, and group, but that the rest of the world can read but not write to. Do not match files that meet these criteria but have other file mode bits set (for example if someone can execute/search the file).

‘-perm -664’

Match files that have read and write permission for their owner, and group, but that the rest of the world can read but not write to, without regard to the presence of any extra file mode bits (for example the executable bit). This matches a file with mode 0777, for example.

‘-perm /222’

Match files that are writable by somebody (their owner, or their group, or anybody else).

‘-perm /022’

Match files that are writable by their group or everyone else - the latter often called other. The files don’t have to be writable by both the group and other to be matched; either will do.

‘-perm /g+w,o+w’

As above.

‘-perm /g=w,o=w’

As above.

‘-perm -022’

Match files that are writable by both their group and everyone else.

‘-perm -g+w,o+w’

As above.

‘-perm -444 -perm /222 ! -perm /111’

Match files that are readable for everybody, have at least one write bit set (i.e., somebody can write to them), but that cannot be executed/searched by anybody. Note that in some shells the ‘!’ must be escaped.

‘-perm -a+r -perm /a+w ! -perm /a+x’

As above.

Warning: If you specify ‘-perm /000’ or ‘-perm /mode’ where the symbolic mode ‘mode’ has no bits set, the test matches all files. Versions of GNU find prior to 4.3.3 matched no files in this situation.

[Test] -context pattern

True if file’s SELinux context matches the pattern pattern. The pattern uses shell glob matching. This predicate is supported only on find versions compiled with SELinux support and only when SELinux is enabled.

2.10 Contents

To search for files based on their contents, you can use the grep program. For example, to find out which C source files in the current directory contain the string ‘thing’, you can do:

     grep -l thing *.[ch]

If you also want to search for the string in files in subdirectories, you can combine grep with find and xargs, like this:

     find . -name '*.[ch]' | xargs grep -l thing

The ‘-l’ option causes grep to print only the names of files that contain the string, rather than the lines that contain it. The string argument (‘thing’) is actually a regular expression, so it can contain metacharacters. This method can be refined a little by using the ‘-r’ option to make xargs not run grep if find produces no output, and using the find action ‘-print0’ and the xargs option ‘-0’ to avoid misinterpreting files whose names contain spaces:

     find . -name '*.[ch]' -print0 | xargs -r -0 grep -l thing

For a fuller treatment of finding files whose contents match a pattern, see the manual page for grep.

2.11 Directories

Here is how to control which directories find searches, and how it searches them. These two options allow you to process a horizontal slice of a directory tree.

[Option] -maxdepth levels

Descend at most levels (a non-negative integer) levels of directories below the command line arguments. Using ‘-maxdepth 0’ means only apply the tests and actions to the command line arguments.

          $ mkdir -p dir/d1/d2/d3/d4/d5/d6
          $ find dir -maxdepth 1
          dir
          dir/d1
          $ find dir -mindepth 5
          dir/d1/d2/d3/d4/d5
          dir/d1/d2/d3/d4/d5/d6
          $ find dir -mindepth 2 -maxdepth 4
          dir/d1/d2
          dir/d1/d2/d3
          dir/d1/d2/d3/d4

[Option] -mindepth levels

Do not apply any tests or actions at levels less than levels (a non-negative integer). Using ‘-mindepth 1’ means process all files except the command line arguments. See ‘-maxdepth’ for examples.

[Option] -depth

Process each directory’s contents before the directory itself. Doing this is a good idea when producing lists of files to archive with cpio or tar. If a directory does not have write permission for its owner, its contents can still be restored from the archive since the directory’s permissions are restored after its contents.

[Option] -d

This is a deprecated synonym for ‘-depth’, for compatibility with Mac OS X, FreeBSD and OpenBSD. The ‘-depth’ option is a POSIX feature, so it is better to use that.

[Action] -prune

If the file is a directory, do not descend into it. The result is true. For example, to skip the directory src/emacs and all files and directories under it, and print the names of the other files found:

     find . -wholename './src/emacs' -prune -o -print

The above command will not print ./src/emacs among its list of results. This however is not due to the effect of the ‘-prune’ action (which only prevents further descent, it doesn’t make sure we ignore that item). Instead, this effect is due to the use of ‘-o’. Since the left hand side of the “or” condition has succeeded for ./src/emacs, it is not necessary to evaluate the right-hand-side (‘-print’) at all for this particular file. If you wanted to print that directory name you could use either an extra ‘-print’ action:

     find . -wholename './src/emacs' -prune -print -o -print

or use the comma operator:

     find . -wholename './src/emacs' -prune , -print

If the ‘-depth’ option is in effect, the subdirectories will have already been visited in any case. Hence ‘-prune’ has no effect in this case. Because ‘-delete’ implies ‘-depth’, using ‘-prune’ in combination with ‘-delete’ may well result in the deletion of more files than you intended.

[Action] -quit

Exit immediately (with return value zero if no errors have occurred). This is different to ‘-prune’ because ‘-prune’ only applies to the contents of pruned directories, while ‘-quit’ simply makes find stop immediately. No child processes will be left running. Any command lines which have been built by ‘-exec ... \+’ or ‘-execdir ... \+’ are invoked before the program is exited. After ‘-quit’ is executed, no more files specified on the command line will be processed. For example, ‘find /tmp/foo /tmp/bar -print -quit’ will print only ‘/tmp/foo’. One common use of ‘-quit’ is to stop searching the file system once we have found what we want. For example, if we want to find just a single file we can do this:

     find / -name needle -print -quit

[Option] -noleaf

Do not optimize by assuming that directories contain 2 fewer subdirectories than their hard link count. This option is needed when searching filesystems that do not follow the Unix directory-link convention, such as CD-ROM or MS-DOS filesystems or AFS volume mount points. Each directory on a normal Unix filesystem has at least 2 hard links: its name and its . entry. Additionally, its subdirectories (if any) each have a .. entry linked to that directory. When find is examining a directory, after it has statted 2 fewer subdirectories than the directory’s link count, it knows that the rest of the entries in the directory are non-directories (leaf files in the directory tree). If only the files’ names need to be examined, there is no need to stat them; this gives a significant increase in search speed.

[Option] -ignore_readdir_race

If a file disappears after its name has been read from a directory but before find gets around to examining the file with stat, don’t issue an error message. If you don’t specify this option, an error message will be issued. Furthermore, find with the ‘-ignore_readdir_race’ option will ignore errors of the ‘-delete’ action in the case the file has disappeared since the parent directory was read: it will not output an error diagnostic, and the return code of the ‘-delete’ action will be true. This option can be useful in system scripts (cron scripts, for example) that examine areas of the filesystem that change frequently (mail queues, temporary directories, and so forth), because this scenario is common for those sorts of directories. Completely silencing error messages from find is undesirable, so this option neatly solves the problem. There is no way to search one part of the filesystem with this option on and part of it with this option off, though. When this option is turned on and find discovers that one of the start-point files specified on the command line does not exist, no error message will be issued.

[Option] -noignore_readdir_race

This option reverses the effect of the ‘-ignore_readdir_race’ option.

2.12 Filesystems

A filesystem is a section of a disk, either on the local host or mounted from a remote host over a network. Searching network filesystems can be slow, so it is common to make find avoid them.

There are two ways to avoid searching certain filesystems. One way is to tell find to only search one filesystem:

[Option] -xdev

[Option] -mount

Don’t descend directories on other filesystems. These options are synonyms.

The other way is to check the type of filesystem each file is on, and not descend directories that are on undesirable filesystem types:

[Test] -fstype type

True if the file is on a filesystem of type type. The valid filesystem types vary among different versions of Unix; an incomplete list of filesystem types that are accepted on some version of Unix or another is:

     autofs ext3 ext4 fuse.sshfs nfs proc sshfs sysfs ufs tmpfs xfs

You can use ‘-printf’ with the ‘%F’ directive to see the types of your filesystems. The ‘%D’ directive shows the device number. See Section 3.2 [Print File Information], page 23. ‘-fstype’ is usually used with ‘-prune’ to avoid searching remote filesystems (see Section 2.11 [Directories], page 19).

2.13 Combining Primaries With Operators

Operators build a complex expression from tests and actions. The operators are, in order of decreasing precedence:

( expr ) Force precedence. True if expr is true.

! expr

-not expr True if expr is false. In some shells, it is necessary to protect the ‘!’ from shell interpretation by quoting it.

expr1 expr2
expr1 -a expr2
expr1 -and expr2

And; expr2 is not evaluated if expr1 is false.

expr1 -o expr2
expr1 -or expr2

Or; expr2 is not evaluated if expr1 is true.

expr1 , expr2

List; both expr1 and expr2 are always evaluated. True if expr2 is true. The value of expr1 is discarded. This operator lets you do multiple independent operations on one traversal, without depending on whether other operations succeeded. The two operations expr1 and expr2 are not always fully independent, since expr1 might have side effects like touching or deleting files, or it might use ‘-prune’ which would also affect expr2.

find searches the directory tree rooted at each file name by evaluating the expression from left to right, according to the rules of precedence, until the outcome is known (the left hand side is false for ‘-and’, true for ‘-or’), at which point find moves on to the next file name.

There are two other tests that can be useful in complex expressions:

[Test] -true

Always true.

[Test] -false

Always false.

3 Actions

There are several ways you can print information about the files that match the criteria you gave in the find expression. You can print the information either to the standard output or to a file that you name. You can also execute commands that have the file names as arguments. You can use those commands as further filters to select files.

3.1 Print File Name

[Action] -print

True; print the entire file name on the standard output, followed by a newline. If there is the faintest possibility that one of the files for which you are searching might contain a newline, you should use ‘-print0’ instead.

[Action] -fprint file

True; print the entire file name into file file, followed by a newline. If file does not exist when find is run, it is created; if it does exist, it is truncated to 0 bytes. The named output file is always created, even if no output is sent to it. The file names /dev/stdout and /dev/stderr are handled specially; they refer to the standard output and standard error output, respectively. If there is the faintest possibility that one of the files for which you are searching might contain a newline, you should use ‘-fprint0’ instead.

3.2 Print File Information

[Action] -ls

True; list the current file in ‘ls -dils’ format on the standard output. The output looks like this:

      204744   17 -rw-r--r--   1 djm      staff       17337 Nov  2  1992 ./lwall-quotes

The fields are:

1. The inode number of the file. See Section 2.4.2 [Hard Links], page 11, for how to find files based on their inode number.

2. the number of blocks in the file. The block counts are of 1K blocks, unless the environment variable POSIXLY_CORRECT is set, in which case 512-byte blocks are used. See Section 2.6 [Size], page 14, for how to find files based on their size.

3. The file’s type and file mode bits. The type is shown as a dash for a regular file; for other file types, a letter like for ‘-type’ is used (see Section 2.7 [Type], page 15). The file mode bits are read, write, and execute/search for the file’s owner, its group, and other users, respectively; a dash means the permission is not granted. See Chapter 5 [File Permissions], page 45, for more details about file permissions. See Section 2.9 [Mode Bits], page 17, for how to find files based on their file mode bits.

4. The number of hard links to the file.

5. The user who owns the file.

6. The file’s group.

7. The file’s size in bytes.

8. The date the file was last modified.

9. The file’s name. ‘-ls’ quotes non-printable characters in the file names using C-like backslash escapes. This may change soon, as the treatment of unprintable characters is harmonised for ‘-ls’, ‘-fls’, ‘-print’, ‘-fprint’, ‘-printf’ and

    ‘-fprintf’.

[Action] -fls file

True; like ‘-ls’ but write to file like ‘-fprint’ (see Section 3.1 [Print File Name], page 23). The named output file is always created, even if no output is sent to it.

[Action] -printf format

True; print format on the standard output, interpreting ‘\’ escapes and ‘%’ directives (more details in the following sections). Field widths and precisions can be specified as with the printf C function. Format flags (like ‘#’ for example) may not work as you expect because many of the fields, even numeric ones, are printed with %s. Numeric flags which are affected in this way include ‘G’, ‘U’, ‘b’, ‘D’, ‘k’ and ‘n’. This difference in behaviour means though that the format flag ‘-’ will work; it forces left-alignment of the field. Unlike ‘-print’, ‘-printf’ does not add a newline at the end of the string. If you want a newline at the end of the string, add a ‘\n’. As an example, an approximate equivalent of ‘-ls’ with null-terminated filenames can be achieved with this -printf format:

     find -printf "%i %4k %M %3n %-8u %-8g %8s %T+ %p\n->%l\0" | cat

A practical reason for doing this would be to get literal filenames in the output, instead of ‘-ls’’s backslash-escaped names. (Using cat here prevents this happening for the ‘%p’ format specifier; see Section 3.3.2.3 [Unusual Characters in File Names], page 33). This format also outputs a uniform timestamp format. As for symbolic links, the format above outputs the target of the symbolic link on a second line, following ‘\n->’. There is nothing following the arrow for file types other than symbolic links. Another approach, for complete consistency, would be to -fprintf the symbolic links into a separate file, so they too can be null-terminated.

[Action] -fprintf file format

True; like ‘-printf’ but write to file like ‘-fprint’ (see Section 3.1 [Print File Name], page 23). The output file is always created, even if no output is ever sent to it.

3.2.1 Escapes

The escapes that ‘-printf’ and ‘-fprintf’ recognise are:

\a Alarm bell.

\b Backspace.

\c Stop printing from this format immediately and flush the output.

\f Form feed.

\n Newline.

\r Carriage return.

\t Horizontal tab.

\v Vertical tab.

\\ A literal backslash (‘\’).

\0 ASCII NUL.

\NNN The character whose ASCII code is NNN (octal).

A ‘\’ character followed by any other character is treated as an ordinary character, so they both are printed, and a warning message is printed to the standard error output (because it was probably a typo).

3.2.2 Format Directives

‘-printf’ and ‘-fprintf’ support the following format directives to print information about the file being processed. The C printf function, field width and precision specifiers are supported, as applied to string (%s) types. That is, you can specify "minimum field width"."maximum field width" for each directive. Format flags (like ‘#’ for example) may not work as you expect because many of the fields, even numeric ones, are printed with %s. The format flag ‘-’ does work; it forces left-alignment of the field.

‘%%’ is a literal percent sign. See Section 3.2.2.7 [Reserved and Unknown Directives], page 28, for a description of how format directives not mentioned below are handled.

A ‘%’ at the end of the format argument causes undefined behaviour since there is no following character. In some locales, it may hide your door keys, while in others it may remove the final page from the novel you are reading.

3.2.2.1 Name Directives

%p File’s name (not the absolute path name, but the name of the file as it was encountered by find - that is, as a relative path from one of the starting points).

%f File’s name with any leading directories removed (only the last element). That is, the basename of the file.

%h Leading directories of file’s name (all but the last element and the slash before it). That is, the dirname of the file. If the file’s name contains no slashes (for example because it was named on the command line and is in the current working directory), then “%h” expands to “.”. This prevents “%h/%f” expanding to “/foo”, which would be surprising and probably not desirable.

%P File’s name with the name of the command line argument under which it was found removed from the beginning.

%H Command line argument under which file was found.

For some corner-cases, the interpretation of the ‘%f’ and ‘%h’ format directives is not obvious. Here is an example including some output:

     $ find \
       . .. / /tmp /tmp/TRACE compile compile/64/tests/find \
       -maxdepth 0 -printf '%p: [%h][%f]\n'

.: [.][.]
..: [.][..]
/: [][/]
/tmp: [][tmp]
/tmp/TRACE: [/tmp][TRACE]
compile: [.][compile]
compile/64/tests/find: [compile/64/tests][find]

3.2.2.2 Ownership Directives

%g File’s group name, or numeric group ID if the group has no name.

%G File’s numeric group ID.

%u File’s user name, or numeric user ID if the user has no name.

%U File’s numeric user ID.

%m File’s mode bits (in octal). If you always want to have a leading zero on the number, use the ’#’ format flag, for example ’%#m’.

The file mode bit numbers used are the traditional Unix numbers, which will be as expected on most systems, but if your system’s file mode bit layout differs from the traditional Unix semantics, you will see a difference between the mode as printed by ‘%m’ and the mode as it appears in struct stat.

%M File’s type and mode bits (in symbolic form, as for ls). This directive is supported in findutils 4.2.5 and later.

3.2.2.3 Size Directives

%k The amount of disk space used for this file in 1K blocks. Since disk space is allocated in multiples of the filesystem block size this is usually greater than %s/1024, but it can also be smaller if the file is a sparse file (that is, it has “holes”).

%b The amount of disk space used for this file in 512-byte blocks. Since disk space is allocated in multiples of the filesystem block size this is usually greater than %s/512, but it can also be smaller if the file is a sparse file (that is, it has “holes”).

%s File’s size in bytes.

%S File’s sparseness. This is calculated as (BLOCKSIZE*st_blocks / st_size). The exact value you will get for an ordinary file of a certain length is systemdependent. However, normally sparse files will have values less than 1.0, and files which use indirect blocks and have few holes may have a value which is greater than 1.0. The value used for BLOCKSIZE is system-dependent, but is usually 512 bytes. If the file size is zero, the value printed is undefined. On systems which lack support for st blocks, a file’s sparseness is assumed to be 1.0.

3.2.2.4 Location Directives

%d File’s depth in the directory tree (depth below a file named on the command line, not depth below the root directory). Files named on the command line have a depth of 0. Subdirectories immediately below them have a depth of 1, and so on.

%D The device number on which the file exists (the st_dev field of struct stat), in decimal.

%F Type of the filesystem the file is on; this value can be used for ‘-fstype’ (see Section 2.11 [Directories], page 19).

%l Object of symbolic link (empty string if file is not a symbolic link).

%i File’s inode number (in decimal).

%n Number of hard links to file.

%y Type of the file as used with ‘-type’. If the file is a symbolic link, ‘l’ will be printed.

%Y Type of the file as used with ‘-type’. If the file is a symbolic link, it is dereferenced. If the file is a broken symbolic link, ‘N’ is printed. When determining the type of the target of a symbolic link, and a loop is encountered, then ‘L’ is printed (e.g. for a symbolic link to itself); ‘?’ is printed for any other error

          (like e.g. ‘permission denied’).

3.2.2.5 Time Directives

Some of these directives use the C ctime function. Its output depends on the current locale, but it typically looks like

     Wed Nov  2 00:42:36 1994

%a File’s last access time in the format returned by the C ctime function.

%Ak File’s last access time in the format specified by k (see Section 3.2.3 [Time Formats], page 28).

%Bk File’s birth time, i.e., its creation time, in the format specified by k (see Section 3.2.3 [Time Formats], page 28). This directive produces an empty string if the underlying operating system or filesystem does not support birth times.

%c File’s last status change time in the format returned by the C ctime function.

%Ck File’s last status change time in the format specified by k (see Section 3.2.3 [Time Formats], page 28).

%t File’s last modification time in the format returned by the C ctime function.

%Tk File’s last modification time in the format specified by k (see Section 3.2.3 [Time Formats], page 28).

3.2.2.6 Other Directives

%Z File’s SELinux context, or empty string if the file has no SELinux context.

3.2.2.7 Reserved and Unknown Directives

The ‘%(’, ‘%{’ and ‘%[’ format directives, with or without field with and precision specifications, are reserved for future use. Don’t use them and don’t rely on current experiment to predict future behaviour. To print ‘(’, simply use ‘(’ rather than ‘%(’. Likewise for ‘{’ and ‘[’.

Similarly, a ‘%’ character followed by any other unrecognised character (i.e., not a known directive or printf field width and precision specifier), is discarded (but the unrecognised character is printed), and a warning message is printed to the standard error output (because it was probably a typo). Don’t rely on this behaviour, because other directives may be added in the future.

3.2.3 Time Formats

Below is an incomplete list of formats for the directives ‘%A’, ‘%B’, ‘%C’, and ‘%T’, which print the file’s timestamps. Please refer to the documentation of strftime for the full list. Some of these formats might not be available on all systems, due to differences in the implementation of the C strftime function.

3.2.3.1 Time Components

The following format directives print single components of the time.

H hour (00..23)

I hour (01..12)

k hour ( 0..23)

l hour ( 1..12)

p locale’s AM or PM

Z time zone (e.g., EDT), or nothing if no time zone is determinable

M minute (00..59)

S second (00..61). There is a fractional part.

@ seconds since Jan. 1, 1970, 00:00 GMT, with fractional part.

The fractional part of the seconds field is of indeterminate length and precision. That is, the length of the fractional part of the seconds field will in general vary between findutils releases and between systems. This means that it is unwise to assume that field has any specific length. The length of this field is not usually a guide to the precision of timestamps in the underlying file system.

3.2.3.2 Date Components

The following format directives print single components of the date.

a locale’s abbreviated weekday name (Sun..Sat)

A locale’s full weekday name, variable length (Sunday..Saturday)

b

h locale’s abbreviated month name (Jan..Dec)

B locale’s full month name, variable length (January..December) m month (01..12) d day of month (01..31) w day of week (0..6) j day of year (001..366) U week number of year with Sunday as first day of week (00..53) W week number of year with Monday as first day of week (00..53) Y year (1970…) y last two digits of year (00..99)

3.2.3.3 Combined Time Formats

The following format directives print combinations of time and date components. r time, 12-hour (hh:mm:ss [AP]M) T time, 24-hour (hh:mm:ss.xxxxxxxxxx) X locale’s time representation (H:M:S). The seconds field includes a fractional

part.

c locale’s date and time in ctime format (Sat Nov 04 12:02:33 EST 1989). This

format does not include any fractional part in the seconds field.

D date (mm/dd/yy) F date (yyyy-mm-dd) x locale’s date representation (mm/dd/yy) + Date and time, separated by ’+’, for example ‘2004-04-28+22:22:05.0000000000’.

The time is given in the current timezone (which may be affected by setting the TZ environment variable). This is a GNU extension. The seconds field includes a fractional part.

3.2.4 Formatting Flags

The ‘%m’ and ‘%d’ directives support the ‘#’, ‘0’ and ‘+’ flags, but the other directives do not, even if they print numbers. Numeric directives that do not support these flags include

‘G’, ‘U’, ‘b’, ‘D’, ‘k’ and ‘n’.

All fields support the format flag ‘-’, which makes fields left-aligned. That is, if the field width is greater than the actual contents of the field, the requisite number of spaces are printed after the field content instead of before it.

3.3 Run Commands

You can use the list of file names created by find or locate as arguments to other commands. In this way you can perform arbitrary actions on the files.

3.3.1 Single File

Here is how to run a command on one file at a time.

[Action] -execdir command ;

Execute command; true if command returns zero. find takes all arguments after ‘-execdir’ to be part of the command until an argument consisting of ‘;’ is reached. It replaces the string ‘{}’ by the current file name being processed everywhere it occurs in the command. Both of these constructions need to be escaped (with a ‘\’) or quoted to protect them from expansion by the shell. The command is executed in the directory which find was searching at the time the action was executed (that is, {} will expand to a file in the local directory).

For example, to compare each C header file in or below the current directory with

the file /tmp/master:
     find . -name '*.h' -execdir diff -u '{}' /tmp/master ';'

If you use ‘-execdir’, you must ensure that the PATH variable contains only absolute directory names. Having an empty element in PATH or explicitly including ‘.’ (or any other non-absolute name) is insecure. GNU find will refuse to run if you use ‘-execdir’ and it thinks your PATH setting is insecure. For example:

‘/bin:/usr/bin:’

Insecure; empty path element (at the end)

‘:/bin:/usr/bin:/usr/local/bin’

Insecure; empty path element (at the start)

‘/bin:/usr/bin::/usr/local/bin’

Insecure; empty path element (two colons in a row)

‘/bin:/usr/bin:.:/usr/local/bin’

Insecure; ‘.’ is a path element (. is not an absolute file name)

‘/bin:/usr/bin:sbin:/usr/local/bin’

Insecure; ‘sbin’ is not an absolute file name

‘/bin:/usr/bin:/sbin:/usr/local/bin’

Secure (if you control the contents of those directories and any access to them)

Another similar option, ‘-exec’ is supported, but is less secure. See Chapter 11 [Security Considerations], page 93, for a discussion of the security problems surrounding ‘-exec’.

[Action] -exec command ;

This insecure variant of the ‘-execdir’ action is specified by POSIX. Like ‘-execdir command ;’ it is true if zero is returned by command. The main difference is that the command is executed in the directory from which find was invoked, meaning that ‘{}’ is expanded to a relative path starting with the name of one of the starting directories, rather than just the basename of the matched file.

While some implementations of find replace the ‘{}’ only where it appears on its own in an argument, GNU find replaces ‘{}’ wherever it appears.

3.3.2 Multiple Files

Sometimes you need to process files one at a time. But usually this is not necessary, and, it is faster to run a command on as many files as possible at a time, rather than once per file. Doing this saves on the time it takes to start up the command each time.

The ‘-execdir’ and ‘-exec’ actions have variants that build command lines containing as many matched files as possible.

[Action] -execdir command {} +

This works as for ‘-execdir command ;’, except that the result is always true, and the ‘{}’ at the end of the command is expanded to a list of names of matching files. This expansion is done in such a way as to avoid exceeding the maximum command line length available on the system. Only one ‘{}’ is allowed within the command, and it must appear at the end, immediately before the ‘+’. A ‘+’ appearing in any position other than immediately after ‘{}’ is not considered to be special (that is, it does not terminate the command).

[Action] -exec command {} +

This insecure variant of the ‘-execdir’ action is specified by POSIX. The main difference is that the command is executed in the directory from which find was invoked, meaning that ‘{}’ is expanded to a relative path starting with the name of one of the starting directories, rather than just the basename of the matched file. The result is always true.

Before find exits, any partially-built command lines are executed. This happens even if the exit was caused by the ‘-quit’ action. However, some types of error (for example not being able to invoke stat() on the current directory) can cause an immediate fatal exit. In this situation, any partially-built command lines will not be invoked (this prevents possible infinite loops).

At first sight, it looks like the list of filenames to be processed can only be at the end of the command line, and that this might be a problem for some commands (cp and rsync for example).

However, there is a slightly obscure but powerful workaround for this problem which takes advantage of the behaviour of sh -c:

     find startpoint -tests ... -exec sh -c 'scp "$@" remote:/dest' sh {} +

In the example above, the filenames we want to work on need to occur on the scp command line before the name of the destination. We use the shell to invoke the command scp "$@" remote:/dest and the shell expands "$@" to the list of filenames we want to process.

Another, but less secure, way to run a command on more than one file at once, is to use the xargs command, which is invoked like this:

     xargs [option...] [command [initial-arguments]]

xargs normally reads arguments from the standard input. These arguments are delimited by blanks (which can be protected with double or single quotes or a backslash) or newlines. It executes the command (the default is echo) one or more times with any initialarguments followed by arguments read from standard input. Blank lines on the standard input are ignored. If the ‘-L’ option is in use, trailing blanks indicate that xargs should consider the following line to be part of this one.

Instead of blank-delimited names, it is safer to use ‘find -print0’ or ‘find -fprint0’ and process the output by giving the ‘-0’ or ‘--null’ option to GNU xargs, GNU tar, GNU cpio, or perl. The locate command also has a ‘-0’ or ‘--null’ option which does the same thing.

You can use shell command substitution (backquotes) to process a list of arguments, like this:

     grep -l sprintf ` find $HOME -name '*.c' -print`

However, that method produces an error if the length of the ‘.c’ file names exceeds the operating system’s command line length limit. xargs avoids that problem by running the command as many times as necessary without exceeding the limit:

     find $HOME -name '*.c' -print | xargs grep -l sprintf

However, if the command needs to have its standard input be a terminal (less, for example), you have to use the shell command substitution method or use either the ‘--arg-file’ option or the ‘--open-tty’ option of xargs.

The xargs command will usually process all of its input, building command lines and executing them. The processing stops earlier and immediately if the tool reads a line containing the end-of-file marker string specified with the ‘--eof’ option, or if one of the launched commands exits with a status of 255. The latter will cause xargs to issue an error message and exit with status 124.

3.3.2.1 Unsafe File Name Handling

Because file names can contain quotes, backslashes, blank characters, and even newlines, it is not safe to process them using xargs in its default mode of operation. But since most files’ names do not contain blanks, this problem occurs only infrequently. If you are only searching through files that you know have safe names, then you need not be concerned about it.

Error messages issued by find and locate quote unusual characters in file names in order to prevent unwanted changes in the terminal’s state.

In many applications, if xargs botches processing a file because its name contains special characters, some data might be lost. The importance of this problem depends on the importance of the data and whether anyone notices the loss soon enough to correct it. However, here is an extreme example of the problems that using blank-delimited names can cause. If the following command is run daily from cron, then any user can remove any file on the system:

     find / -name '#*' -atime +7 -print | xargs rm

For example, you could do something like this:

     eg$ echo > '#
     vmunix'

and then cron would delete /vmunix, if it ran xargs with / as its current directory.

To delete other files, for example /u/joeuser/.plan, you could do this:

     eg$ mkdir '#

     '
     eg$ cd '#
     '
     eg$ mkdir u u/joeuser u/joeuser/.plan'
     '
     eg$ echo > u/joeuser/.plan'
     /#foo'
     eg$ cd ..
     eg$ find . -name '#*' -print | xargs echo
     ./# ./# /u/joeuser/.plan /#foo

3.3.2.2 Safe File Name Handling

Here is how to make find output file names so that they can be used by other programs without being mangled or misinterpreted. You can process file names generated this way by giving the ‘-0’ or ‘--null’ option to GNU xargs, GNU tar, GNU cpio, or perl.

[Action] -print0

True; print the entire file name on the standard output, followed by a null character.

[Action] -fprint0 file

True; like ‘-print0’ but write to file like ‘-fprint’ (see Section 3.1 [Print File Name], page 23). The output file is always created.

As of findutils version 4.2.4, the locate program also has a ‘--null’ option which does the same thing. For similarity with xargs, the short form of the option ‘-0’ can also be used.

If you want to be able to handle file names safely but need to run commands which want to be connected to a terminal on their input, you can use the ‘--open-tty’ option to xargs or the ‘--arg-file’ option to xargs like this:

     find / -name xyzzy -print0 > list
     xargs --null --arg-file=list munge

The example above runs the munge program on all the files named xyzzy that we can find, but munge’s input will still be the terminal (or whatever the shell was using as standard input). If your shell has the “process substitution” feature ‘<(...)’, you can do this in just one step:

     xargs --null --arg-file=<(find / -name xyzzy -print0) munge

3.3.2.3 Unusual Characters in File Names

As discussed above, you often need to be careful about how the names of files are handled by find and other programs. If the output of find is not going to another program but instead is being shown on a terminal, this can still be a problem. For example, some character sequences can reprogram the function keys on some terminals. See Chapter 11 [Security Considerations], page 93, for a discussion of other security problems relating to

find.

Unusual characters are handled differently by various actions, as described below.

‘-print0’
‘-fprint0’

Always print the exact file name, unchanged, even if the output is going to a terminal.

‘-ok’

‘-okdir’ Always print the exact file name, unchanged. This will probably change in a future release.

‘-ls’

‘-fls’ Unusual characters are always escaped. White space, backslash, and double quote characters are printed using C-style escaping (for example ‘\f’, ‘\"’). Other unusual characters are printed using an octal escape. Other printable characters (for ‘-ls’ and ‘-fls’ these are the characters between octal 041 and 0176) are printed as-is.

‘-printf’
‘-fprintf’

If the output is not going to a terminal, it is printed as-is. Otherwise, the result depends on which directive is in use:

%D, %F, %H, %Y, %y These expand to values which are not under control of files’ owners, and so are printed as-is.

%a, %b, %c, %d, %g, %G, %i, %k, %m, %M, %n, %s, %t, %u, %U These have values which are under the control of files’ owners but which cannot be used to send arbitrary data to the terminal, and so these are printed as-is.

%f, %h, %l, %p, %P The output of these directives is quoted if the output is going to a terminal. The setting of the LC_CTYPE environment variable is used to determine which characters need to be quoted. This quoting is performed in the same way as for GNU ls. This is not the same quoting mechanism as the one used for ‘-ls’ and ‘fls’. If you are able to decide what format to use for the output of find then it is normally better to use ‘\0’ as a terminator than to use newline, as file names can contain white space and newline characters.

‘-print’

‘-fprint’ Quoting is handled in the same way as for the ‘%p’ directive of ‘-printf’ and ‘-fprintf’. If you are using find in a script or in a situation where the matched files might have arbitrary names, you should consider using ‘-print0’ instead

of ‘-print’.

The locate program quotes and escapes unusual characters in file names in the same way as find’s ‘-print’ action.

The behaviours described above may change soon, as the treatment of unprintable characters is harmonised for ‘-ls’, ‘-fls’, ‘-print’, ‘-fprint’, ‘-printf’ and ‘-fprintf’.

3.3.2.4 Limiting Command Size

xargs gives you control over how many arguments it passes to the command each time it executes it. By default, it uses up to ARG_MAX - 2k, or 128k, whichever is smaller, characters per command. It uses as many lines and arguments as fit within that limit. The following options modify those values.

--no-run-if-empty

-r If the standard input does not contain any nonblanks, do not run the command. By default, the command is run once even if there is no input. This option is a GNU extension.

--max-lines[=max-lines]
-L max-lines
-l[max-lines]

Use at most max-lines nonblank input lines per command line; max-lines defaults to 1 if omitted; omitting the argument is not allowed in the case of the ‘-L’ option. Trailing blanks cause an input line to be logically continued on the next input line, for the purpose of counting the lines. Implies ‘-x’. The preferred name for this option is ‘-L’ as this is specified by POSIX.

--max-args=max-args
-n max-args

Use at most max-args arguments per command line. Fewer than max-args arguments will be used if the size (see the ‘-s’ option) is exceeded, unless the ‘-x’ option is given, in which case xargs will exit.

--max-chars=max-chars
-s max-chars

Use at most max-chars characters per command line, including the command initial arguments and the terminating nulls at the ends of the argument strings. If you specify a value for this option which is too large or small, a warning message is printed and the appropriate upper or lower limit is used instead. You can use ‘--show-limits’ option to understand the command-line limits applying to xargs and how this is affected by any other options. The POSIX limits shown when you do this have already been adjusted to take into account the size of your environment variables. The largest allowed value is system-dependent, and is calculated as the argument length limit for exec, less the size of your environment, less 2048 bytes of headroom. If this value is more than 128KiB, 128Kib is used as the default value; otherwise, the default value is the maximum.

3.3.2.5 Controlling Parallelism

Normally, xargs runs one command at a time. This is called "serial" execution; the commands happen in a series, one after another. If you’d like xargs to do things in "parallel", you can ask it to do so, either when you invoke it, or later while it is running. Running several commands at one time can make the entire operation go more quickly, if the commands are independent, and if your system has enough resources to handle the load. When parallelism works in your application, xargs provides an easy way to get your work done faster.

--max-procs=max-procs
-P max-procs

Run up to max-procs processes at a time; the default is 1. If max-procs is 0, xargs will run as many processes as possible at a time. Use the ‘-n’, ‘-s’, or ‘-L’ option with ‘-P’; otherwise chances are that the command will be run only once. If a child process exits with status 255, xargs will still wait for all child processes to exit (before version 4.9.0 this might not happen).

If xargs is run without the ‘-P’ option, it will not change the handling of the SIGUSR1 and SIGUSR2 signals. This means they will terminate the xargs program unless those signals were set to be ignored in the parent process of xargs. If you do not want parallel execution but you also do not want these signals to be fatal, you can specify -P 1.

Suppose you have a directory tree of large image files and a makeallsizes script that takes a single file name and creates various sized images from it (thumbnail-sized, webpage-sized, printer-sized, and the original large file). The script is doing enough work that it takes significant time to run, even on a single image. You could run:

     find originals -name '*.jpg' | xargs -L 1 makeallsizes

This will run makeallsizes filename once for each .jpg file in the originals directory. However, if your system has two central processors, this script will only keep one of them busy. Instead, you could probably finish in about half the time by running:

     find originals -name '*.jpg' | xargs -L 1 -P 2 makeallsizes

xargs will run the first two commands in parallel, and then whenever one of them terminates, it will start another one, until the entire job is done.

The same idea can be generalized to as many processors as you have handy. It also generalizes to other resources besides processors. For example, if xargs is running commands that are waiting for a response from a distant network connection, running a few in parallel may reduce the overall latency by overlapping their waiting time.

If you are running commands in parallel, you need to think about how they should arbitrate access to any resources that they share. For example, if more than one of them tries to print to stdout, the output will be produced in an indeterminate order (and very likely mixed up) unless the processes collaborate in some way to prevent this. Using some kind of locking scheme is one way to prevent such problems. In general, using a locking scheme will help ensure correct output but reduce performance. If you don’t want to tolerate the performance difference, simply arrange for each process to produce a separate output file (or otherwise use separate resources).

xargs also allows “turning up” or “turning down” its parallelism in the middle of a run. Suppose you are keeping your four-processor system busy for hours, processing thousands of images using -P 4. Now, in the middle of the run, you or someone else wants you to reduce your load on the system, so that something else will run faster. If you interrupt xargs, your job will be half-done, and it may take significant manual work to resume it only for the remaining images. If you suspend xargs using your shell’s job controls (e.g. control-Z), then it will get no work done while suspended.

Find out the process ID of the xargs process, either from your shell or with the ps command. After you send it the signal SIGUSR2, xargs will run one fewer command in parallel. If you send it the signal SIGUSR1, it will run one more command in parallel. For example:

     shell$ xargs <allimages -L 1 -P 4 makeallsizes &
     [4] 27643
        ... at some later point ...
     shell$ kill -USR2 27643
     shell$ kill -USR2 %4

The first kill command will cause xargs to wait for two commands to terminate before starting the next command (reducing the parallelism from 4 to 3). The second kill will reduce it from 3 to 2. (%4 works in some shells as a shorthand for the process ID of the background job labeled [4].)

Similarly, if you started a long xargs job without parallelism, you can easily switch it to start running two commands in parallel by sending it a SIGUSR1.

xargs will never terminate any existing commands when you ask it to run fewer processes. It merely waits for the excess commands to finish. If you ask it to run more commands, it will start the next one immediately (if it has more work to do). If the degree of parallelism is already 1, sending SIGUSR2 will have no further effect (since --max-procs=0 means that there should be no limit on the number of processes to run).

There is an implementation-defined limit on the number of processes. This limit is shown with xargs --show-limits. The limit is at least 127 on all systems (and on the author’s system it is 2147483647).

If you send several identical signals quickly, the operating system does not guarantee that each of them will be delivered to xargs. This means that you can’t rapidly increase or decrease the parallelism by more than one command at a time. You can avoid this problem by sending a signal, observing the result, then sending the next one; or merely by delaying for a few seconds between signals (unless your system is very heavily loaded).

Whether or not parallel execution will work well for you depends on the nature of the commmand you are running in parallel, on the configuration of the system on which you are running the command, and on the other work being done on the system at the time.

3.3.2.6 Interspersing File Names

xargs can insert the name of the file it is processing between arguments you give for the command. Unless you also give options to limit the command size (see Section 3.3.2.4 [Limiting Command Size], page 35), this mode of operation is equivalent to ‘find -exec’ (see Section 3.3.1 [Single File], page 30).

-I replace-str
--replace[=replace-str]
-i[replace-str]

Replace occurrences of replace-str in the initial arguments with names read from standard input. Also, unquoted blanks do not terminate arguments; instead, the input is split at newlines only. If replace-str is omitted (omitting it is allowed only for ‘-i’ and ‘--replace’), it defaults to ‘{}’ (like for ‘find -exec’). Implies ‘-x’ and ‘-L 1’. The ‘-i’ option is deprecated in favour of the ‘-I’ option.

As an example, to sort each file in the bills directory, leaving the output in that file name with .sorted appended, you could do:

find bills -type f | xargs -I XX sort -o XX.sorted XX

The equivalent command using ‘find -execdir’ is:

find bills -type f -execdir sort -o '{}.sorted' '{}' ';'

When you use the ‘-I’ option, each line read from the input is buffered internally. This means that there is an upper limit on the length of input line that xargs will accept when used with the ‘-I’ option. To work around this limitation, you can use the ‘-s’ option to increase the amount of buffer space that xargs uses, and you can also use an extra invocation of xargs to ensure that very long lines do not occur. For example:

somecommand | xargs -s 50000 echo | xargs -I '{}' -s 100000 rm '{}'

Here, the first invocation of xargs has no input line length limit because it doesn’t use the ‘-I’ option. The second invocation of xargs does have such a limit, but we have ensured that it never encounters a line which is longer than it can handle.

This is not an ideal solution. Instead, the ‘-I’ option should not impose a line length limit (apart from any limit imposed by the operating system) and so one might consider this limitation to be a bug. A better solution would be to allow xargs -I to automatically move to a larger value for the ‘-s’ option when this is needed.

This sort of problem doesn’t occur with the output of find because it emits just one filename per line.

3.3.3 Querying

To ask the user whether to execute a command on a single file, you can use the find primary ‘-okdir’ instead of ‘-execdir’, and the find primary ‘-ok’ instead of ‘-exec’:

[Action] -okdir command ;

Like ‘-execdir’ (see Section 3.3.1 [Single File], page 30), but ask the user first. If the user does not agree to run the command, just return false. Otherwise, run it, with standard input redirected from /dev/null. This action may not be specified together with the ‘-files0-from’ option. The response to the prompt is matched against a pair of regular expressions to determine if it is a yes or no response. These regular expressions are obtained from the system (nl_langinfo items YESEXPR and NOEXPR are used) if the POSIXLY_CORRECT environment variable is set and the system has such patterns available. Otherwise, find’s message translations are used. In either case, the LC_MESSAGES environment variable will determine the regular expressions used to determine if the answer is affirmative or negative. The interpretation of the regular expressions themselves will be affected by the environment variables LC_CTYPE (character classes) and LC_COLLATE (character ranges and equivalence classes).

[Action] -ok command ;

This insecure variant of the ‘-okdir’ action is specified by POSIX. The main difference is that the command is executed in the directory from which find was invoked, meaning that ‘{}’ is expanded to a relative path starting with the name of one of the starting directories, rather than just the basename of the matched file. If the command is run, its standard input is redirected from /dev/null. This action may not be specified together with the ‘-files0-from’ option.

When processing multiple files with a single command, to query the user you give xargs the following option. When using this option, you might find it useful to control the number of files processed per invocation of the command (see Section 3.3.2.4 [Limiting Command Size], page 35).

--interactive

-p Prompt the user about whether to run each command line and read a line from the terminal. Only run the command line if the response starts with ‘y’ or ‘Y’. Implies ‘-t’.

3.4 Delete Files

[Action] -delete

Delete files or directories; true if removal succeeded. If the removal failed, an error message is issued and find’s exit status will be nonzero (when it eventually exits).

Warning: Don’t forget that find evaluates the command line as an expression, so putting ‘-delete’ first will make find try to delete everything below the starting points you specified.

The use of the ‘-delete’ action on the command line automatically turns on the ‘-depth’ option. As in turn ‘-depth’ makes ‘-prune’ ineffective, the ‘-delete’ action cannot usefully be combined with ‘-prune’. Often, the user might want to test a find command line with ‘-print’ prior to adding ‘-delete’ for the actual removal run. To avoid surprising results, it is usually best to remember to use ‘-depth’ explicitly during those earlier test runs. See Section 9.3 [Cleaning Up], page 82, for a deeper discussion about good use cases of the ‘-delete’ action and those with surprising results. The ‘-delete’ action will fail to remove a directory unless it is empty. Together with the ‘-ignore_readdir_race’ option, find will ignore errors of the ‘-delete’ action in the case the file has disappeared since the parent directory was read: it will not output an error diagnostic, not change the exit code to nonzero, and the return code of the ‘-delete’ action will be true.

3.5 Adding Tests

You can test for file attributes that none of the find builtin tests check. To do this, use xargs to run a program that filters a list of files printed by find. If possible, use find builtin tests to pare down the list, so the program run by xargs has less work to do. The tests builtin to find will likely run faster than tests that other programs perform.

For reasons of efficiency it is often useful to limit the number of times an external program has to be run. For this reason, it is often a good idea to implement “extended” tests by using xargs.

For example, here is a way to print the names of all of the unstripped binaries in the /usr/local directory tree. Builtin tests avoid running file on files that are not regular files or are not executable.

find /usr/local -type f -perm /a=x | xargs file |
  grep 'not stripped' | cut -d: -f1

The cut program removes everything after the file name from the output of file.

However, using xargs can present important security problems (see Chapter 11 [Security Considerations], page 93). These can be avoided by using ‘-execdir’. The ‘-execdir’ action is also a useful way of putting your own test in the middle of a set of other tests or actions for find (for example, you might want to use ‘-prune’).

To place a special test somewhere in the middle of a find expression, you can use ‘-execdir’ (or, less securely, ‘-exec’) to run a program that performs the test. Because ‘-execdir’ evaluates to the exit status of the executed program, you can use a program (which can be a shell script) that tests for a special attribute and make it exit with a true (zero) or false (non-zero) status. It is a good idea to place such a special test after the builtin tests, because it starts a new process which could be avoided if a builtin test evaluates to false.

Here is a shell script called unstripped that checks whether its argument is an unstripped binary file:

     #! /bin/sh
     file "$1" | grep "not stripped" >/dev/null

This script relies on the shell exiting with the status of the last command in the pipeline, in this case grep. The grep command exits with a true status if it found any matches, false if not. Here is an example of using the script (assuming it is in your search path). It lists the stripped executables (and shell scripts) in the file sbins and the unstripped ones in

ubins.
     find /usr/local -type f -perm /a=x \
       \( -execdir unstripped '{}' \; -fprint ubins -o -fprint sbins \)

4 File Name Databases

The file name databases used by locate contain lists of files that were in particular directory trees when the databases were last updated. The file name of the default database is determined when locate and updatedb are configured and installed. The frequency with which the databases are updated and the directories for which they contain entries depend on how often updatedb is run, and with which arguments.

You can obtain some statistics about the databases by using ‘locate --statistics’.

4.1 Database Locations

There can be multiple file name databases. Users can select which databases locate searches using the LOCATE_PATH environment variable or a command line option. The system administrator can choose the file name of the default database, the frequency with which the databases are updated, and the directories for which they contain entries. File name databases are updated by running the updatedb program, typically nightly.

In networked environments, it often makes sense to build a database at the root of each filesystem, containing the entries for that filesystem. updatedb is then run for each filesystem on the fileserver where that filesystem is on a local disk, to prevent thrashing the network.

See Section 8.3 [Invoking updatedb], page 65, for the description of the options to updatedb. These options can be used to specify which directories are indexed by each database file.

The default location for the locate database depends on how findutils is built, but the findutils installation accompanying this manual uses the default location

/usr/local/var/locatedb.

If no database exists at /usr/local/var/locatedb but the user did not specify where to look (by using ‘-d’ or setting LOCATE_PATH), then locate will also check for a “secure”

database in /var/lib/slocate/slocate.db.

4.2 Database Formats

The file name databases contain lists of files that were in particular directory trees when the databases were last updated. The file name database format changed starting with GNU locate version 4.0 to allow machines with different byte orderings to share the databases.

GNU locate can read both the old pre-findutils-4.0 database format and the ‘LOCATE02’ database format. Support for the old database format will shortly be removed from locate. It has already been removed from updatedb.

If you run ‘locate --statistics’, the resulting summary indicates the type of each locate database. You select which database format updatedb will use with the

‘--dbformat’ option.

The ‘slocate’ database format is very similar to ‘LOCATE02’ and is also supported (in

both updatedb and locate).

4.2.1 LOCATE02 Database Format

updatedb runs a program called frcode to front-compress the list of file names, which reduces the database size by a factor of 4 to 5. Front-compression (also known as incremental encoding) works as follows.

The database entries are a sorted list (case-insensitively, for users’ convenience). Since the list is sorted, each entry is likely to share a prefix (initial string) with the previous entry. Each database entry begins with an offset-differential count byte, which is the additional number of characters of prefix of the preceding entry to use beyond the number that the preceding entry is using of its predecessor. (The counts can be negative.) Following the count is a null-terminated ASCII remainder – the part of the name that follows the shared prefix.

If the offset-differential count is larger than can be stored in a byte (+/-127), the byte has the value 0x80 and the count follows in a 2-byte word, with the high byte first (network byte order).

Every database begins with a dummy entry for a file called LOCATE02, which locate checks for to ensure that the database file has the correct format; it ignores the entry in doing the search.

Databases cannot be concatenated together, even if the first (dummy) entry is trimmed from all but the first database. This is because the offset-differential count in the first entry of the second and following databases will be wrong.

In the output of ‘locate --statistics’, the new database format is referred to as

‘LOCATE02’.

4.2.2 Sample LOCATE02 Database

Sample input to frcode:

     /usr/src
     /usr/src/cmd/aardvark.c
     /usr/src/cmd/armadillo.c
     /usr/tmp/zoo

Length of the longest prefix of the preceding entry to share:

     0 /usr/src
     8 /cmd/aardvark.c
     14 rmadillo.c
     5 tmp/zoo

Output from frcode, with trailing nulls changed to newlines and count bytes made printable:

     0 LOCATE02
     0 /usr/src
     8 /cmd/aardvark.c
     6 rmadillo.c
     -9 tmp/zoo

(6 = 14 - 8, and -9 = 5 - 14)

4.2.3 slocate Database Format

The slocate program uses a database format similar to, but not quite the same as, GNU locate. The first byte of the database specifies its security level. If the security level is 0, slocate will read, match and print filenames on the basis of the information in the database only. However, if the security level byte is 1, slocate omits entries from its output if the invoking user is unable to access them. The second byte of the database is zero. The second byte is immediately followed by the first database entry. The first entry in the database is not preceded by any differential count or dummy entry. Instead the differential count for the first item is assumed to be zero.

Starting with the second entry (if any) in the database, data is interpreted as for the GNU LOCATE02 format.

4.2.4 Old Database Format

The old database format is used by Unix locate and find programs and pre-4.0 releases of GNU findutils. locate understands this format, though updatedb will no longer produce it.

The old format differs from ‘LOCATE02’ in the following ways. Instead of each entry starting with an offset-differential count byte and ending with a null, byte values from 0 through 28 indicate offset-differential counts from -14 through 14. The byte value indicating that a long offset-differential count follows is 0x1e (30), not 0x80. The long counts are stored in host byte order, which is not necessarily network byte order, and host integer word size, which is usually 4 bytes. They also represent a count 14 less than their value. The database lines have no termination byte; the start of the next line is indicated by its first byte having a value <= 30.

In addition, instead of starting with a dummy entry, the old database format starts with a 256 byte table containing the 128 most common bigrams in the file list. A bigram is a pair of adjacent bytes. Bytes in the database that have the high bit set are indexes (with the high bit cleared) into the bigram table. The bigram and offset-differential count coding makes these databases 20-25% smaller than the new format, but makes them not 8-bit clean. Any byte in a file name that is in the ranges used for the special codes is replaced in the database by a question mark, which not coincidentally is the shell wildcard to match a single character. The old format therefore cannot faithfully store entries with non-ASCII characters.

Because the long counts are stored as native-order machine words, the database format is not easily used in environments which differ in terms of byte order. If locate databases are to be shared between machines, the ‘LOCATE02’ database format should be used. This has other benefits as discussed above. However, the length of the filename currently being processed can normally be used to place reasonable limits on the long counts and so this information is used by locate to help it guess the byte ordering of the old format database. Unless it finds evidence to the contrary, locate will assume that the byte order of the database is the same as the native byte order of the machine running locate. The output of ‘locate --statistics’ also includes information about the byte order of old-format databases.

The output of ‘locate --statistics’ will give an incorrect count of the number of file names containing newlines or high-bit characters for old-format databases.

Old versions of GNU locate fail to correctly handle very long file names, possibly leading to security problems relating to a heap buffer overrun. See Section 11.4 [Security Considerations for locate], page 98, for a detailed explanation.

4.3 Newline Handling

Within the database, file names are terminated with a null character. This is the case for both the old and the new format.

When the new database format is being used, the compression technique used to generate the database though relies on the ability to sort the list of files before they are presented to

frcode.

If the system’s sort command allows separating its input list of files with null characters via the ‘-z’ option, this option is used and therefore updatedb and locate will both correctly handle file names containing newlines. If the sort command lacks support for this, the list of files is delimited with the newline character, meaning that parts of file names containing newlines will be incorrectly sorted. This can result in both incorrect matches and incorrect failures to match.

5 File Permissions

Each file has a set of permissions that control the kinds of access that users have to that file. The permissions for a file are also called its access mode. They can be represented either in symbolic form or as an octal number.

5.1 Structure of File Permissions

There are three kinds of permissions that a user can have for a file:

1. permission to read the file. For directories, this means permission to list the contents of the directory.

2. permission to write to (change) the file. For directories, this means permission to create and remove files in the directory.

3. permission to execute the file (run it as a program). For directories, this means permission to access files in the directory.

There are three categories of users who may have different permissions to perform any of the above operations on a file:

1. the file’s owner;

2. other users who are in the file’s group;

3. everyone else.

Files are given an owner and group when they are created. Usually the owner is the current user and the group is the group of the directory the file is in, but this varies with the operating system, the file system the file is created on, and the way the file is created. You can change the owner and group of a file by using the chown and chgrp commands.

In addition to the three sets of three permissions listed above, a file’s permissions have three special components, which affect only executable files (programs) and, on some systems, directories:

1. Set the process’s effective user ID to that of the file upon execution (called the setuid bit). No effect on directories.

2. Set the process’s effective group ID to that of the file upon execution (called the setgid bit). For directories on some systems, put files created in the directory into the same group as the directory, no matter what group the user who creates them is in.

3. prevent users from removing or renaming a file in a directory unless they own the file or the directory; this is called the restricted deletion flag for the directory. For regular files on some systems, save the program’s text image on the swap device so it will load more quickly when run; this is called the sticky bit.

In addition to the permissions listed above, there may be file attributes specific to the file system, e.g: access control lists (ACLs), whether a file is compressed, whether a file can be modified (immutability), whether a file can be dumped. These are usually set using programs specific to the file system. For example:

ext2 On GNU and GNU/Linux the file permissions (“attributes”) specific to the ext2 file system are set using chattr.

FFS On FreeBSD the file permissions (“flags”) specific to the FFS file system are set using chrflags.

Although a file’s permission “bits” allow an operation on that file, that operation may still fail, because:

• the file-system-specific permissions do not permit it;

• the file system is mounted as read-only.

For example, if the immutable attribute is set on a file, it cannot be modified, regardless of the fact that you may have just run chmod a+w FILE.

5.2 Symbolic Modes

Symbolic modes represent changes to files’ permissions as operations on single-character symbols. They allow you to modify either all or selected parts of files’ permissions, optionally based on their previous values, and perhaps on the current umask as well (see Section 5.2.6 [Umask and Protection], page 49).

The format of symbolic modes is:

     [ugoa...][+-=]perms...[,...]

where perms is either zero or more letters from the set ‘rwxXst’, or a single letter from the set ‘ugo’.

The following sections describe the operators and other details of symbolic modes.

5.2.1 Setting Permissions

The basic symbolic operations on a file’s permissions are adding, removing, and setting the permission that certain users have to read, write, and execute the file. These operations have the following format:

     users operation permissions

The spaces between the three parts above are shown for readability only; symbolic modes cannot contain spaces.

The users part tells which users’ access to the file is changed. It consists of one or more of the following letters (or it can be empty; see Section 5.2.6 [Umask and Protection], page 49, for a description of what happens then). When more than one of these letters is given, the order that they are in does not matter.

u the user who owns the file;

g other users who are in the file’s group;

o all other users;

a all users; the same as ‘ugo’.

The operation part tells how to change the affected users’ access to the file, and is one of the following symbols:

+ to add the permissions to whatever permissions the users already have for the file;

- to remove the permissions from whatever permissions the users already have for the file; = to make the permissions the only permissions that the users have for the file.

The permissions part tells what kind of access to the file should be changed; it is normally zero or more of the following letters. As with the users part, the order does not matter when more than one letter is given. Omitting the permissions part is useful only with the ‘=’ operation, where it gives the specified users no access at all to the file.

r the permission the users have to read the file;

w the permission the users have to write to the file;

x the permission the users have to execute the file.

For example, to give everyone permission to read and write a file, but not to execute it, use:

     a=rw

To remove write permission for all users other than the file’s owner, use:

     go-w

The above command does not affect the access that the owner of the file has to it, nor does it affect whether other users can read or execute the file.

To give everyone except a file’s owner no permission to do anything with that file, use the mode below. Other users could still remove the file, if they have write permission on the directory it is in.

     go=

Another way to specify the same thing is:

     og-rwx

5.2.2 Copying Existing Permissions

You can base a file’s permissions on its existing permissions. To do this, instead of using a series of ‘r’, ‘w’, or ‘x’ letters after the operator, you use the letter ‘u’, ‘g’, or ‘o’. For example, the mode

     o+g

adds the permissions for users who are in a file’s group to the permissions that other users have for the file. Thus, if the file started out as mode 664 (‘rw-rw-r--’), the above mode would change it to mode 666 (‘rw-rw-rw-’). If the file had started out as mode 741 (‘rwxr----x’), the above mode would change it to mode 745 (‘rwxr--r-x’). The ‘-’ and ‘=’ operations work analogously.

5.2.3 Changing Special Permissions

In addition to changing a file’s read, write, and execute permissions, you can change its special permissions. See Section 5.1 [Mode Structure], page 45, for a summary of these permissions.

To change a file’s permission to set the user ID on execution, use ‘u’ in the users part of the symbolic mode and ‘s’ in the permissions part.

To change a file’s permission to set the group ID on execution, use ‘g’ in the users part of the symbolic mode and ‘s’ in the permissions part.

To change a file’s permission to set the restricted deletion flag or sticky bit, omit the users part of the symbolic mode (or use ‘a’) and put ‘t’ in the permissions part.

For example, to add set-user-ID permission to a program, you can use the mode:

     u+s

To remove both set-user-ID and set-group-ID permission from it, you can use the mode:

     ug-s

To set the restricted deletion flag or sticky bit, you can use the mode:

     +t

The combination ‘o+s’ has no effect. On GNU systems the combinations ‘u+t’ and ‘g+t’ have no effect, and ‘o+t’ acts like plain ‘+t’.

The ‘=’ operator is not very useful with special permissions; for example, the mode:

     o=t

does set the restricted deletion flag or sticky bit, but it also removes all read, write, and execute permissions that users not in the file’s group might have had for it.

5.2.4 Conditional Executability

There is one more special type of symbolic permission: if you use ‘X’ instead of ‘x’, execute permission is affected only if the file is a directory or already had execute permission.

For example, this mode:

     a+X

gives all users permission to search directories, or to execute files if anyone could execute them before.

5.2.5 Making Multiple Changes

The format of symbolic modes is actually more complex than described above (see Section 5.2.1 [Setting Permissions], page 46). It provides two ways to make multiple changes to files’ permissions.

The first way is to specify multiple operation and permissions parts after a users part in the symbolic mode.

For example, the mode:

     og+rX-w

gives users other than the owner of the file read permission and, if it is a directory or if someone already had execute permission to it, gives them execute permission; and it also denies them write permission to the file. It does not affect the permission that the owner of the file has for it. The above mode is equivalent to the two modes:

     og+rX
     og-w

The second way to make multiple changes is to specify more than one simple symbolic mode, separated by commas. For example, the mode:

     a+r,go-w

gives everyone permission to read the file and removes write permission on it for all users except its owner. Another example:

     u=rwx,g=rx,o=

sets all of the non-special permissions for the file explicitly. (It gives users who are not in the file’s group no permission at all for it.)

The two methods can be combined. The mode:

     a+r,g+x-w

gives all users permission to read the file, and gives users who are in the file’s group permission to execute it, as well, but not permission to write to it. The above mode could be written in several different ways; another is:

     u+r,g+rx,o+r,g-w

5.2.6 The Umask and Protection

If the users part of a symbolic mode is omitted, it defaults to ‘a’ (affect all users), except that any permissions that are set in the system variable umask are not affected. The value of umask can be set using the umask command. Its default value varies from system to system.

Omitting the users part of a symbolic mode is generally not useful with operations other than ‘+’. It is useful with ‘+’ because it allows you to use umask as an easily customizable protection against giving away more permission to files than you intended to.

As an example, if umask has the value 2, which removes write permission for users who are not in the file’s group, then the mode:

     +w

adds permission to write to the file to its owner and to other users who are in the file’s group, but not to other users. In contrast, the mode:

     a+w

ignores umask, and does give write permission for the file to all users.

5.3 Numeric Modes

As an alternative to giving a symbolic mode, you can give an octal (base 8) number that represents the new mode. This number is always interpreted in octal; you do not have to add a leading 0, as you do in C. Mode 0055 is the same as mode 55.

A numeric mode is usually shorter than the corresponding symbolic mode, but it is limited in that it cannot take into account a file’s previous permissions; it can only set them absolutely.

The permissions granted to the user, to other users in the file’s group, and to other users not in the file’s group each require three bits, which are represented as one octal digit. The three special permissions also require one bit each, and they are as a group represented as another octal digit. Here is how the bits are arranged, starting with the lowest valued bit:

     Value in  Corresponding
     Mode      Permission
               Other users not in the file's group:

        1      Execute
        2      Write
        4      Read
               Other users in the file's group:
       10      Execute
       20      Write
       40      Read
               The file's owner:
      100      Execute
      200      Write
      400      Read
               Special permissions:
     1000      Restricted deletion flag or sticky bit
     2000      Set group ID on execution
     4000      Set user ID on execution

For example, numeric mode 4755 corresponds to symbolic mode ‘u=rwxs,go=rx’, and numeric mode 664 corresponds to symbolic mode ‘ug=rw,o=r’. Numeric mode 0 corresponds to symbolic mode ‘a=’.

6 Date input formats

First, a quote:

Our units of temporal measurement, from seconds on up to months, are so complicated, asymmetrical and disjunctive so as to make coherent mental reckoning in time all but impossible. Indeed, had some tyrannical god contrived to enslave our minds to time, to make it all but impossible for us to escape subjection to sodden routines and unpleasant surprises, he could hardly have done better than handing down our present system. It is like a set of trapezoidal building blocks, with no vertical or horizontal surfaces, like a language in which the simplest thought demands ornate constructions, useless particles and lengthy circumlocutions. Unlike the more successful patterns of language and science, which enable us to face experience boldly or at least level-headedly, our system of temporal calculation silently and persistently encourages our terror of time. … It is as though architects had to measure length in feet, width in meters and height in ells; as though basic instruction manuals demanded a knowledge of five different languages. It is no wonder then that we often look into our own immediate past or future, last Tuesday or a week from Sunday, with feelings of helpless confusion. … —Robert Grudin, Time and the Art of Living.

This section describes the textual date representations that GNU programs accept. These are the strings you, as a user, can supply as arguments to the various programs. The C interface (via the parse_datetime function) is not described here.

6.1 General date syntax

A date is a string, possibly empty, containing many items separated by whitespace. The whitespace may be omitted when no ambiguity arises. The empty string means the beginning of today (i.e., midnight). Order of the items is immaterial. A date string may contain many flavors of items:

• calendar date items

• time of day items

• time zone items

• combined date and time of day items

• day of the week items

• relative items

• pure numbers.

We describe each of these item types in turn, below.

A few ordinal numbers may be written out in words in some contexts. This is most useful for specifying day of the week items or relative items (see below). Among the most commonly used ordinal numbers, the word ‘last’ stands for −1, ‘this’ stands for 0, and ‘first’ and ‘next’ both stand for 1. Because the word ‘second’ stands for the unit of time there is no way to write the ordinal number 2, but for convenience ‘third’ stands for 3, ‘fourth’ for 4, ‘fifth’ for 5, ‘sixth’ for 6, ‘seventh’ for 7, ‘eighth’ for 8, ‘ninth’ for 9, ‘tenth’ for 10, ‘eleventh’ for 11 and ‘twelfth’ for 12.

When a month is written this way, it is still considered to be written numerically, instead of being “spelled in full”; this changes the allowed strings.

In the current implementation, only English is supported for words and abbreviations

like ‘AM’, ‘DST’, ‘EST’, ‘first’, ‘January’, ‘Sunday’, ‘tomorrow’, and ‘year’.

The output of the date command is not always acceptable as a date string, not only because of the language problem, but also because there is no standard meaning for time zone items like ‘IST’. When using date to generate a date string intended to be parsed later, specify a date format that is independent of language and that does not use time zone items other than ‘UTC’ and ‘Z’. Here are some ways to do this:

$ LC_ALL=C TZ=UTC0 date
Tue Nov 15 02:02:42 UTC 2022
$ TZ=UTC0 date +'%Y-%m-%d %H:%M:%SZ'
2022-11-15 02:02:42Z
$ date --rfc-3339=ns  # --rfc-3339 is a GNU extension.
2022-11-14 21:02:42.000000000-05:00
$ date --rfc-email  # a GNU extension
Mon, 14 Nov 2022 21:02:42 -0500
$ date +'%Y-%m-%d %H:%M:%S %z'  # %z is a GNU extension.
2022-11-14 21:02:42 -0500
$ date +'@%s.%N'  # %s and %N are GNU extensions.
@1668477762.692722128

Alphabetic case is completely ignored in dates. Comments may be introduced between round parentheses, as long as included parentheses are properly nested. Hyphens not followed by a digit are currently ignored. Leading zeros on numbers are ignored.

Invalid dates like ‘2022-02-29’ or times like ‘24:00’ are rejected. In the typical case of a host that does not support leap seconds, a time like ‘23:59:60’ is rejected even if it corresponds to a valid leap second.

6.2 Calendar date items

A calendar date item specifies a day of the year. It is specified differently, depending on whether the month is specified numerically or literally. All these strings specify the same calendar date:

2022-11-14     # ISO 8601.
22-11-14       # Assume 19xx for 69 through 99,
               # 20xx for 00 through 68 (not recommended).
11/14/2022     # Common U.S. writing.
14 November 2022
14 Nov 2022    # Three-letter abbreviations always allowed.
November 14, 2022
14-nov-2022
14nov2022

The year can also be omitted. In this case, the last specified year is used, or the current year if none. For example:

     11/14
     nov 14

Here are the rules.

For numeric months, the ISO 8601 format ‘year-month-day’ is allowed, where year is any positive number, month is a number between 01 and 12, and day is a number between 01 and 31. A leading zero must be present if a number is less than ten. If year is 68 or smaller, then 2000 is added to it; otherwise, if year is less than 100, then 1900 is added to it. The construct ‘month/day/year’, popular in the United States, is accepted. Also ‘month/day’, omitting the year.

Literal months may be spelled out in full: ‘January’, ‘February’, ‘March’, ‘April’, ‘May’,

‘June’, ‘July’, ‘August’, ‘September’, ‘October’, ‘November’ or ‘December’. Literal months

may be abbreviated to their first three letters, possibly followed by an abbreviating dot. It is also permitted to write ‘Sept’ instead of ‘September’.

When months are written literally, the calendar date may be given as any of the following:

     day month year
     day month
     month day year
     day-month-year

Or, omitting the year:

     month day

6.3 Time of day items

A time of day item in date strings specifies the time on a given day. Here are some examples, all of which represent the same time:

     20:02:00.000000
     20:02
     8:02pm
     20:02-0500      # In EST (U.S. Eastern Standard Time).

More generally, the time of day may be given as ‘hour:minute:second’, where hour is a number between 0 and 23, minute is a number between 0 and 59, and second is a number between 0 and 59 possibly followed by ‘.’ or ‘,’ and a fraction containing one or more digits. Alternatively, ‘:second’ can be omitted, in which case it is taken to be zero. On the rare hosts that support leap seconds, second may be 60.

If the time is followed by ‘am’ or ‘pm’ (or ‘a.m.’ or ‘p.m.’), hour is restricted to run from 1 to 12, and ‘:minute’ may be omitted (taken to be zero). ‘am’ indicates the first half of the day, ‘pm’ indicates the second half of the day. In this notation, 12 is the predecessor of 1: midnight is ‘12am’ while noon is ‘12pm’. (This is the zero-oriented interpretation of ‘12am’ and ‘12pm’, as opposed to the old tradition derived from Latin which uses ‘12m’ for noon and ‘12pm’ for midnight.)

The time may alternatively be followed by a time zone correction, expressed as ‘shhmm’, where s is ‘+’ or ‘-’, hh is a number of zone hours and mm is a number of zone minutes. The zone minutes term, mm, may be omitted, in which case the one- or two-digit correction is interpreted as a number of hours. You can also separate hh from mm with a colon. When a time zone correction is given this way, it forces interpretation of the time relative to Coordinated Universal Time (UTC), overriding any previous specification for the time zone or the local time zone. For example, ‘+0530’ and ‘+05:30’ both stand for the time zone 5.5 hours ahead of UTC (e.g., India). This is the best way to specify a time zone correction by fractional parts of an hour. The maximum zone correction is 24 hours.

Either ‘am’/‘pm’ or a time zone correction may be specified, but not both.

6.4 Time zone items

A time zone item specifies an international time zone, indicated by a small set of letters, e.g., ‘UTC’ or ‘Z’ for Coordinated Universal Time. Any included periods are ignored. By following a non-daylight-saving time zone by the string ‘DST’ in a separate word (that is, separated by some white space), the corresponding daylight saving time zone may be specified. Alternatively, a non-daylight-saving time zone can be followed by a time zone correction, to add the two values. This is normally done only for ‘UTC’; for example, ‘UTC+05:30’ is equivalent to ‘+05:30’.

Time zone items other than ‘UTC’ and ‘Z’ are obsolescent and are not recommended, because they are ambiguous; for example, ‘EST’ has a different meaning in Australia than in the United States, and ‘A’ has different meaning as a military time zone than as an obsolete RFC 822 time zone. Instead, it’s better to use unambiguous numeric time zone corrections like ‘-0500’, as described in the previous section.

If neither a time zone item nor a time zone correction is supplied, timestamps are interpreted using the rules of the default time zone (see Section 6.10 [Specifying time zone rules], page 56).

6.5 Combined date and time of day items

The ISO 8601 date and time of day extended format consists of an ISO 8601 date, a ‘T’ character separator, and an ISO 8601 time of day. This format is also recognized if the ‘T’ is replaced by a space.

In this format, the time of day should use 24-hour notation. Fractional seconds are allowed, with either comma or period preceding the fraction. ISO 8601 fractional minutes and hours are not supported. Typically, hosts support nanosecond timestamp resolution; excess precision is silently discarded.

Here are some examples:

     2022-09-24T20:02:00.052-05:00
     2022-12-31T23:59:59,999999999+11:00
     1970-01-01 00:00Z

6.6 Day of week items

The explicit mention of a day of the week will forward the date (only if necessary) to reach that day of the week in the future.

Days of the week may be spelled out in full: ‘Sunday’, ‘Monday’, ‘Tuesday’, ‘Wednesday’, ‘Thursday’, ‘Friday’ or ‘Saturday’. Days may be abbreviated to their first three letters, optionally followed by a period. The special abbreviations ‘Tues’ for ‘Tuesday’, ‘Wednes’ for ‘Wednesday’ and ‘Thur’ or ‘Thurs’ for ‘Thursday’ are also allowed.

A number may precede a day of the week item to move forward supplementary weeks. It is best used in expression like ‘third monday’. In this context, ‘last day’ or ‘next day’ is also acceptable; they move one week before or after the day that day by itself would represent.

A comma following a day of the week item is ignored.

6.7 Relative items in date strings

Relative items adjust a date (or the current date if none) forward or backward. The effects of relative items accumulate. Here are some examples:

     1 year
     1 year ago
     3 years
     2 days

The unit of time displacement may be selected by the string ‘year’ or ‘month’ for moving by whole years or months. These are fuzzy units, as years and months are not all of equal duration. More precise units are ‘fortnight’ which is worth 14 days, ‘week’ worth 7 days, ‘day’ worth 24 hours, ‘hour’ worth 60 minutes, ‘minute’ or ‘min’ worth 60 seconds, and ‘second’ or ‘sec’ worth one second. An ‘s’ suffix on these units is accepted and ignored.

The unit of time may be preceded by a multiplier, given as an optionally signed number. Unsigned numbers are taken as positively signed. No number at all implies 1 for a multiplier. Following a relative item by the string ‘ago’ is equivalent to preceding the unit by a multiplier with value −1.

The string ‘tomorrow’ is worth one day in the future (equivalent to ‘day’), the string ‘yesterday’ is worth one day in the past (equivalent to ‘day ago’).

The strings ‘now’ or ‘today’ are relative items corresponding to zero-valued time displacement, these strings come from the fact a zero-valued time displacement represents the current time when not otherwise changed by previous items. They may be used to stress other items, like in ‘12:00 today’. The string ‘this’ also has the meaning of a zero-valued time displacement, but is preferred in date strings like ‘this thursday’.

When a relative item causes the resulting date to cross a boundary where the clocks were adjusted, typically for daylight saving time, the resulting date and time are adjusted accordingly.

The fuzz in units can cause problems with relative items. For example, ‘2022-12-31 -1 month’ might evaluate to 2022-12-01, because 2022-11-31 is an invalid date. To determine the previous month more reliably, you can ask for the month before the 15th of the current month. For example:

     $ date -R
     Thu, 31 Dec 2022 13:02:39 -0400
     $ date --date='-1 month' +'Last month was %B?'
     Last month was December?
     $ date --date="$(date +%Y-%m-15) -1 month" +'Last month was %B!'
     Last month was November!

Also, take care when manipulating dates around clock changes such as daylight saving leaps. In a few cases these have added or subtracted as much as 24 hours from the clock, so it is often wise to adopt universal time by setting the TZ environment variable to ‘UTC0’ before embarking on calendrical calculations.

6.8 Pure numbers in date strings

The precise interpretation of a pure decimal number depends on the context in the date string.

If the decimal number is of the form yyyymmdd and no other calendar date item (see Section 6.2 [Calendar date items], page 52) appears before it in the date string, then yyyy is read as the year, mm as the month number and dd as the day of the month, for the specified calendar date.

If the decimal number is of the form hhmm and no other time of day item appears before it in the date string, then hh is read as the hour of the day and mm as the minute of the hour, for the specified time of day. mm can also be omitted.

If both a calendar date and a time of day appear to the left of a number in the date string, but no relative item, then the number overrides the year.

6.9 Seconds since the Epoch

If you precede a number with ‘@’, it represents an internal timestamp as a count of seconds. The number can contain an internal decimal point (either ‘.’ or ‘,’); any excess precision not supported by the internal representation is truncated toward minus infinity. Such a number cannot be combined with any other date item, as it specifies a complete timestamp.

Internally, computer times are represented as a count of seconds since an Epoch—a welldefined point of time. On GNU and POSIX systems, the Epoch is 1970-01-01 00:00:00 UTC, so ‘@0’ represents this time, ‘@1’ represents 1970-01-01 00:00:01 UTC, and so forth. GNU and most other POSIX-compliant systems support such times as an extension to POSIX, using negative counts, so that ‘@-1’ represents 1969-12-31 23:59:59 UTC.

Most modern systems count seconds with 64-bit two’s-complement integers of seconds with nanosecond subcounts, which is a range that includes the known lifetime of the universe with nanosecond resolution. Some obsolescent systems count seconds with 32-bit two’scomplement integers and can represent times from 1901-12-13 20:45:52 through 2038-01-19 03:14:07 UTC. A few systems sport other time ranges.

On most hosts, these counts ignore the presence of leap seconds. For example, on most hosts ‘@1483228799’ represents 2016-12-31 23:59:59 UTC, ‘@1483228800’ represents 2017-01-01 00:00:00 UTC, and there is no way to represent the intervening leap second 2016-12-31 23:59:60 UTC.

6.10 Specifying time zone rules

Normally, dates are interpreted using the rules of the current time zone, which in turn are specified by the TZ environment variable, or by a system default if TZ is not set. To specify a different set of default time zone rules that apply just to one date, start the date with a string of the form ‘TZ="rule"’. The two quote characters (‘"’) must be present in the date, and any quotes or backslashes within rule must be escaped by a backslash.

For example, with the GNU date command you can answer the question “What time is it in New York when a Paris clock shows 6:30am on October 31, 2022?” by using a date beginning with ‘TZ="Europe/Paris"’ as shown in the following shell transcript:

     $ export TZ="America/New_York"
     $ date --date='TZ="Europe/Paris" 2022-10-31 06:30'
     Mon Oct 31 01:30:00 EDT 2022

In this example, the --date operand begins with its own TZ setting, so the rest of that operand is processed according to ‘Europe/Paris’ rules, treating the string ‘2022-11-14 06:30’ as if it were in Paris. However, since the output of the date command is processed according to the overall time zone rules, it uses New York time. (Paris was normally six hours ahead of New York in 2022, but this example refers to a brief Halloween period when the gap was five hours.)

A TZ value is a rule that typically names a location in the ‘tz’ database (https:// www.iana.org/time-zones). A recent catalog of location names appears in the TWiki

Date and Time Gateway (https://twiki.org/cgi-bin/xtra/tzdatepick.html).

A few non-GNU hosts require a colon before a location name in a TZ setting, e.g.,

‘TZ=":America/New_York"’.

The ‘tz’ database includes a wide variety of locations ranging from ‘Africa/Abidjan’ to ‘Pacific/Tongatapu’, but if you are at sea and have your own private time zone, or if you are using a non-GNU host that does not support the ‘tz’ database, you may need to use a POSIX rule instead. The previously-mentioned POSIX rule ‘UTC0’ says that the time zone abbreviation is ‘UTC’, the zone is zero hours away from Greenwich, and there is no daylight saving time. POSIX rules can also specify nonzero Greenwich offsets. For example, the following shell transcript answers the question “What time is it five and a half hours east of Greenwich when a clock seven hours west of Greenwich shows 9:50pm on July 12, 2022?”

     $ TZ="<+0530>-5:30" date --date='TZ="<-07>+7" 2022-07-12 21:50'
     Wed Jul 13 10:20:00 +0530 2022

This example uses the somewhat-confusing POSIX convention for rules. ‘TZ="<-07>+7"’ says that the time zone abbreviation is ‘-07’ and the time zone is 7 hours west of Greenwich, and ‘TZ="<+0530>-5:30"’ says that the time zone abbreviation is ‘+0530’ and the time zone is 5 hours 30 minutes east of Greenwich. (One should never use a setting like ‘TZ="UTC-5"’, since this would incorrectly imply that local time is five hours east of Greenwich and the time zone is called “UTC”.) Although trickier POSIX TZ settings like ‘TZ="<-05>+5<-04>,M3.2.0/2,M11.1.0/2"’ can specify some daylight saving regimes, location-based settings like ‘TZ="America/New_York"’ are typically simpler and more accurate historically. See Section “Specifying the Time Zone with TZ” in The GNU C Library.

6.11 Authors of parse_datetime

parse_datetime started life as getdate, as originally implemented by Steven M. Bellovin (smb@research.att.com) while at the University of North Carolina at Chapel Hill. The code was later tweaked by a couple of people on Usenet, then completely overhauled by Rich $alz (rsalz@bbn.com) and Jim Berets (jberets@bbn.com) in August, 1990. Various revisions for the GNU system were made by David MacKenzie, Jim Meyering, Paul Eggert and others, including renaming it to get_date to avoid a conflict with the alternative Posix function getdate, and a later rename to parse_datetime. The Posix function getdate can parse more locale-specific dates using strptime, but relies on an environment variable and external file, and lacks the thread-safety of parse_datetime.

This chapter was originally produced by François Pinard (pinard@iro.umontreal.ca) from the parse_datetime.y source code, and then edited by K. Berry (kb@cs.umb.edu).

7 Configuration

The findutils source distribution includes a configure script which examines the system and generates files required to build findutils. See the files README and INSTALL.

A number of options can be specified on the configure command line, and many of these are straightforward, adequately documented in the --help output, or not normally useful. Options which are useful or which are not obvious are explained here.

7.1 Leaf Optimisation

Files in Unix file systems have a link count which indicates how many names point to the same inode. Directories in Unix filssytems have a .. entry which functions as a hard link to the parent directory and a . entry which functions as a link to the directory itself. The .. entry of the root directory also points to the root. This means that find can deduce the number of subdirectories a directory has, simply by subtracting 2 from the directory’s link count. This allows find saving stat calls which would otherwise be needed to discover which directory entries are subdirectories.

File systems which don’t have these semantics should simply return a value less than 2 in the st_nlinks member of struct stat in response to a successful call to stat.

If you are building find for a system on which the value of st_nlinks is unreliable, you can specify --disable-leaf-optimisation to configure to prevent this assumption being made.

7.2 d type Optimisation

When this feature is enabled, find takes advantage of the fact that on some systems readdir will return the type of a file in struct dirent.

8 Reference

Below are summaries of the command line syntax for the programs discussed in this manual.

8.1 Invoking find

find [-H] [-L] [-P] [-D debugoptions] [-Olevel] [file...] [expression]

find searches the directory tree rooted at each file name file by evaluating the expression on each file it finds in the tree.

The command line may begin with the ‘-H’, ‘-L’, ‘-P’, ‘-D’ and ‘-O’ options. These are followed by a list of files or directories that should be searched. If no files to search are specified, the current directory (.) is used.

This list of files to search is followed by a list of expressions describing the files we wish to search for. The first part of the expression is recognised by the fact that it begins with ‘-’ followed by some other letters (for example ‘-print’), or is either ‘(’ or ‘!’. Any arguments after it are the rest of the expression.

If no expression is given, the expression ‘-print’ is used.

The find command exits with status zero if all files matched are processed successfully, greater than zero if errors occur.

The find program also recognises two options for administrative use:

‘--help’ Print a summary of the command line usage and exit.

‘--version’

Print the version number of find and exit.

The ‘-version’ option is a synonym for ‘--version’

8.1.1 Filesystem Traversal Options

The options ‘-H’, ‘-L’ or ‘-P’ may be specified at the start of the command line (if none of these is specified, ‘-P’ is assumed). If you specify more than one of these options, the last one specified takes effect (but note that the ‘-follow’ option is equivalent to ‘-L’).

-P Never follow symbolic links (this is the default), except in the case of the ‘-xtype’ predicate.

-L Always follow symbolic links, except in the case of the ‘-xtype’ predicate.

-H Follow symbolic links specified in the list of files to search, or which are otherwise specified on the command line.

If find would follow a symbolic link, but cannot for any reason (for example, because it has insufficient permissions or the link is broken), it falls back on using the properties of the symbolic link itself. Section 2.4.1 [Symbolic Links], page 10, for a more complete description of how symbolic links are handled.

8.1.2 Warning Messages

If there is an error on the find command line, an error message is normally issued. However, there are some usages that are inadvisable but which find should still accept. Under these circumstances, find may issue a warning message.

By default, warnings are enabled only if find is being run interactively (specifically, if the standard input is a terminal) and the POSIXLY_CORRECT environment variable is not set. Warning messages can be controlled explicitly by the use of options on the command line:

-warn Issue warning messages where appropriate.

-nowarn Do not issue warning messages.

These options take effect at the point on the command line where they are specified. Therefore it’s not useful to specify ‘-nowarn’ at the end of the command line. The warning messages affected by the above options are triggered by: − Use of the ‘-d’ option which is deprecated; please use ‘-depth’ instead, since the latter

is POSIX-compliant. − Specifying an option (for example ‘-mindepth’) after a non-option (for example ‘-type’

or ‘-print’) on the command line. − Use of the ‘-name’ or ‘-iname’ option with a slash character in the pattern. Since the

name predicates only compare against the basename of the visited files, the only file that can match a slash is the root directory itself.

The default behaviour above is designed to work in that way so that existing shell scripts don’t generate spurious errors, but people will be made aware of the problem.

Some warning messages are issued for less common or more serious problems, and consequently cannot be turned off: − Use of an unrecognised backslash escape sequence with ‘-fprintf’ − Use of an unrecognised formatting directive with ‘-fprintf’

8.1.3 Optimisation Options

The ‘-Olevel’ option sets find’s optimisation level to level. The default optimisation level is 1.

At certain optimisation levels (but not by default), find reorders tests to speed up execution while preserving the overall effect; that is, predicates with side effects are not reordered relative to each other. The optimisations performed at each optimisation level are as follows.

0 Currently equivalent to optimisation level 1.

1 This is the default optimisation level and corresponds to the traditional behaviour. Expressions are reordered so that tests based only on the names of files (for example‘ -name’ and ‘-regex’) are performed first.

2 Any ‘-type’ or ‘-xtype’ tests are performed after any tests based only on the names of files, but before any tests that require information from the inode. On many modern versions of Unix, file types are returned by readdir() and so these predicates are faster to evaluate than predicates which need to stat the file first.

If you use the ‘-fstype FOO’ predicate and specify a filsystem type ‘FOO’ which is not known (that is, present in /etc/mtab) at the time find starts, that predicate is equivalent to ‘-false’.

3 At this optimisation level, the full cost-based query optimizer is enabled. The order of tests is modified so that cheap (i.e., fast) tests are performed first and more expensive ones are performed later, if necessary. Within each cost band, predicates are evaluated earlier or later according to whether they are likely to succeed or not. For ‘-o’, predicates which are likely to succeed are evaluated earlier, and for ‘-a’, predicates which are likely to fail are evaluated earlier.

The re-ordering of operations performed by the cost-based optimizer can result in uservisible behaviour change. For example, the ‘-readable’ and ‘-empty’ predicates are sensitive to re-ordering. If they are run in the order ‘-empty -readable’, an error message will be issued for unreadable directories. If they are run in the order ‘-readable -empty’, no error message will be issued. This is the reason why such operation re-ordering is not performed at the default optimisation level.

8.1.4 Debug Options

The ‘-D’ option makes find produce diagnostic output. Much of the information is useful only for diagnosing problems, and so most people will not find this option helpful.

The list of debug options should be comma separated. Compatibility of the debug options is not guaranteed between releases of findutils. For a complete list of valid debug options, see the output of find -D help.

Valid debug options include:

‘tree’ Show the expression tree in its original and optimized form.

‘stat’ Print messages as files are examined with the stat and lstat system calls. The find program tries to minimise such calls.

‘opt’ Prints diagnostic information relating to the optimisation of the expression tree; see the ‘-O’ option.

‘rates’ Prints a summary indicating how often each predicate succeeded or failed.

‘all’ Enable all of the other debug options (but ‘help’).

‘help’ Explain the debugging options.

8.1.5 Find Expressions

The final part of the find command line is a list of expressions. See [Primary Index], page 114, for a summary of all of the tests, actions, and options that the expression can contain. If the expression is missing, ‘-print’ is assumed.

8.2 Invoking locate

locate [option...] pattern...

For each pattern given locate searches one or more file name databases returning each match of pattern.

--all

-A Print only names which match all non-option arguments, not those matching one or more non-option arguments.

--basename

-b The specified pattern is matched against just the last component of the name of a file in the locate database. This last component is also called the “base name”. For example, the base name of /tmp/mystuff/foo.old.c is foo.old.c. If the pattern contains metacharacters, it must match the base name exactly. If not, it must match part of the base name.

--count

-c Instead of printing the matched file names, just print the total number of matches found, unless ‘--print’ (‘-p’) is also present.

--database=path

-d path Instead of searching the default locate database /usr/local/var/locatedb, locate searches the file name databases in path, which is a colon-separated list of database file names. You can also use the environment variable LOCATE_PATH to set the list of database files to search. The option overrides the environment variable if both are used. Empty elements in path (that is, a leading or trailing colon, or two colons in a row) are taken to stand for the default database. A database can be supplied on stdin, using ‘-’ as an element of ‘path’. If more than one element of ‘path’ is ‘-’, later instances are ignored (but a warning message is printed).

--existing

-e Only print out such names which currently exist (instead of such names which existed when the database was created). Note that this may slow down the program a lot, if there are many matches in the database. The way in which broken symbolic links are treated is affected by the ‘-L’, ‘-P’ and ‘-H’ options. Please note that it is possible for the file to be deleted after locate has checked that it exists, but before you use it. This option is automatically turned on when reading an slocate database in secure mode (see Section 4.2.3 [slocate Database Format], page 43).

--non-existing

-E Only print out such names which currently do not exist (instead of such names which existed when the database was created). Note that this may slow down the program a lot, if there are many matches in the database. The way in which broken symbolic links are treated is affected by the ‘-L’, ‘-P’ and ‘-H’ options. Please note that locate checks that the file does not exist, but a file of the same name might be created after locate’s check but before you read locate’s output.

--follow

-L If testing for the existence of files (with the ‘-e’ or ‘-E’ options), consider broken symbolic links to be non-existing. This is the default behaviour.

--nofollow
-P

-H If testing for the existence of files (with the ‘-e’ or ‘-E’ options), treat broken symbolic links as if they were existing files. The ‘-H’ form of this option is provided purely for similarity with find; the use of ‘-P’ is recommended over ‘-H’.

--ignore-case

-i Ignore case distinctions in both the pattern and the file names.

--limit=N

-l N Limit the number of results printed to N. When used with the ‘--count’ option, the value printed will never be larger than this limit.

--max-database-age=D

Normally, locate will issue a warning message when it searches a database which is more than 8 days old. This option changes that value to something other than 8. The effect of specifying a negative value is undefined.

--mmap

-m Accepted but does nothing. The option is supported only to provide compatibility with BSD’s locate.

--null

-0 Results are separated with the ASCII NUL character rather than the newline character. To get the full benefit of this option, use the new locate database format (that is the default anyway).

--print

-p Print search results when they normally would not be due to use of

‘--statistics’ (‘-S’) or ‘--count’ (‘-c’).
--wholename

-w The specified pattern is matched against the whole name of the file in the locate database. If the pattern contains metacharacters, it must match exactly. If not, it must match part of the whole file name. This is the default behaviour.

--regex

-r Instead of using substring or shell glob matching, the pattern specified on the command line is understood to be a regular expression. GNU Emacs-style regular expressions are assumed unless the ‘--regextype’ option is also given. File names from the locate database are matched using the specified regular expression. If the ‘-i’ flag is also given, matching is case-insensitive. Matches are performed against the whole path name, and so by default a pathname will be matched if any part of it matches the specified regular expression. The regular expression may use ‘^’ or ‘$’ to anchor a match at the beginning or end of a pathname.

--regextype

This option changes the regular expression syntax and behaviour used by the ‘--regex’ option. Section 8.5 [Regular Expressions], page 70, for more information on the regular expression dialects understood by GNU findutils.

--stdio

-s Accepted but does nothing. The option is supported only to provide compatibility with BSD’s locate.

--statistics

-S Print some summary information for each locate database. No search is performed unless non-option arguments are given. Although the BSD version of locate also has this option, the format of the output is different.

--help Print a summary of the command line usage for locate and exit.

--version

Print the version number of locate and exit.

8.3 Invoking updatedb

updatedb [option...]

updatedb creates and updates the database of file names used by locate. updatedb generates a list of files similar to the output of find and then uses utilities for optimizing the database for performance. updatedb is often run periodically as a cron job and configured with environment variables or command options. Typically, operating systems have a shell script that “exports” configurations for variable definitions and uses another shell script that “sources” the configuration file into the environment and then executes updatedb in the environment.

--findoptions='OPTION...'

Global options to pass on to find. The environment variable FINDOPTIONS also sets this value. Default is none.

--localpaths='path...'

Non-network directories to put in the database. Default is /.

--netpaths='path...'

Network (NFS, AFS, RFS, etc.) directories to put in the database. The environment variable NETPATHS also sets this value. Default is none.

--prunepaths='path...'

Directories to omit from the database, which would otherwise be included. The environment variable PRUNEPATHS also sets this value. Default is /tmp /usr/tmp /var/tmp /afs. The paths are used as regular expressions (with find ... -regex, so you need to specify these paths in the same way that find will encounter them. This means for example that the paths must not include trailing slashes.

--prunefs='path...'

Filesystems to omit from the database, which would otherwise be included. Note that files are pruned when a filesystem is reached; Any filesystem mounted under an undesired filesystem will be ignored. The environment variable PRUNEFS also sets this value. Default is nfs NFS proc.

--output=dbfile

The database file to build. The default is system-dependent, but when this document was formatted it was /usr/local/var/locatedb.

--localuser=user

The user to search the non-network directories as, using su. Default is to search the non-network directories as the current user. You can also use the environment variable LOCALUSER to set this user.

--netuser=user

The user to search network directories as, using su. Default user is daemon. You can also use the environment variable NETUSER to set this user.

--dbformat=FORMAT

Generate the locate database in format FORMAT. Supported database formats include LOCATE02 (which is the default) and slocate. The slocate format exists for compatibility with slocate. See Section 4.2 [Database Formats], page 41, for a detailed description of each format.

--help Print a summary of the command line usage and exit.

--version

Print the version number of updatedb and exit.

8.4 Invoking xargs

xargs [option...] [command [initial-arguments]]

xargs exits with the following status:

0 if it succeeds

123 if any invocation of the command exited with status 1-125

124 if the command exited with status 255

125 if the command is killed by a signal

126 if the command cannot be run

127 if the command is not found

1 if some other error occurred.

Exit codes greater than 128 are used by the shell to indicate that a program died due to a fatal signal.

8.4.1 xargs options

-- Delimit the option list. Later arguments, if any, are treated as operands even if they begin with ‘-’. For example, ‘xargs -- --help’ runs the command ‘--help’ (found in PATH) instead of printing the usage text, and ‘xargs -- --mycommand’ runs the command ‘--mycommand’ instead of rejecting this as unrecognized option.

--arg-file=inputfile

-a inputfile Read names from the file inputfile instead of standard input. If you use this option, the standard input stream remains unchanged when commands are run. Otherwise, stdin is redirected from /dev/null.

--null

-0 Input file names are terminated by a null character instead of by whitespace, and any quotes and backslash characters are not considered special (every character is taken literally). Disables the end of file string, which is treated like any other argument.

--delimiter delim
-d delim

Input file names are terminated by the specified character delim instead of by whitespace, and any quotes and backslash characters are not considered special (every character is taken literally). Disables the logical end-of-file marker string, which is treated like any other argument.

The specified delimiter may be a single character, a C-style character escape such as ‘\n’, or an octal or hexadecimal escape code. Octal and hexadecimal escape codes are understood as for the printf command. Multibyte characters are not supported.

-E eof-str
--eof[=eof-str]
-e[eof-str]

Set the logical end-of-file marker string to eof-str. If the logical end-of-file marker string occurs as a line of input, the rest of the input is ignored. If eof-str is omitted (‘-e’) or blank (either ‘-e’ or ‘-E’), there is no logical end-of-file marker string. The ‘-e’ form of this option is deprecated in favour of the POSIX-compliant ‘-E’ option, which you should use instead. As of GNU xargs version 4.2.9, the default behaviour of xargs is not to have a logical end-of-file marker string. The POSIX standard (IEEE Std 1003.1, 2004 Edition) allows this.

The logical end-of-file marker string is not treated specially if the ‘-d’ or the ‘-0’ options are in effect. That is, when either of these options are in effect, the whole input file will be read even if ‘-E’ was used.

--help Print a summary of the options to xargs and exit.

-I replace-str
--replace[=replace-str]
-i[replace-str]

Replace occurrences of replace-str in the initial arguments with names read from standard input. Also, unquoted blanks do not terminate arguments; instead, the input is split at newlines only. If replace-str is omitted (omitting it is allowed only for ‘-i’ and ‘--replace’), it defaults to ‘{}’ (like for ‘find -exec’). Implies ‘-x’ and ‘-L 1’. The ‘-i’ option is deprecated in favour of the ‘-I’ option.

-L max-lines
--max-lines[=max-lines]
-l[max-lines]

Use at most max-lines non-blank input lines per command line. For ‘-l’, maxlines defaults to 1 if omitted. For ‘-L’, the argument is mandatory. Trailing blanks cause an input line to be logically continued on the next input line, for the purpose of counting the lines. Implies ‘-x’. The ‘-l’ form of this option is deprecated in favour of the POSIX-compliant ‘-L’ option.

--max-args=max-args
-n max-args

Use at most max-args arguments per command line. Fewer than max-args arguments will be used if the size (see the ‘-s’ option) is exceeded, unless the ‘-x’ option is given, in which case xargs will exit.

--open-tty

-o Reopen stdin as /dev/tty in the child process before executing the command, thus allowing that command to be associated to the terminal while xargs reads from a different stream, e.g. from a pipe. This is useful if you want xargs to run an interactive application.

     grep -lZ PATTERN * | xargs -0o -n1 vi
--interactive

-p Prompt the user about whether to run each command line and read a line from the terminal. Only run the command line if the response starts with ‘y’ or ‘Y’. Implies ‘-t’.

--no-run-if-empty

-r If the standard input is completely empty, do not run the command. By default, the command is run once even if there is no input.

--max-chars=max-chars
-s max-chars

Use at most max-chars characters per command line, including the command, initial arguments and any terminating nulls at the ends of the argument strings.

--show-limits

Display the limits on the command-line length which are imposed by the operating system, xargs’ choice of buffer size and the ‘-s’ option. Pipe the input from /dev/null (and perhaps specify ‘--no-run-if-empty’) if you don’t want xargs to do anything.

--verbose

-t Print the command line on the standard error output before executing it.

--version

Print the version number of xargs and exit.

--exit

-x Exit if the size (see the ‘-s’ option) is exceeded.

--max-procs=max-procs
-P max-procs

Run simultaneously up to max-procs processes at once; the default is 1. If max-procs is 0, xargs will run as many processes as possible simultaneously. See Section 3.3.2.5 [Controlling Parallelism], page 35, for information on dynamically controlling parallelism.

--process-slot-var=environment-variable-name

Set the environment variable environment-variable-name to a unique value in each running child process. Each value is a decimal integer. Values are reused once child processes exit. This can be used in a rudimentary load distribution scheme, for example.

8.4.2 Conflicting options

The options ‘--max-lines’ (‘-L’, ‘-l’), ‘--replace’ (‘-I’, ‘-i’) and ‘--max-args’ (‘-n’) are mutually exclusive.

If some of them are specified at the same time, then xargs will generally use the option specified last on the command line, i.e., it will reset the value of the offending option (given before) to its default value. Additionally, xargs will issue a warning diagnostic on stderr.

     $ seq 4 | xargs -L2 -n3
     xargs: warning: options --max-lines and --max-args/-n are \
       mutually exclusive, ignoring previous --max-lines value
     1 2 3
     4

The exception to this rule is that the special max-args value 1 is ignored after the ‘--replace’ option and its short-option aliases ‘-I’ and ‘-i’, because it would not actually conflict.

     $ seq 2 | xargs --replace -n1 echo a-{}-b
     a-1-b
     a-2-b

8.4.3 Invoking the shell from xargs

Normally, xargs will exec the command you specified directly, without invoking a shell. This is normally the behaviour one would want. It’s somewhat more efficient and avoids problems with shell metacharacters, for example. However, sometimes it is necessary to manipulate the environment of a command before it is run, in a way that xargs does not directly support.

Invoking a shell from xargs is a good way of performing such manipulations. However, some care must be taken to prevent problems, for example unwanted interpretation of shell metacharacters.

This command moves a set of files into an archive directory:

     find /foo -maxdepth 1 -atime +366 -exec mv {} /archive \;

However, this will only move one file at a time. We cannot in this case use -exec ... + because the matched file names are added at the end of the command line, while the destination directory would need to be specified last. We also can’t use xargs in the obvious way for the same reason. One way of working around this problem is to make use of the special properties of GNU mv; it has a -t option that allows specifying the target directory before the list of files to be moved. However, while this technique works for GNU mv, it doesn’t solve the more general problem.

Here is a more general technique for solving this problem:

     find /foo -maxdepth 1 -atime +366 -print0 |
     xargs -r0 sh -c 'mv "$@" /archive' move

Here, a shell is being invoked. There are two shell instances to think about. The first is the shell which launches the xargs command (this might be the shell into which you are typing, for example). The second is the shell launched by xargs (in fact it will probably launch several, one after the other, depending on how many files need to be archived). We’ll refer to this second shell as a subshell.

Our example uses the -c option of sh. Its argument is a shell command to be executed by the subshell. Along with the rest of that command, the $@ is enclosed by single quotes to make sure it is passed to the subshell without being expanded by the parent shell. It is also enclosed with double quotes so that the subshell will expand $@ correctly even if one of the file names contains a space or newline.

The subshell will use any non-option arguments as positional parameters (that is, in the expansion of $@). Because xargs launches the sh -c subshell with a list of files, those files will end up as the expansion of $@.

You may also notice the ‘move’ at the end of the command line. This is used as the value of $0 by the subshell. We include it because otherwise the name of the first file to be moved would be used instead. If that happened it would not be included in the subshell’s expansion of $@, and so it wouldn’t actually get moved.

Another reason to use the sh -c construct could be to perform redirection:

     find /usr/include -name '*.h' | xargs grep -wl mode_t |
     xargs -r sh -c 'exec emacs "$@" < /dev/tty' Emacs

Notice that we use the shell builtin exec here. That’s simply because the subshell needs to do nothing once Emacs has been invoked. Therefore instead of keeping a sh process around for no reason, we just arrange for the subshell to exec Emacs, saving an extra process creation.

Although GNU xargs and the implementations on some other platforms like BSD support the ‘-o’ option to achieve the same, the above is the portable way to redirect stdin to

/dev/tty.

Sometimes, though, it can be helpful to keep the shell process around:

     find /foo -maxdepth 1 -atime +366 -print0 |
     xargs -r0 sh -c 'mv "$@" /archive || exit 255' move

Here, the shell will exit with status 255 if any mv failed. This causes xargs to stop immediately.

8.5 Regular Expressions

The ‘-regex’ and ‘-iregex’ tests of find allow matching by regular expression, as does the

‘--regex’ option of locate.

Your locale configuration affects how regular expressions are interpreted. See Section 8.6 [Environment Variables], page 80, for a description of how your locale setup affects the interpretation of regular expressions.

There are also several different types of regular expression, and these are interpreted differently. Normally, the type of regular expression used by find and locate is almost identical to that used in GNU Emacs. The single difference is that in find and locate, a ‘.’ will match a newline character.

Both find and locate provide an option which allows selecting an alternative regular expression syntax; for find this is the ‘-regextype’ option, and for locate this is the

‘--regextype’ option.

These options take a single argument, which indicates the specific regular expression syntax and behaviour that should be used. This should be one of the following:

8.5.1 ‘findutils-default’ regular expression syntax

The character ‘.’ matches any single character.

‘+’ indicates that the regular expression should match one or more occurrences of the previous atom or regexp.

‘?’ indicates that the regular expression should match zero or one occurrence of the previous atom or regexp.

‘\+’ matches a ‘+’

‘\?’ matches a ‘?’.

Bracket expressions are used to match ranges of characters. Bracket expressions where the range is backward, for example ‘[z-a]’, are ignored. Within square brackets, ‘\’ is taken literally. Character classes are not supported, so for example you would need to use

‘[0-9]’ instead of ‘[[:digit:]]’.

GNU extensions are supported:

1. ‘\w’ matches a character within a word

2. ‘\W’ matches a character which is not within a word

3. ‘\<’ matches the beginning of a word

4. ‘\>’ matches the end of a word

5. ‘\b’ matches a word boundary

6. ‘\B’ matches characters which are not a word boundary

7. ‘\` ’ matches the beginning of the whole input

8. ‘\'’ matches the end of the whole input

Grouping is performed with backslashes followed by parentheses ‘\(’, ‘\)’. A backslash followed by a digit acts as a back-reference and matches the same thing as the previous grouped expression indicated by that number. For example ‘\2’ matches the second group expression. The order of group expressions is determined by the position of their opening parenthesis ‘\(’.

The alternation operator is ‘\|’.

The character ‘^’ only represents the beginning of a string when it appears:

1. At the beginning of a regular expression 2. After an open-group, signified by ‘\(’

3. After the alternation operator ‘\|’ The character ‘$’ only represents the end of a string when it appears:

1. At the end of a regular expression

2. Before a close-group, signified by ‘\)’

3. Before the alternation operator ‘\|’ ‘*’, ‘+’ and ‘?’ are special at any point in a regular expression except:

1. At the beginning of a regular expression

2. After an open-group, signified by ‘\(’

3. After the alternation operator ‘\|’ The longest possible match is returned; this applies to the regular expression as a whole

and (subject to this constraint) to subexpressions within groups.

8.5.2 ‘emacs’ regular expression syntax

The character ‘.’ matches any single character except newline. ‘+’ indicates that the regular expression should match one or more occurrences of

the previous atom or regexp.

‘?’ indicates that the regular expression should match zero or one occurrence of

the previous atom or regexp.

‘\+’ matches a ‘+’ ‘\?’ matches a ‘?’.

Bracket expressions are used to match ranges of characters. Bracket expressions where the range is backward, for example ‘[z-a]’, are ignored. Within square brackets, ‘\’ is taken literally. Character classes are not supported, so for example you would need to use

‘[0-9]’ instead of ‘[[:digit:]]’.

GNU extensions are supported:

1. ‘\w’ matches a character within a word

2. ‘\W’ matches a character which is not within a word

3. ‘\<’ matches the beginning of a word

4. ‘\>’ matches the end of a word

5. ‘\b’ matches a word boundary

6. ‘\B’ matches characters which are not a word boundary

7. ‘\` ’ matches the beginning of the whole input

8. ‘\'’ matches the end of the whole input Grouping is performed with backslashes followed by parentheses ‘\(’, ‘\)’. A backslash

followed by a digit acts as a back-reference and matches the same thing as the previous grouped expression indicated by that number. For example ‘\2’ matches the second group expression. The order of group expressions is determined by the position of their opening parenthesis ‘\(’.

The alternation operator is ‘\|’.

The character ‘^’ only represents the beginning of a string when it appears:

1. At the beginning of a regular expression

2. After an open-group, signified by ‘\(’

3. After the alternation operator ‘\|’ The character ‘$’ only represents the end of a string when it appears:

1. At the end of a regular expression

2. Before a close-group, signified by ‘\)’

3. Before the alternation operator ‘\|’ ‘*’, ‘+’ and ‘?’ are special at any point in a regular expression except:

1. At the beginning of a regular expression

2. After an open-group, signified by ‘\(’

3. After the alternation operator ‘\|’ The longest possible match is returned; this applies to the regular expression as a whole

and (subject to this constraint) to subexpressions within groups.

8.5.3 ‘gnu-awk’ regular expression syntax

The character ‘.’ matches any single character. ‘+’ indicates that the regular expression should match one or more occurrences of

the previous atom or regexp.

‘?’ indicates that the regular expression should match zero or one occurrence of

the previous atom or regexp.

‘\+’ matches a ‘+’ ‘\?’ matches a ‘?’.

Bracket expressions are used to match ranges of characters. Bracket expressions where the range is backward, for example ‘[z-a]’, are invalid. Within square brackets, ‘\’ can be used to quote the following character. Character classes are supported; for example ‘[[:digit:]]’ will match a single decimal digit.

GNU extensions are supported:

1. ‘\w’ matches a character within a word

2. ‘\W’ matches a character which is not within a word

3. ‘\<’ matches the beginning of a word

4. ‘\>’ matches the end of a word

5. ‘\b’ matches a word boundary

6. ‘\B’ matches characters which are not a word boundary

7. ‘\` ’ matches the beginning of the whole input

8. ‘\'’ matches the end of the whole input

Grouping is performed with parentheses ‘()’. An unmatched ‘)’ matches just itself. A backslash followed by a digit acts as a back-reference and matches the same thing as the previous grouped expression indicated by that number. For example ‘\2’ matches the second group expression. The order of group expressions is determined by the position of their opening parenthesis ‘(’.

The alternation operator is ‘|’.

The characters ‘^’ and ‘$’ always represent the beginning and end of a string respectively, except within square brackets. Within brackets, ‘^’ can be used to invert the membership of the character class being specified.

‘*’, ‘+’ and ‘?’ are special at any point in a regular expression except:

1. At the beginning of a regular expression

2. After an open-group, signified by ‘(’

3. After the alternation operator ‘|’

Intervals are specified by ‘{’ and ‘}’. Invalid intervals are treated as literals, for example ‘a{1’ is treated as ‘a\{1’

The longest possible match is returned; this applies to the regular expression as a whole and (subject to this constraint) to subexpressions within groups.

8.5.4 ‘grep’ regular expression syntax

The character ‘.’ matches any single character.

‘\+’ indicates that the regular expression should match one or more occurrences of the previous atom or regexp.

‘\?’ indicates that the regular expression should match zero or one occurrence of the previous atom or regexp.

‘+ and ?’ match themselves.

Bracket expressions are used to match ranges of characters. Bracket expressions where the range is backward, for example ‘[z-a]’, are invalid. Within square brackets, ‘\’ is taken literally. Character classes are supported; for example ‘[[:digit:]]’ will match a single decimal digit.

GNU extensions are supported:

1. ‘\w’ matches a character within a word

2. ‘\W’ matches a character which is not within a word

3. ‘\<’ matches the beginning of a word

4. ‘\>’ matches the end of a word

5. ‘\b’ matches a word boundary

6. ‘\B’ matches characters which are not a word boundary

7. ‘\` ’ matches the beginning of the whole input

8. ‘\'’ matches the end of the whole input

Grouping is performed with backslashes followed by parentheses ‘\(’, ‘\)’. A backslash followed by a digit acts as a back-reference and matches the same thing as the previous grouped expression indicated by that number. For example ‘\2’ matches the second group expression. The order of group expressions is determined by the position of their opening parenthesis ‘\(’.

The alternation operator is ‘\|’.

The character ‘^’ only represents the beginning of a string when it appears:

1. At the beginning of a regular expression

2. After an open-group, signified by ‘\(’

3. After a newline

4. After the alternation operator ‘\|’

The character ‘$’ only represents the end of a string when it appears:

1. At the end of a regular expression

2. Before a close-group, signified by ‘\)’

3. Before a newline

4. Before the alternation operator ‘\|’

‘\*’, ‘\+’ and ‘\?’ are special at any point in a regular expression except:

1. At the beginning of a regular expression

2. After an open-group, signified by ‘\(’

3. After a newline

4. After the alternation operator ‘\|’

Intervals are specified by ‘\{’ and ‘\}’. Invalid intervals such as ‘a\{1z’ are not accepted. The longest possible match is returned; this applies to the regular expression as a whole

and (subject to this constraint) to subexpressions within groups.

8.5.5 ‘posix-awk’ regular expression syntax

The character ‘.’ matches any single character except the null character.

‘+’ indicates that the regular expression should match one or more occurrences of the previous atom or regexp.

‘?’ indicates that the regular expression should match zero or one occurrence of the previous atom or regexp.

‘\+’ matches a ‘+’

‘\?’ matches a ‘?’.

Bracket expressions are used to match ranges of characters. Bracket expressions where the range is backward, for example ‘[z-a]’, are invalid. Within square brackets, ‘\’ can be used to quote the following character. Character classes are supported; for example ‘[[:digit:]]’ will match a single decimal digit.

GNU extensions are not supported and so ‘\w’, ‘\W’, ‘\<’, ‘\>’, ‘\b’, ‘\B’, ‘\` ’, and ‘\'’ match ‘w’, ‘W’, ‘<’, ‘>’, ‘b’, ‘B’, ‘` ’, and ‘'’ respectively.

Grouping is performed with parentheses ‘()’. An unmatched ‘)’ matches just itself. A backslash followed by a digit acts as a back-reference and matches the same thing as the previous grouped expression indicated by that number. For example ‘\2’ matches the second group expression. The order of group expressions is determined by the position of their opening parenthesis ‘(’.

The alternation operator is ‘|’.

The characters ‘^’ and ‘$’ always represent the beginning and end of a string respectively, except within square brackets. Within brackets, ‘^’ can be used to invert the membership of the character class being specified.

‘*’, ‘+’ and ‘?’ are special at any point in a regular expression except the following places, where they are not allowed:

1. At the beginning of a regular expression

2. After an open-group, signified by ‘(’

3. After the alternation operator ‘|’

Intervals are specified by ‘{’ and ‘}’. Invalid intervals are treated as literals, for example ‘a{1’ is treated as ‘a\{1’

The longest possible match is returned; this applies to the regular expression as a whole and (subject to this constraint) to subexpressions within groups.

8.5.6 ‘awk’ regular expression syntax

The character ‘.’ matches any single character except the null character.

‘+’ indicates that the regular expression should match one or more occurrences of the previous atom or regexp.

‘?’ indicates that the regular expression should match zero or one occurrence of the previous atom or regexp.

‘\+’ matches a ‘+’

‘\?’ matches a ‘?’.

Bracket expressions are used to match ranges of characters. Bracket expressions where the range is backward, for example ‘[z-a]’, are invalid. Within square brackets, ‘\’ can be used to quote the following character. Character classes are supported; for example ‘[[:digit:]]’ will match a single decimal digit.

GNU extensions are not supported and so ‘\w’, ‘\W’, ‘\<’, ‘\>’, ‘\b’, ‘\B’, ‘\` ’, and ‘\'’ match ‘w’, ‘W’, ‘<’, ‘>’, ‘b’, ‘B’, ‘` ’, and ‘'’ respectively.

Grouping is performed with parentheses ‘()’. An unmatched ‘)’ matches just itself. A backslash followed by a digit matches that digit.

The alternation operator is ‘|’.

The characters ‘^’ and ‘$’ always represent the beginning and end of a string respectively, except within square brackets. Within brackets, ‘^’ can be used to invert the membership of the character class being specified.

‘*’, ‘+’ and ‘?’ are special at any point in a regular expression except:

1. At the beginning of a regular expression

2. After an open-group, signified by ‘(’

3. After the alternation operator ‘|’

The longest possible match is returned; this applies to the regular expression as a whole and (subject to this constraint) to subexpressions within groups.

8.5.7 ‘posix-basic’ regular expression syntax

The character ‘.’ matches any single character except the null character. ‘\+’ indicates that the regular expression should match one or more occurrences of

the previous atom or regexp.

‘\?’ indicates that the regular expression should match zero or one occurrence of

the previous atom or regexp.

‘+ and ?’ match themselves.

Bracket expressions are used to match ranges of characters. Bracket expressions where the range is backward, for example ‘[z-a]’, are invalid. Within square brackets, ‘\’ is taken literally. Character classes are supported; for example ‘[[:digit:]]’ will match a single decimal digit.

GNU extensions are supported:

1. ‘\w’ matches a character within a word

2. ‘\W’ matches a character which is not within a word

3. ‘\<’ matches the beginning of a word

4. ‘\>’ matches the end of a word

5. ‘\b’ matches a word boundary

6. ‘\B’ matches characters which are not a word boundary

7. ‘\` ’ matches the beginning of the whole input

8. ‘\'’ matches the end of the whole input Grouping is performed with backslashes followed by parentheses ‘\(’, ‘\)’. A backslash

followed by a digit acts as a back-reference and matches the same thing as the previous grouped expression indicated by that number. For example ‘\2’ matches the second group expression. The order of group expressions is determined by the position of their opening parenthesis ‘\(’.

The alternation operator is ‘\|’.

The character ‘^’ only represents the beginning of a string when it appears:

1. At the beginning of a regular expression

2. After an open-group, signified by ‘\(’

3. After the alternation operator ‘\|’ The character ‘$’ only represents the end of a string when it appears:

1. At the end of a regular expression

2. Before a close-group, signified by ‘\)’

3. Before the alternation operator ‘\|’ ‘\*’, ‘\+’ and ‘\?’ are special at any point in a regular expression except:

1. At the beginning of a regular expression 2. After an open-group, signified by ‘\(’

3. After the alternation operator ‘\|’ Intervals are specified by ‘\{’ and ‘\}’. Invalid intervals such as ‘a\{1z’ are not accepted. The longest possible match is returned; this applies to the regular expression as a whole

and (subject to this constraint) to subexpressions within groups.

8.5.8 ‘posix-egrep’ regular expression syntax

The character ‘.’ matches any single character. ‘+’ indicates that the regular expression should match one or more occurrences of

the previous atom or regexp.

‘?’ indicates that the regular expression should match zero or one occurrence of

the previous atom or regexp.

‘\+’ matches a ‘+’ ‘\?’ matches a ‘?’.

Bracket expressions are used to match ranges of characters. Bracket expressions where the range is backward, for example ‘[z-a]’, are invalid. Within square brackets, ‘\’ is taken literally. Character classes are supported; for example ‘[[:digit:]]’ will match a single decimal digit.

GNU extensions are supported:

1. ‘\w’ matches a character within a word

2. ‘\W’ matches a character which is not within a word

3. ‘\<’ matches the beginning of a word

4. ‘\>’ matches the end of a word

5. ‘\b’ matches a word boundary

6. ‘\B’ matches characters which are not a word boundary

7. ‘\` ’ matches the beginning of the whole input

8. ‘\'’ matches the end of the whole input Grouping is performed with parentheses ‘()’. An unmatched ‘)’ matches just itself.

A backslash followed by a digit acts as a back-reference and matches the same thing as the previous grouped expression indicated by that number. For example ‘\2’ matches the second group expression. The order of group expressions is determined by the position of their opening parenthesis ‘(’.

The alternation operator is ‘|’.

The characters ‘^’ and ‘$’ always represent the beginning and end of a string respectively, except within square brackets. Within brackets, ‘^’ can be used to invert the membership of the character class being specified.

The characters ‘*’, ‘+’ and ‘?’ are special anywhere in a regular expression.

Intervals are specified by ‘{’ and ‘}’. Invalid intervals are treated as literals, for example ‘a{1’ is treated as ‘a\{1’

The longest possible match is returned; this applies to the regular expression as a whole and (subject to this constraint) to subexpressions within groups.

8.5.9 ‘egrep’ regular expression syntax

This is a synonym for posix-egrep.

8.5.10 ‘posix-extended’ regular expression syntax

The character ‘.’ matches any single character except the null character. ‘+’ indicates that the regular expression should match one or more occurrences of

the previous atom or regexp.

‘?’ indicates that the regular expression should match zero or one occurrence of

the previous atom or regexp.

‘\+’ matches a ‘+’ ‘\?’ matches a ‘?’.

Bracket expressions are used to match ranges of characters. Bracket expressions where the range is backward, for example ‘[z-a]’, are invalid. Within square brackets, ‘\’ is taken literally. Character classes are supported; for example ‘[[:digit:]]’ will match a single decimal digit.

GNU extensions are supported:

1. ‘\w’ matches a character within a word

2. ‘\W’ matches a character which is not within a word

3. ‘\<’ matches the beginning of a word

4. ‘\>’ matches the end of a word

5. ‘\b’ matches a word boundary

6. ‘\B’ matches characters which are not a word boundary

7. ‘\` ’ matches the beginning of the whole input

8. ‘\'’ matches the end of the whole input Grouping is performed with parentheses ‘()’. An unmatched ‘)’ matches just itself.

A backslash followed by a digit acts as a back-reference and matches the same thing as the previous grouped expression indicated by that number. For example ‘\2’ matches the second group expression. The order of group expressions is determined by the position of their opening parenthesis ‘(’.

The alternation operator is ‘|’.

The characters ‘^’ and ‘$’ always represent the beginning and end of a string respectively, except within square brackets. Within brackets, ‘^’ can be used to invert the membership of the character class being specified.

‘*’, ‘+’ and ‘?’ are special at any point in a regular expression except the following places, where they are not allowed:

1. At the beginning of a regular expression

2. After an open-group, signified by ‘(’

3. After the alternation operator ‘|’ Intervals are specified by ‘{’ and ‘}’. Invalid intervals such as ‘a{1z’ are not accepted. The longest possible match is returned; this applies to the regular expression as a whole

and (subject to this constraint) to subexpressions within groups.

8.6 Environment Variables

LANG Provides a default value for the internationalisation variables that are unset or null.

LC_ALL If set to a non-empty string value, override the values of all the other internationalisation variables.

LC_COLLATE

The POSIX standard specifies that this variable affects the pattern matching to be used for the ‘-name’ option. GNU find uses the GNU version of the fnmatch library function. This variable also affects the interpretation of the response to -ok; while the LC_MESSAGES variable selects the actual pattern used to interpret the response to -ok, the interpretation of any bracket expressions in the pattern will be affected by the LC_COLLATE variable.

LC_CTYPE This variable affects the treatment of character classes used in regular expression and with the ‘-name’ test, if the fnmatch function supports this. This variable also affects the interpretation of any character classes in the regular expressions used to interpret the response to the prompt issued by -ok. The LC_CTYPE environment variable will also affect which characters are considered to be unprintable when filenames are printed (see Section 3.3.2.3 [Unusual Characters in File Names], page 33).

LC_MESSAGES

Determines the locale to be used for internationalised messages, including the interpretation of the response to the prompt made by the -ok action.

NLSPATH Determines the location of the internationalisation message catalogues.

PATH Affects the directories which are searched to find the executables invoked by ‘-exec’, ‘-execdir’ ‘-ok’ and ‘-okdir’. If the PATH environment variable includes the current directory (by explicitly including ‘.’ or by having an empty element), and the find command line includes ‘-execdir’ or ‘-okdir’, find will refuse to run. See Chapter 11 [Security Considerations], page 93, for a more detailed discussion of security matters.

POSIXLY_CORRECT

Determines the block size used by ‘-ls’ and ‘-fls’. If POSIXLY_CORRECT is set, blocks are units of 512 bytes. Otherwise they are units of 1024 bytes. Setting this variable also turns off warning messages (that is, implies ‘-nowarn’) by default, because POSIX requires that apart from the output for ‘-ok’, all messages printed on stderr are diagnostics and must result in a non-zero exit status. When POSIXLY_CORRECT is set, the response to the prompt made by the -ok action is interpreted according to the system’s message catalogue, as opposed to according to find’s own message translations.

TZ Affects the time zone used for some of the time-related format directives of

‘-printf’ and ‘-fprintf’.

9 Common Tasks

The sections that follow contain some extended examples that both give a good idea of the power of these programs, and show you how to solve common real-world problems.

9.1 Viewing And Editing

To view a list of files that meet certain criteria, simply run your file viewing program with the file names as arguments. Shells substitute a command enclosed in backquotes with its output, so the whole command looks like this:

     less ` find /usr/include -name '*.h' | xargs grep -l mode_t`

You can edit those files by giving an editor name instead of a file viewing program:

     emacs ` find /usr/include -name '*.h' | xargs grep -l mode_t`

Because there is a limit to the length of any individual command line, there is a limit to the number of files that can be handled in this way. We can get around this difficulty by using xargs like this:

     find /usr/include -name '*.h' | xargs grep -l mode_t > todo
     xargs --arg-file=todo emacs

Here, xargs will run emacs as many times as necessary to visit all of the files listed in the file todo. Generating a temporary file is not always convenient, though. This command does much the same thing without needing one:

     find /usr/include -name '*.h' | xargs grep -l mode_t |
     xargs sh -c 'emacs "$@" < /dev/tty' Emacs

The example above illustrates a useful trick; Using sh -c you can invoke a shell command from xargs. The $@ in the command line is expanded by the shell to a list of arguments as provided by xargs. The single quotes in the command line protect the $@ against expansion by your interactive shell (which will normally have no arguments and thus expand $@ to nothing). The capitalised ‘Emacs’ on the command line is used as $0 by the shell that xargs launches.

Please note that the implementations in GNU xargs and at least BSD support the ‘-o’ option as extension to achieve the same, while the above is the portable way to redirect

stdin to /dev/tty.

9.2 Archiving

You can pass a list of files produced by find to a file archiving program. GNU tar and cpio can both read lists of file names from the standard input – either delimited by nulls (the safe way) or by blanks (the lazy, risky default way). To use null-delimited names, give them the ‘--null’ option. You can store a file archive in a file, write it on a tape, or send it over a network to extract on another machine.

One common use of find to archive files is to send a list of the files in a directory tree to cpio. Use ‘-depth’ so if a directory does not have write permission for its owner, its contents can still be restored from the archive since the directory’s permissions are restored after its contents. Here is an example of doing this using cpio; you could use a more complex find expression to archive only certain files.

     find . -depth -print0 |

       cpio --create --null --format=crc --file=/dev/nrst0

You could restore that archive using this command:

     cpio --extract --null --make-dir --unconditional \
       --preserve --file=/dev/nrst0

Here are the commands to do the same things using tar:

     find . -depth -print0 |
       tar --create --null --files-from=- --file=/dev/nrst0
     tar --extract --null --preserve-perm --same-owner \
       --file=/dev/nrst0

Here is an example of copying a directory from one machine to another:

     find . -depth -print0 | cpio -0o -Hnewc |
       rsh other-machine "cd ` pwd` && cpio -i0dum"

9.3 Cleaning Up

This section gives examples of removing unwanted files in various situations. Here is a command to remove the CVS backup files created when an update requires a merge:

     find . -name '.#*' -print0 | xargs -0r rm -f

If your find command removes directories, you may find that you get a spurious error message when find tries to recurse into a directory that has now been removed. Using the ‘-depth’ option will normally resolve this problem.

It is also possible to use the ‘-delete’ action:

     find . -depth -name '.#*' -delete

You can run this command to clean out your clutter in /tmp. You might place it in the file your shell runs when you log out (.bash_logout, .logout, or .zlogout, depending on which shell you use).

     find /tmp -depth -user "$LOGNAME" -type f -delete

To remove old Emacs backup and auto-save files, you can use a command like the following. It is especially important in this case to use null-terminated file names because Emacs packages like the VM mailer often create temporary file names with spaces in them,

like #reply to David J. MacKenzie<1>#.
     find ~ \( -name '*~' -o -name '#*#' \) -print0 |
       xargs --no-run-if-empty --null rm -vf

Removing old files from /tmp is commonly done from cron:

     find /tmp /var/tmp -depth -not        -type d -mtime +3 -delete
     find /tmp /var/tmp -depth -mindepth 1 -type d -empty    -delete

The second find command above cleans out empty directories depth-first (‘-delete’ implies ‘-depth’ anyway), hoping that the parents become empty and can be removed too. It uses ‘-mindepth’ to avoid removing /tmp itself if it becomes totally empty.

Lastly, an example of a program that almost certainly does not do what the user intended:

     find dirname -delete -name quux

If the user hoped to delete only files named quux they will get an unpleasant surprise; this command will attempt to delete everything at or below the starting point dirname. This is because find evaluates the items on the command line as an expression. The find program will normally execute an action if the preceding action succeeds. Here, there is no action or test before the ‘-delete’ so it will always be executed. The ‘-name quux’ test will be performed for files we successfully deleted, but that test has no effect since ‘-delete’ also disables the default ‘-print’ operation. So the above example will probably delete a lot of files the user didn’t want to delete.

This command is also likely to do something you did not intend:

     find dirname -path dirname/foo -prune -o -delete

Because ‘-delete’ turns on ‘-depth’, the ‘-prune’ action has no effect and files in dirname/foo will be deleted too.

9.4 Strange File Names

find can help you remove or rename a file with strange characters in its name. People are sometimes stymied by files whose names contain characters such as spaces, tabs, control characters, or characters with the high bit set. The simplest way to remove such files is:

     rm -i some*pattern*that*matches*the*problem*file

rm asks you whether to remove each file matching the given pattern. If you are using an old shell, this approach might not work if the file name contains a character with the high bit set; the shell may strip it off. A more reliable way is:

     find . -maxdepth 1 tests -okdir rm '{}' \;

where tests uniquely identify the file. The ‘-maxdepth 1’ option prevents find from wasting time searching for the file in any subdirectories; if there are no subdirectories, you may omit it. A good way to uniquely identify the problem file is to figure out its inode number; use

     ls -i

Suppose you have a file whose name contains control characters, and you have found that its inode number is 12345. This command prompts you for whether to remove it:

     find . -maxdepth 1 -inum 12345 -okdir rm -f '{}' \;

If you don’t want to be asked, perhaps because the file name may contain a strange character sequence that will mess up your screen when printed, then use ‘-execdir’ instead

of ‘-okdir’.

If you want to rename the file instead, you can use mv instead of rm:

     find . -maxdepth 1 -inum 12345 -okdir mv '{}' new-file-name \;

9.5 Fixing Permissions

Suppose you want to make sure that everyone can write to the directories in a certain directory tree. Here is a way to find directories lacking either user or group write permission (or both), and fix their permissions:

     find . -type d -not -perm -ug=w | xargs chmod ug+w

You could also reverse the operations, if you want to make sure that directories do not have world write permission.

9.6 Classifying Files

If you want to classify a set of files into several groups based on different criteria, you can use the comma operator to perform multiple independent tests on the files. Here is an example:

find / -type d \( -perm -o=w -fprint allwrite , \
  -perm -o=x -fprint allexec \)
echo "Directories that can be written to by everyone:"
cat allwrite
echo ""
echo "Directories with search permissions for everyone:"
cat allexec

find has only to make one scan through the directory tree (which is one of the most time consuming parts of its work).

10 Worked Examples

The tools in the findutils package, and in particular find, have a large number of options. This means that quite often, there is more than one way to do things. Some of the options and facilities only exist for compatibility with other tools, and findutils provides improved ways of doing things.

This chapter describes a number of useful tasks that are commonly performed, and compares the different ways of achieving them.

10.1 Deleting Files

One of the most common tasks that find is used for is locating files that can be deleted. This might include:

• Files last modified more than 3 years ago which haven’t been accessed for at least 2 years

• Files belonging to a certain user

• Temporary files which are no longer required

This example concentrates on the actual deletion task rather than on sophisticated ways of locating the files that need to be deleted. We’ll assume that the files we want to delete are old files underneath /var/tmp/stuff.

10.1.1 The Traditional Way

The traditional way to delete files in /var/tmp/stuff that have not been modified in over 90 days would have been:

      find /var/tmp/stuff -mtime +90 -exec /bin/rm {} \;

The above command uses ‘-exec’ to run the /bin/rm command to remove each file. This approach works and in fact would have worked in Version 7 Unix in 1979. However, there are a number of problems with this approach.

The most obvious problem with the approach above is that it causes find to fork every time it finds a file that needs to delete, and the child process then has to use the exec system call to launch /bin/rm. All this is quite inefficient. If we are going to use /bin/rm to do this job, it is better to make it delete more than one file at a time.

The most obvious way of doing this is to use the shell’s command expansion feature:

      /bin/rm ` find /var/tmp/stuff -mtime +90 -print`

or you could use the more modern form

      /bin/rm $(find /var/tmp/stuff -mtime +90 -print)

The commands above are much more efficient than the first attempt. However, there is a problem with them. The shell has a maximum command length which is imposed by the operating system (the actual limit varies between systems). This means that while the command expansion technique will usually work, it will suddenly fail when there are lots of files to delete. Since the task is to delete unwanted files, this is precisely the time we don’t want things to go wrong.

10.1.2 Making Use of xargs

So, is there a way to be more efficient in the use of fork() and exec() without running up against this limit? Yes, we can be almost optimally efficient by making use of the xargs command. The xargs command reads arguments from its standard input and builds them into command lines. We can use it like this:

      find /var/tmp/stuff -mtime +90 -print | xargs /bin/rm
   For example if the files found by find are /var/tmp/stuff/A, /var/tmp/stuff/B and

/var/tmp/stuff/C then xargs might issue the commands

      /bin/rm /var/tmp/stuff/A /var/tmp/stuff/B
      /bin/rm /var/tmp/stuff/C

The above assumes that xargs has a very small maximum command line length. The real limit is much larger but the idea is that xargs will run /bin/rm as many times as necessary to get the job done, given the limits on command line length.

This usage of xargs is pretty efficient, and the xargs command is widely implemented (all modern versions of Unix offer it). So far then, the news is all good. However, there is bad news too.

10.1.3 Unusual characters in filenames

Unix-like systems allow any characters to appear in file names with the exception of the ASCII NUL character and the slash. Slashes can occur in path names (as the directory separator) but not in the names of actual directory entries. This means that the list of files that xargs reads could in fact contain white space characters – spaces, tabs and newline characters. Since by default, xargs assumes that the list of files it is reading uses white space as an argument separator, it cannot correctly handle the case where a filename actually includes white space. This makes the default behaviour of xargs almost useless for handling arbitrary data.

To solve this problem, GNU findutils introduced the ‘-print0’ action for find. This uses the ASCII NUL character to separate the entries in the file list that it produces. This is the ideal choice of separator since it is the only character that cannot appear within a path name. The ‘-0’ option to xargs makes it assume that arguments are separated with ASCII NUL instead of white space. It also turns off another misfeature in the default behaviour of xargs, which is that it pays attention to quote characters in its input. Some versions of xargs also terminate when they see a lone ‘_’ in the input, but GNU find no longer does that (since it has become an optional behaviour in the Unix standard).

So, putting find -print0 together with xargs -0 we get this command:

      find /var/tmp/stuff -mtime +90 -print0 | xargs -0 /bin/rm

The result is an efficient way of proceeding that correctly handles all the possible characters that could appear in the list of files to delete. This is good news. However, there is, as I’m sure you’re expecting, also more bad news. The problem is that this is not a portable construct; although other versions of Unix (notably BSD-derived ones) support ‘-print0’, it’s not universal. So, is there a more universal mechanism?

10.1.4 Going back to -exec

There is indeed a more universal mechanism, which is a slight modification to the ‘-exec’ action. The normal ‘-exec’ action assumes that the command to run is terminated with a semicolon (the semicolon normally has to be quoted in order to protect it from interpretation as the shell command separator). The SVR4 edition of Unix introduced a slight variation, which involves terminating the command with ‘+’ instead:

      find /var/tmp/stuff -mtime +90 -exec /bin/rm {} \+

The above use of ‘-exec’ causes find to build up a long command line and then issue it. This can be less efficient than some uses of xargs; for example xargs allows building up new command lines while the previous command is still executing, and allows specifying a number of commands to run in parallel. However, the find ... -exec ... + construct has the advantage of wide portability. GNU findutils did not support ‘-exec ... +’ until version 4.2.12; one of the reasons for this is that it already had the ‘-print0’ action in any case.

10.1.5 A more secure version of -exec

The command above seems to be efficient and portable. However, within it lurks a security problem. The problem is shared with all the commands we’ve tried in this worked example so far, too. The security problem is a race condition; that is, if it is possible for somebody to manipulate the filesystem that you are searching while you are searching it, it is possible for them to persuade your find command to cause the deletion of a file that you can delete but they normally cannot.

The problem occurs because the ‘-exec’ action is defined by the POSIX standard to invoke its command with the same working directory as find had when it was started. This means that the arguments which replace the {} include a relative path from find’s starting point down the file that needs to be deleted. For example,

      find /var/tmp/stuff -mtime +90 -exec /bin/rm {} \+

might actually issue the command:

      /bin/rm /var/tmp/stuff/A /var/tmp/stuff/B /var/tmp/stuff/passwd

Notice the file /var/tmp/stuff/passwd. Likewise, the command:

      cd /var/tmp && find stuff -mtime +90 -exec /bin/rm {} \+

might actually issue the command:

      /bin/rm stuff/A stuff/B stuff/passwd

If an attacker can rename stuff to something else (making use of their write permissions in /var/tmp) they can replace it with a symbolic link to /etc. That means that the /bin/rm command will be invoked on /etc/passwd. If you are running your find command as root, the attacker has just managed to delete a vital file. All they needed to do to achieve this was replace a subdirectory with a symbolic link at the vital moment.

There is however, a simple solution to the problem. This is an action which works a lot like -exec but doesn’t need to traverse a chain of directories to reach the file that it needs to work on. This is the ‘-execdir’ action, which was introduced by the BSD family of operating systems. The command,

      find /var/tmp/stuff -mtime +90 -execdir /bin/rm {} \+

might delete a set of files by performing these actions:

1. Change directory to /var/tmp/stuff/foo

 2. Invoke /bin/rm ./file1 ./file2 ./file3

3. Change directory to /var/tmp/stuff/bar

 4. Invoke /bin/rm ./file99 ./file100 ./file101

This is a much more secure method. We are no longer exposed to a race condition. For many typical uses of find, this is the best strategy. It’s reasonably efficient, but the length of the command line is limited not just by the operating system limits, but also by how many files we actually need to delete from each directory.

Is it possible to do any better? In the case of general file processing, no. However, in the specific case of deleting files it is indeed possible to do better.

10.1.6 Using the -delete action

The most efficient and secure method of solving this problem is to use the ‘-delete’ action:

      find /var/tmp/stuff -mtime +90 -delete

This alternative is more efficient than any of the ‘-exec’ or ‘-execdir’ actions, since it entirely avoids the overhead of forking a new process and using exec to run /bin/rm. It is also normally more efficient than xargs for the same reason. The file deletion is performed from the directory containing the entry to be deleted, so the ‘-delete’ action has the same security advantages as the ‘-execdir’ action has.

The ‘-delete’ action was introduced by the BSD family of operating systems.

10.1.7 Improving things still further

Is it possible to improve things still further? Not without either modifying the system library to the operating system or having more specific knowledge of the layout of the filesystem and disk I/O subsystem, or both.

The find command traverses the filesystem, reading directories. It then issues a separate system call for each file to be deleted. If we could modify the operating system, there are potential gains that could be made:

• We could have a system call to which we pass more than one filename for deletion • Alternatively, we could pass in a list of inode numbers (on GNU/Linux systems,

readdir() also returns the inode number of each directory entry) to be deleted.

The above possibilities sound interesting, but from the kernel’s point of view it is difficult to enforce standard Unix access controls for such processing by inode number. Such a facility would probably need to be restricted to the superuser.

Another way of improving performance would be to increase the parallelism of the process. For example if the directory hierarchy we are searching is actually spread across a number of disks, we might somehow be able to arrange for find to process each disk in parallel. In practice GNU find doesn’t have such an intimate understanding of the system’s filesystem layout and disk I/O subsystem.

However, since the system administrator can have such an understanding they can take advantage of it like so:

      find /var/tmp/stuff1 -mtime +90 -delete &
      find /var/tmp/stuff2 -mtime +90 -delete &
      find /var/tmp/stuff3 -mtime +90 -delete &
      find /var/tmp/stuff4 -mtime +90 -delete &
      wait

In the example above, four separate instances of find are used to search four subdirectories in parallel. The wait command simply waits for all of these to complete. Whether this approach is more or less efficient than a single instance of find depends on a number of things:

• Are the directories being searched in parallel actually on separate disks? If not, this parallel search might just result in a lot of disk head movement and so the speed might even be slower.

• Other activity - are other programs also doing things on those disks?

10.1.8 Conclusion

The fastest and most secure way to delete files with the help of find is to use ‘-delete’. Using xargs -0 -P N can also make effective use of the disk, but it is not as secure.

In the case where we’re doing things other than deleting files, the most secure alternative is ‘-execdir ... +’, but this is not as portable as the insecure action ‘-exec ... +’.

The ‘-delete’ action is not completely portable, but the only other possibility which is as secure (‘-execdir’) is no more portable. The most efficient portable alternative is ‘-exec ...+’, but this is insecure and isn’t supported by versions of GNU findutils prior to 4.2.12.

10.2 Copying A Subset of Files

Suppose you want to copy some files from /source-dir to /dest-dir, but there are a small number of files in /source-dir you don’t want to copy.

One option of course is cp /source-dir /dest-dir followed by deletion of the unwanted material under /dest-dir. But often that can be inconvenient, because for example we would have copied a large amount of extraneous material, or because /dest-dir is too small. Naturally there are many other possible reasons why this strategy may be unsuitable.

So we need to have some way of identifying which files we want to copy, and we need to have a way of copying that file list. The second part of this condition is met by cpio -p. Of course, we can identify the files we wish to copy by using find. Here is a command that solves our problem:

     cd /source-dir
     find . -name '.snapshot' -prune -o \( \! -name '*~' -print0 \) |
     cpio -pmd0   /dest-dir

The first part of the find command here identifies files or directories named .snapshot and tells find not to recurse into them (since they do not need to be copied). The combination -name '.snapshot' -prune yields false for anything that didn’t get pruned, but it is exactly those files we want to copy. Therefore we need to use an OR (‘-o’) condition to introduce the rest of our expression. The remainder of the expression simply arranges for the name of any file not ending in ‘~’ to be printed.

Using -print0 ensures that white space characters in file names do not pose a problem. The cpio command does the actual work of copying files. The program as a whole fails if the cpio program returns nonzero. If the find command returns non-zero on the other hand, the Unix shell will not diagnose a problem (since find is not the last command in the pipeline).

10.3 Updating A Timestamp File

Suppose we have a directory full of files which is maintained with a set of automated tools; perhaps one set of tools updates them and another set of tools uses the result. In this situation, it might be useful for the second set of tools to know if the files have recently been changed. It might be useful, for example, to have a ’timestamp’ file which gives the timestamp on the newest file in the collection.

We can use find to achieve this, but there are several different ways to do it.

10.3.1 Updating the Timestamp The Wrong Way

The obvious but wrong answer is just to use ‘-newer’:

      find subdir -newer timestamp -exec touch -r {} timestamp \;

This does the right sort of thing but has a bug. Suppose that two files in the subdirectory have been updated, and that these are called file1 and file2. The command above will update timestamp with the modification time of file1 or that of file2, but we don’t know which one. Since the timestamps on file1 and file2 will in general be different, this could well be the wrong value.

One solution to this problem is to modify find to recheck the modification time of timestamp every time a file is to be compared against it, but that will reduce the performance of find.

10.3.2 Using the test utility to compare timestamps

The test command can be used to compare timestamps:

      find subdir -exec test {} -nt timestamp \; -exec touch -r {} timestamp \;

This will ensure that any changes made to the modification time of timestamp that take place during the execution of find are taken into account. This resolves our earlier problem, but unfortunately this runs much more slowly.

10.3.3 A combined approach

We can of course still use ‘-newer’ to cut down on the number of calls to test:

      find subdir -newer timestamp -and \
           -exec test {} -nt timestamp \; -and \
           -exec touch -r {} timestamp \;

Here, the ‘-newer’ test excludes all the files which are definitely older than the timestamp, but all the files which are newer than the old value of the timestamp are compared against the current updated timestamp.

This is indeed faster in general, but the speed difference will depend on how many updated files there are.

10.3.4 Using -printf and sort to compare timestamps

It is possible to use the ‘-printf’ action to abandon the use of test entirely:

      newest=$(find subdir -newer timestamp -printf "%A@:%p\n" |
                 sort -n |
                 tail -n1 |
                 cut -d: -f2- )
      touch -r "${newest:-timestamp}" timestamp

The command above works by generating a list of the timestamps and names of all the files which are newer than the timestamp. The sort, tail and cut commands simply pull out the name of the file with the largest timestamp value (that is, the latest file). The touch command is then used to update the timestamp,

The "${newest:-timestamp}" expression simply expands to the value of $newest if that variable is set, but to timestamp otherwise. This ensures that an argument is always given to the ‘-r’ option of the touch command.

This approach seems quite efficient, but unfortunately it has a problem. Many operating systems now keep file modification time information at a granularity which is finer than one second. Findutils version 4.3.3 and later will print a fractional part with %A@, but older versions will not.

10.3.5 Solving the problem with make

Another tool which often works with timestamps is make. We can use find to generate a Makefile file on the fly and then use make to update the timestamps:

      makefile=$(mktemp)
      find subdir \
              \( \! -xtype l \) \
              -newer timestamp \
              -printf "timestamp:: %p\n\ttouch -r %p timestamp\n\n" > "$makefile"
      make -f "$makefile"
      rm   -f "$makefile"

Unfortunately although the solution above is quite elegant, it fails to cope with white space within file names, and adjusting it to do so would require a rather complex shell script.

10.3.6 Coping with odd filenames too

We can fix both of these problems (looping and problems with white space), and do things more efficiently too. The following command works with newlines and doesn’t need to sort the list of filenames.

      find subdir -newer timestamp -printf "%A@:%p\0" |
         perl -0 newest.pl |
         xargs --no-run-if-empty --null --replace \
            find {} -maxdepth 0 -newer timestamp -exec touch -r {} timestamp \;

The first find command generates a list of files which are newer than the original timestamp file, and prints a list of them with their timestamps. The newest.pl script simply filters out all the filenames which have timestamps which are older than whatever the newest file is:

      #! /usr/bin/perl -0
      my @newest = ();
      my $latest_stamp = undef;
      while (<>) {
          my ($stamp, $name) = split(/:/);
          if (!defined($latest_stamp) || ($tstamp > $latest_stamp)) {
              $latest_stamp = $stamp;
              @newest = ();
          }
          if ($tstamp >= $latest_stamp) {
              push @newest, $name;
          }

      }
      print join("\0", @newest);

This prints a list of zero or more files, all of which are newer than the original timestamp file, and which have the same timestamp as each other, to the nearest second. The second find command takes each resulting file one at a time, and if that is newer than the timestamp file, the timestamp is updated.

10.4 Finding the Shallowest Instance

Suppose you maintain local copies of sources from various projects, each with their own choice of directory organisation and source code management (SCM) tool. You need to periodically synchronize each project with its upstream tree. As the number local repositories grows, so does the work involved in maintaining synchronization. SCM utilities typically create some sort of administrative directory: .svn for Subversion, CVS for CVS, and so on. These directories can be used as a key to search for the bases of the project source trees. Suppose we have the following directory structure:

      repo/project1/CVS
      repo/gnu/project2/.svn
      repo/gnu/project3/.svn
      repo/gnu/project3/src/.svn
      repo/gnu/project3/doc/.svn
      repo/project4/.git

One would expect to update each of the projectX directories, but not their subdirectories (src, doc, etc.). To locate the project roots, we would need to find the least deeply nested directories containing an SCM-related subdirectory. The following command discovers those roots efficiently. It is efficient because it avoids searching subdirectories inside projects whose SCM directory we already found.

      find repo/ \
      -exec test -d {}/.svn \; -or \
      -exec test -d {}/.git \; -or \
      -exec test -d {}/CVS \; -print -prune

In this example, test is used to tell if we are currently examining a directory which appears to the a project’s root directory (because it has an SCM subdirectory). When we find a project root, there is no need to search inside it, and -prune makes sure that we descend no further.

For large, complex trees like the Linux kernel, this will prevent searching a large portion of the structure, saving a good deal of time.

11 Security Considerations

Security considerations are important if you are using find or xargs to search for or process files that don’t belong to you or which other people have control. Security considerations relating to locate may also apply if you have files which you do not want others to see.

The most severe forms of security problems affecting find and related programs are when third parties bring about a situation allowing them to do something they would normally not be able to accomplish. This is called privilege elevation. This might include deleting files they would not normally be able to delete. It is common for the operating system to periodically invoke find for self-maintenance purposes. These invocations of find are particularly problematic from a security point of view as these are often invoked by the superuser and search the entire filesystem hierarchy. Generally, the severity of any associated problem depends on what the system is going to do with the files found by find.

11.1 Levels of Risk

There are some security risks inherent in the use of find, xargs and (to a lesser extent) locate. The severity of these risks depends on what sort of system you are using:

High risk Multi-user systems where you do not control (or trust) the other users, and on which you execute find, including areas where those other users can manipulate the filesystem (for example beneath /home or /tmp).

Medium Risk Systems where the actions of other users can create file names chosen by them, but to which they don’t have access while find is being run. This access might include leaving programs running (shell background jobs, at or cron tasks, for example). On these sorts of systems, carefully written commands (avoiding use of ‘-print’ for example) should not expose you to a high degree of risk. Most systems fall into this category.

Low Risk Systems to which untrusted parties do not have access, cannot create file names of their own choice (even remotely) and which contain no security flaws which might enable an untrusted third party to gain access. Most systems do not fall into this category because there are many ways in which external parties can affect the names of files that are created on your system. The system on which I am writing this for example automatically downloads software updates from the Internet; the names of the files in which these updates exist are chosen by third parties^{1}.

In the discussion above, “risk” denotes the likelihood that someone can cause find, xargs, locate or some other program which is controlled by them to do something you did not intend. The levels of risk suggested do not take any account of the consequences of this sort of event. That is, if you operate a “low risk” type system, but the consequences of a security problem are disastrous, then you should still give serious thought to all the possible security problems, many of which of course will not be discussed here – this section of the manual is intended to be informative but not comprehensive or exhaustive.

1 Of course, I trust these parties to a large extent anyway, because I install software provided by them; I

choose to trust them in this way, and that’s a deliberate choice

If you are responsible for the operation of a system where the consequences of a security problem could be very important, you should do two things:

1. Define a security policy which defines who is allowed to do what on your system.

2. Seek competent advice on how to enforce your policy, detect breaches of that policy, and take account of any potential problems that might fall outside the scope of your policy.

11.2 Security Considerations for find

Some of the actions find might take have a direct effect; these include -exec and -delete. However, it is also common to use -print explicitly or implicitly, and so if find produces the wrong list of file names, that can also be a security problem; consider the case for example where find is producing a list of files to be deleted.

We normally assume that the find command line expresses the file selection criteria and actions that the user had in mind – that is, the command line is “trusted” data.

From a security analysis point of view, the output of find should be correct; that is, the output should contain only the names of those files which meet the user’s criteria specified on the command line. This applies for the -exec and -delete actions; one can consider these to be part of the output.

On the other hand, the contents of the filesystem can be manipulated by other people, and hence we regard this as “untrusted” data. This implies that the find command line is a filter which converts the untrusted contents of the filesystem into a correct list of output files.

The filesystem will in general change while find is searching it; in fact, most of the potential security problems with find relate to this issue in some way.

Race conditions are a general class of security problem where the relative ordering of actions taken by find (for example) and something else are critically important in getting the correct and expected result^{2} .

For find, an attacker might move or rename files or directories in the hope that an action might be taken against a file which was not normally intended to be affected. Alternatively, this sort of attack might be intended to persuade find to search part of the filesystem which would not normally be included in the search (defeating the -prune action for example).

11.2.1 Problems with -exec and filenames

It is safe in many cases to use the ‘-execdir’ action with any file name. Because ‘-execdir’ prefixes the arguments it passes to programs with ‘./’, you will not accidentally pass an argument which is interpreted as an option. For example the file -f would be passed to rm as ./-f, which is harmless.

However, your degree of safety does depend on the nature of the program you are running. For example constructs such as these two commands

     # risky
     find -exec sh -c "something {}" \;
     find -execdir sh -c "something {}" \;

2 This is more or less the definition of the term “race condition”

are very dangerous. The reason for this is that the ‘{}’ is expanded to a filename which might contain a semicolon or other characters special to the shell. If for example someone creates the file /tmp/foo; rm -rf $HOME then the two commands above could delete someone’s home directory.

So for this reason do not run any command which will pass untrusted data (such as the names of files) to commands which interpret arguments as commands to be further interpreted (for example ‘sh’).

In the case of the shell, there is a clever workaround for this problem:

     # safer
     find -exec sh -c 'something "$@"' sh {} \;
     find -execdir sh -c 'something "$@"' sh {} \;

This approach is not guaranteed to avoid every problem, but it is much safer than substituting data of an attacker’s choice into the text of a shell command.

11.2.2 Changing the Current Working Directory

As find searches the filesystem, it finds subdirectories and then searches within them by changing its working directory. First, find reaches and recognises a subdirectory. It then decides if that subdirectory meets the criteria for being searched; that is, any ‘-xdev’ or ‘-prune’ expressions are taken into account. The find program will then change working directory and proceed to search the directory.

A race condition attack might take the form that once the checks relevant to ‘-xdev’ and ‘-prune’ have been done, an attacker might rename the directory that was being considered, and put in its place a symbolic link that actually points somewhere else.

The idea behind this attack is to fool find into going into the wrong directory. This would leave find with a working directory chosen by an attacker, bypassing any protection apparently provided by ‘-xdev’ and ‘-prune’, and any protection provided by being able to not list particular directories on the find command line. This form of attack is particularly problematic if the attacker can predict when the find command will be run, as is the case with cron tasks for example.

GNU find has specific safeguards to prevent this general class of problem. The exact form of these safeguards depends on the properties of your system.

11.2.2.1 O_NOFOLLOW

If your system supports the O_NOFOLLOW flag^{3} to the open(2) system call, find uses it to safely change directories. The target directory is first opened and then find changes working directory with the fchdir() system call. This ensures that symbolic links are not followed, preventing the sort of race condition attack in which use is made of symbolic links.

If for any reason this approach does not work, find will fall back on the method which is normally used if O_NOFOLLOW is not supported.

You can tell if your system supports O_NOFOLLOW by running

     find --version | grep Features

3 GNU/Linux (kernel version 2.1.126 and later) and FreeBSD (3.0-CURRENT and later) support this

This will tell you the version number and which features are enabled. For example, if I run this on my system now, this gives:

     Features enabled: D_TYPE O_NOFOLLOW(enabled) LEAF_OPTIMISATION \
     FTS(FTS_CWDFD) CBO(level=2)

Here, you can see that I am running a version of find which was built from the development (git) code prior to the release of findutils-4.5.12, and that several features including O_NOFOLLOW are present. O_NOFOLLOW is qualified with “enabled”. This simply means that the current system seems to support O_NOFOLLOW. This check is needed because it is possible to build find on a system that defines O_NOFOLLOW and then run it on a system that ignores the O_NOFOLLOW flag. We try to detect such cases at startup by checking the operating system and version number; when this happens you will see ‘O_NOFOLLOW(disabled)’ instead.

11.2.2.2 Systems without O_NOFOLLOW

The strategy for preventing this type of problem on systems that lack support for the O_NOFOLLOW flag is more complex. Each time find changes directory, it examines the directory it is about to move to, issues the chdir() system call, and then checks that it has ended up in the subdirectory it expected. If all is as expected, processing continues as normal. However, there are two main reasons why the directory might change: the use of an automounter and someone removing the old directory and replacing it with something else while find is trying to descend into it.

Where a filesystem “automounter” is in use it can be the case that the use of the chdir() system call can itself cause a new filesystem to be mounted at that point. On systems that do not support O_NOFOLLOW, this will cause find’s security check to fail.

However, this does not normally represent a security problem, since the automounter configuration is normally set up by the system administrator. Therefore, if the chdir() sanity check fails, find will make one more attempt^{4}. If that succeeds, execution carries on as normal. This is the usual case for automounters.

Where an attacker is trying to exploit a race condition, the problem may not have gone away on the second attempt. If this is the case, find will issue a warning message and then ignore that subdirectory. When this happens, actions such as ‘-exec’ or ‘-print’ may already have taken place for the problematic subdirectory. This is because find applies tests and actions to directories before searching within them (unless ‘-depth’ was specified).

Because of the nature of the directory-change operation and security check, in the worst case the only things that find would have done with the directory are to move into it and back out to the original parent. No operations would have been performed within that directory.

11.2.3 Race Conditions with -exec

The ‘-exec’ action causes another program to be run. It passes to the program the name of the file which is being considered at the time. The invoked program will typically then perform some action on that file. Once again, there is a race condition which can be exploited here. We shall take as a specific example the command

     find /tmp -path /tmp/umsp/passwd -exec /bin/rm

4 This may not be the case for the fts-based executable

In this simple example, we are identifying just one file to be deleted and invoking /bin/rm to delete it. A problem exists because there is a time gap between the point where find decides that it needs to process the ‘-exec’ action and the point where the /bin/rm command actually issues the unlink() system call to delete the file from the filesystem. Within this time period, an attacker can rename the /tmp/umsp directory, replacing it with a symbolic link to /etc. There is no way for /bin/rm to determine that it is working on the same file that find had in mind. Once the symbolic link is in place, the attacker has persuaded find to cause the deletion of the /etc/passwd file, which is not the effect intended by the command which was actually invoked.

One possible defence against this type of attack is to modify the behaviour of ‘-exec’ so that the /bin/rm command is run with the argument ./passwd and a suitable choice of working directory. This would allow the normal sanity check that find performs to protect against this form of attack too. Unfortunately, this strategy cannot be used as the POSIX standard specifies that the current working directory for commands invoked with ‘-exec’ must be the same as the current working directory from which find was invoked. This means that the ‘-exec’ action is inherently insecure and can’t be fixed.

GNU find implements a more secure variant of the ‘-exec’ action, ‘-execdir’. The ‘-execdir’ action ensures that it is not necessary to dereference subdirectories to process target files. The current directory used to invoke programs is the same as the directory in which the file to be processed exists (/tmp/umsp in our example, and only the basename of the file to be processed is passed to the invoked command, with a ‘./’ prepended (giving ./passwd in our example).

The ‘-execdir’ action refuses to do anything if the current directory is included in the PATH environment variable. This is necessary because ‘-execdir’ runs programs in the same directory in which it finds files – in general, such a directory might be writable by untrusted users. For similar reasons, ‘-execdir’ does not allow ‘{}’ to appear in the name of the command to be run.

11.2.4 Race Conditions with -print and -print0

The ‘-print’ and ‘-print0’ actions can be used to produce a list of files matching some criteria, which can then be used with some other command, perhaps with xargs. Unfortunately, this means that there is an unavoidable time gap between find deciding that one or more files meet its criteria and the relevant command being executed. For this reason, the ‘-print’ and ‘-print0’ actions are just as insecure as ‘-exec’.

In fact, since the construction

     find ...  -print | xargs ...

does not cope correctly with newlines or other “white space” in file names, and copes poorly with file names containing quotes, the ‘-print’ action is less secure even than

‘-print0’.

11.3 Security Considerations for xargs

The description of the race conditions affecting the ‘-print’ action of find shows that xargs cannot be secure if it is possible for an attacker to modify a filesystem after find has started but before xargs has completed all its actions.

However, there are other security issues that exist even if it is not possible for an attacker to have access to the filesystem in real time. Firstly, if it is possible for an attacker to create files with names of their choice on the filesystem, then xargs is insecure unless the ‘-0’ option is used. If a file with the name /home/someuser/foo/bar\n/etc/passwd exists (assume that ‘\n’ stands for a newline character), then find ... -print can be persuaded to print three separate lines:

     /home/someuser/foo/bar
     /etc/passwd

If it finds a blank line in the input, xargs will ignore it. Therefore, if some action is to be taken on the basis of this list of files, the /etc/passwd file would be included even if this was not the intent of the person running find. There are circumstances in which an attacker can use this to their advantage. The same consideration applies to file names containing ordinary spaces rather than newlines, except that of course the list of file names will no longer contain an “extra” newline.

This problem is an unavoidable consequence of the default behaviour of the xargs command, which is specified by the POSIX standard. The only ways to avoid this problem are either to avoid all use of xargs in favour for example of ‘find -exec’ or (where available) ‘find -execdir’, or to use the ‘-0’ option, which ensures that xargs considers file names to be separated by ASCII NUL characters rather than whitespace. However, useful as this option is, the POSIX standard does not make it mandatory.

POSIX also specifies that xargs interprets quoting and trailing whitespace specially in filenames, too. This means that using find ... -print | xargs ... can cause the commands run by xargs to receive a list of file names which is not the same as the list printed by find. The interpretation of quotes and trailing whitespace is turned off by the ‘-0’ argument to xargs, which is another reason to use that option.

11.4 Security Considerations for locate

11.4.1 Race Conditions

It is fairly unusual for the output of locate to be fed into another command. However, if this were to be done, this would raise the same set of security issues as the use of ‘find ... -print’. Although the problems relating to whitespace in file names can be resolved by using locate’s ‘-0’ option, this still leaves the race condition problems associated with ‘find ... -print0’. There is no way to avoid these problems in the case of locate.

11.5 Summary

Where untrusted parties can create files on the system, or affect the names of files that are created, all uses for find, locate and xargs have known security problems except the following:

Informational use only Uses where the programs are used to prepare lists of file names upon which no further action will ever be taken.

‘-delete’ Use of the ‘-delete’ action with find to delete files which meet specified criteria

‘-execdir’

Use of the ‘-execdir’ action with find where the PATH environment variable contains directories which contain only trusted programs.

11.6 Further Reading on Security

While there are a number of books on computer security, there are also useful articles on the web that touch on the issues described above:

https://goo.gl/DAvh

This article describes some of the unfortunate effects of allowing free choice of file names.

https://cwe.mitre.org/data/definitions/78.html

Describes OS Command Injection

https://cwe.mitre.org/data/definitions/73.html

Describes problems arising from allowing remote computers to send requests which specify file names of their choice

https://cwe.mitre.org/data/definitions/116.html

Describes problems relating to encoding file names and escaping characters. This article is relevant to findutils because for command lines processed via the shell, the encoding and escaping rules are already set by the shell. For

example command lines like find ... -print | some-shell-script require

specific care.

https://xkcd.com/327/

A humorous and pithy summary of the broader problem.

12 Error Messages

This section describes some of the error messages sometimes made by find, xargs, or locate, explains them and in some cases provides advice as to what you should do about this.

This manual is written in English. The GNU findutils software features translations of error messages for many languages. For this reason the error messages produced by the programs are made to be as self-explanatory as possible. This approach avoids leaving people to figure out which test an English-language error message corresponds to. Error messages which are self-explanatory will not normally be mentioned in this document. For those messages mentioned in this document, only the English-language version of the message will be listed.

12.1 Error Messages From find

Most error messages produced by find are self-explanatory. Error messages sometimes include a filename. When this happens, the filename is quoted in order to prevent any unusual characters in the filename making unwanted changes in the state of the terminal.

‘invalid predicate ` -foo'’

This means that the find command line included something that started with a dash or other special character. The find program tried to interpret this as a test, action or option, but didn’t recognise it. If it was intended to be a test, check what was specified against the documentation. If, on the other hand, the string is the name of a file which has been expanded from a wildcard (for example because you have a ‘*’ on the command line), consider using ‘./*’ or just ‘.’ instead.

‘unexpected extra predicate’

This usually happens if you have an extra bracket on the command line (for

example ‘find . -print \)’).
‘Warning: filesystem /path/foo has recently been mounted’
‘Warning: filesystem /path/foo has recently been unmounted’

These messages might appear when find moves into a directory and finds that the device number and inode are different from what it expected them to be. If the directory find has moved into is on a network filesystem (NFS), it will not issue this message, because automount frequently mounts new filesystems on directories as you move into them (that is how it knows you want to use the filesystem). So, if you do see this message, be wary – automount may not have been responsible. Consider the possibility that someone else is manipulating the filesystem while find is running. Some people might do this in order to mislead find or persuade it to look at one set of files when it thought it was looking at another set.

‘/path/foo changed during execution of find (old device number 12345, new
device number 6789, filesystem type is <whatever>) [ref XXX]’

This message is issued when find moves into a directory and ends up somewhere it didn’t expect to be. This happens in one of two circumstances. Firstly, this happens when automount intervenes on a system where find doesn’t know how to determine what the current set of mounted filesystems is. Secondly, this can happen when the device number of a directory appears to change during a change of current directory, but find is moving up the filesystem hierarchy rather than down into it. In order to prevent find wandering off into some unexpected part of the filesystem, we stop it at this point.

‘Don't know how to use getmntent() to read ` /etc/mtab'. This is a bug.’

This message is issued when a problem similar to the above occurs on a system where find doesn’t know how to figure out the current list of mount points.

Ask for help on bug-findutils@gnu.org.
‘/path/foo/bar changed during execution of find (old inode number 12345, new
inode number 67893, filesystem type is <whatever>) [ref XXX]"),’

This message is issued when find moves into a directory and discovers that the inode number of that directory is different from the inode number that it obtained when it examined the directory previously. This usually means that while find was deep in a directory hierarchy doing a time consuming operation, somebody has moved one of the parent directories to another location in the same filesystem. This may or may not have been done maliciously. In any case, find stops at this point to avoid traversing parts of the filesystem that it wasn’t

intended to. You can use ls -li or find /path -inum 12345 -o -inum 67893

to find out more about what has happened.

‘sanity check of the fnmatch() library function failed.’

Please submit a bug report. You may well be asked questions about your system, and if you compiled the findutils code yourself, you should keep your copy of the build tree around. The likely explanation is that your system has a buggy implementation of fnmatch that looks enough like the GNU version to fool configure, but which doesn’t work properly.

‘cannot fork’

This normally happens if you use the -exec action or something similar (-ok and so forth) but the system has run out of free process slots. This is either because the system is very busy and the system has reached its maximum process limit, or because you have a resource limit in place and you’ve reached it. Check the system for runaway processes (with ps, if possible). Some process slots are normally reserved for use by ‘root’.

‘some-program terminated by signal 99’

Some program which was launched with -exec or similar was killed with a fatal signal. This is just an advisory message.

12.2 Error Messages From xargs

‘environment is too large for exec’

This message means that you have so many environment variables set (or such large values for them) that there is no room within the system-imposed limits on program command line argument length to invoke any program. This is an unlikely situation and is more likely result of an attempt to test the limits of xargs, or break it. Please try unsetting some environment variables, or exiting the current shell. You can also use ‘xargs --show-limits’ to understand the relevant sizes.

‘argument list too long’

You are using the ‘-I’ option and xargs doesn’t have enough space to build a command line because it has read a really large item and it doesn’t fit. You may be able to work around this problem with the ‘-s’ option, but the default size is pretty large. This is a rare situation and is more likely an attempt to test the limits of xargs, or break it. Otherwise, you will need to try to shorten the problematic argument or not use xargs.

‘argument line too long’

You are using the ‘-L’ or ‘-l’ option and one of the input lines is too long. You may be able to work around this problem with the ‘-s’ option, but the default size is pretty large. If you can modify the your xargs command not to use ‘-L’ or ‘-l’, that will be more likely to result in success.

‘cannot fork’

See the description of the similar message for find.

‘<program>: exited with status 255; aborting’

When a command run by xargs exits with status 255, xargs is supposed to stop. If this is not what you intended, wrap the program you are trying to invoke in a shell script which doesn’t return status 255.

‘<program>: terminated by signal 99’

See the description of the similar message for find.

‘cannot set SIGUSR1 signal handler’

xargs is having trouble preparing for you to be able to send it signals to increase or decrease the parallelism of its processing. If you don’t plan to send it those signals, this warning can be ignored (though if you’re a programmer, you may want to help us figure out why xargs is confused by your operating system).

‘failed to redirect standard input of the child process’

xargs redirects the stdin stream of the command to be run to either /dev/null or to /dev/tty for the ‘-o’ option. See the manual of the system call dup2(2).

12.3 Error Messages From locate

‘warning: database /usr/local/var/locatedb is more than 8 days old’

The locate program relies on a database which is periodically built by the updatedb program. That hasn’t happened in a long time. To fix this problem, run updatedb manually. This can often happen on systems that are generally not left on, so the periodic “cron” task which normally does this doesn’t get a chance to run.

‘locate database /usr/local/var/locatedb is corrupt or invalid’

This should not happen. Re-run updatedb. If that works, but locate still

produces this error, run locate --version and updatedb --version. These

should produce the same output. If not, you are using a mixed toolset; check your PATH environment variable and your shell aliases (if you have any). If both programs claim to be GNU versions, this is a bug; all versions of these programs should interoperate without problem. Ask for help on bug-findutils@gnu.

org.

12.4 Error Messages From updatedb

The updatedb program (and the programs it invokes) do issue error messages, but none seem to be candidates for guidance. If you are having a problem understanding one of these,

ask for help on bug-findutils@gnu.org.

13 History

The xargs and find programs have separate origins but are collected together in Findutils because they are often used together. While today they also share a small amount of implementation, this wasn’t originally the case. The locate program started out as a feature of find but today it is a separate program.

13.1 History of find

A find program appeared in Version 5 Unix as part of the Programmer’s Workbench project and was written by Dick Haight. Doug McIlroy’s A Research UNIX Reader: Annotated Excerpts from the Programmer’s Manual, 1971-1986 provides some additional details; you

can read it on-line at https://www.cs.dartmouth.edu/~doug/reader.pdf.

GNU find was originally written by Eric Decker, with enhancements by David MacKenzie, Jay Plett, and Tim Wood. The idea for ‘find -print0’ and ‘xargs -0’ came from Dan Bernstein.

13.2 History of xargs

The xargs program was invented by Herb Gellis at Bell Labs. In his own words:

Hi James, Thanks for reaching out. Yes I invented xargs way back before we even released UNIX to the general public when it was running on PDP-11 machines with little memory, and capable shell programs were not there yet - kind of like early IBM PC-DOS command lines. The name came about, first, by noticing at the time there were no commands beginning with ‘x’ (silly reason, I know), and then came up with, basically, "eXecute command with ARGumentS". This obviously allowed one to process files, sequentially, including batches of files, while the UNIX command line buffer was very tiny. I don’t remember exactly how small but possibly only 512 bytes. The very first use intended was to allow compiling C programs that were broken into many small routines whose total name length would exceed the command line buffer. Hope this settles the matter! Oh, another arcane factoid about xargs at the beginning was that I was able to keep it smaller than 4k (I think that was the amount) which at the time was the maximum size of a file segment on PDP-11/UNIX, so that the program could be loaded completely on the first segment, without having to go back and get chains of further segments - Hence fast!

In 2023, GNU xargs is unfortunately much larger, around 75KiB when stripped.

GNU xargs isn’t derived from the original Bell Labs program. It was originally written by Mike Rendell, with enhancements by David MacKenzie.

13.3 History of locate

4.3-BSD introduced the fast-find feature, in which the command find needle would look for a file named ‘needle’. This took advantage of the fact that, at the time, there was no valid two-argument find invocation. The implementation was much faster than searching the whole file system in real time, because it used a pre-built file name database. This functionality is described in more detail in Finding Files Fast by James Woods (Usenix ;login, Volume 8 Issue 1, pages. 8-10, 1983).

Standardisation of find led to this functionality being moved into the locate program in 4.4-BSD. The command find needle now unambiguously means “start searching at the file needle and print the names of the files you encounter”.

GNU locate and its associated utilities were originally written by James Woods, with enhancements by David MacKenzie.

Appendix A GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 https://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ASCII without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See https://www.gnu.org/licenses/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled `  ` GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

find Primary Index

This is a list of all of the primaries (tests, actions, and options) that make up find expressions for selecting files. See Section 2.1 [find Expressions], page 4, for more information on expressions.

!

! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

(

() . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

,

, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

–

-a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 -amin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 -anewer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 -atime. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -cmin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -cnewer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 -context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 -ctime. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 -daystart . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 -depth. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 -empty. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 31 -execdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 31 -executable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 -false. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 -files0-from . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 -fls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 -fprint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 -fprint0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 -fprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 -fstype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 -gid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 -group. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 -ignore_readdir_race . . . . . . . . . . . . . . . . . . . . . . . . 21 -ilname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -iname. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 -inum. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 -ipath. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -iregex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -iwholename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -links. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 -lname. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -ls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 -maxdepth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 -mindepth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 -mmin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -mount. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 -mtime. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 -newer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 -newerXY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -nogroup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 -noignore_readdir_race . . . . . . . . . . . . . . . . . . . . . . 21 -noleaf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 -not . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 -nouser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 -o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 -ok . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 -okdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 -or . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 -path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 -perm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 -print. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 -print0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 -printf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 -prune. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 -quit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 -readable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 -regex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -regextype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -samefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 -true. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 -type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -uid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 -used. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 -user. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 -wholename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 -writable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 -xdev. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 -xtype. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

A

ago in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 am in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

B

bugs, reporting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

D

day in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

find Primary Index 115

F

first in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 fortnight in date strings . . . . . . . . . . . . . . . . . . . . . . 55

H

hour in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

L

last day. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 last in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

M

midnight in date strings . . . . . . . . . . . . . . . . . . . . . . . 53 minute in date strings. . . . . . . . . . . . . . . . . . . . . . . . . . 55 month in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

N

next day. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 next in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 noon in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 now in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

O

option delimiter, -- . . . . . . . . . . . . . . . . . . . . . . . . . 66

P

parse_datetime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 pm in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

T

this in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 today in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 tomorrow in date strings . . . . . . . . . . . . . . . . . . . . . . . 55

W

week in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

Y

year in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 yesterday in date strings . . . . . . . . . . . . . . . . . . . . . . 55